A Complete Bibliography of Publications in Transactions of the American Fisheries

Society: 1960–1969

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

09 June 2023 Version 1.02

Title word cross-reference

14 [GW67, MZC66, RD65]. ^{35} [MZC66]. × [McC60, Sch62].

1-Naphthol [SMB67]. 10-gallon [Hal68]. 1107 [Abr65, Cha66]. 1620 [Fas67, PD67, Swi64a, Swi64b, SS66]. 1950-Brood [All69]. 1959 [Ano60a]. 1960 [Ano61a]. 1961 [Ano62c]. 1962 [Ano63b]. 1963 [Ano64a]. 1964 [Ano65b]. 1965 [Ano66a]. 1966 [Ano67c].

2 [MS67]. 222 [EB60, MK68].

3-Trifluormethyl-4-nitrophenol [AK62]. 3-trifluoromethyl-4-nitrophenol [Smi67]. 3200 [SH67]. 3600 [MC69, VR67].

4 [MS67]. 4-D [MS67]. 4-Styrylpyridine [HT64b].

709 [PG65]. 7090 [Dah66, PB67a]. 7090/7094 [Dah66, PB67a]. 7094 [Abr65, Dah66, PB67a].

Abatement [KM64]. Aberrant [MGC64]. Ability [CGE62, LD68, TBG69, Wea65]. Abnormal [BA65, Daw69, MB65]. Abnormalities [GF66]. above [MMPS69]. ABS [FSC66, ST63]. Absorption [Lau69, TBG69]. Abundance [IM62, Mac66, MB62, Rah67]. Acanthocephalan [Hna69]. Acceptable [MS67]. Acclimation [LD68, Tag69]. Account [Win66]. Accumulation [BHD^{+}64, Hos64, MI67]. Accuracy [Phi69]. Acetate [HS67]. Acid [DLS66, FB69, HAS69]. Acroneuria [CP68]. Activity [Fri65, KTW61, Mef68]. Acute [AYM69b, DLS66, Hun69, Kat61, SMB67, WB69]. Adding [Tan60]. Additional [Mou64a]. Adhesions [Fre65]. Adjacent [Ric67, Sha69b]. Adult [Bur69b, GV68, Hel66, Hna69, Moo60, Nig69, Pow64a, Rah68, Wag69b, Wea65, Wes64]. Advanced [App66, Cho64]. Adverse [FG66]. Aequipecten [Hof68]. aestivalis [Mar69]. Africa [Cok68]. African [SF63]. after [Pow65, Tau63]. Agassiz [II66, Tar69]. agassizi [Hil68]. Age [All66a, Bai64, Bjo61, Boo65, BL60, Bul60, Bur69a, CW61, Cha66, CS64, Coo67b, Eds60, Mar69, May69, MM60, MS64, Mra64, Nor67, Par69, Pri69a, Rah65, Ree67a, Reg62b, Ric67, RH68, Rot63, SD63, VR67, Wat64b]. Age-Group [Par69]. Ages [Van61]. Aholehole [MNB65]. Air [Fra68, FB69, Tal64]. Alabama [BT62, Bro65, BC65b, HC60, Swi60]. alalunga [Bea69]. Alarm [Ree69]. Alaska [Alt65, HSH62, Hea62, LW62, Lar64, MS62, ND65, NB65, Pow64a, Ree67a, Roo60, Tho62]. Albacore [Bea69, Joh62]. albidus [Mar67]. Albinism [Daw67b]. Albinistic [Joh68b]. aleuticus [McL68]. Alevin [Gun67]. Alevins [DB64, KS65]. Alewife [Nor67, Rot63]. Alewives [Bro61b, Hav61, Lac69]. Algonquin [Mar66]. Alien [McD68]. Alkaline [Cro64, McC62, MT68]. Alkyl [LM63, ST63]. alletteratus [BJ67]. Allometric [PG64]. along [Goo67, Ket68]. Alosa [GV68, Mar69, Nor67, Rot63, Wea65]. alpinus [EW65]. Alsea [WWC63]. Alteration [SS62a]. Alterations [Els68]. Alternate [CWSM60]. Alternate-Year [CWSM60]. Alternative [Van69]. ambloplitis [HP65]. America [Hub68, Ket68, McF69, Sha69a]. American [Ano60a, Ano61a, Ano62c, Ano63b, Ano64a, Ano65b, Ano66a, Ano67c, Ano62a, Ano63a, Bai64, Bur69a, Che65, Daw67b, GV68, Gor61, HA64, Hod66, Rot61, Rup65, SFH69, Sak66a, Sak66b, Wal61, Wal63, Wea65]. americanus [Sch62, BSH65, SFH69, SP69a]. Amia [HW67]. Among [Alb69, Cro64, GC62, Het68]. Anadromous [Dav67, Hav61, MS62]. Analog [Sil69]. Analysis [Dah66, Pat69, Roe65, SS66]. Anesthetic [HT64b, MK68]. Angler [LS60, Ste66, Von61]. Anglers [She62]. Angling [AH69, Mac66, SA62, SA65, SA66, WCH62, Win66]. anguillicaudatus [Sch60]. Animal [Sea66]. Animals [JM63]. Annual [Ano60a, Ano61a, Ano62c, Ano63b, Ano64a, Ano65b, Ano66a, Ano67c]. annularis [Han65]. Annuli [BC63b, Hat61]. Annulus [SP69a]. Anomalies [MH68]. anomalum [LP60]. Antimycin [Fin69a, HGC67, PB67b]. Aplodinotus [But65, Pri66, Pri67]. Apostle [Dry66, DB68]. Apparatus [BF68b]. Apparently [BS69]. Appearance [SE60]. Application [Kee65, RH69, SH67, Von61]. Applicator [Del68]. Applied [GG65]. Applying [Jon66]. Aquaculture [TYII69]. Aquaria [BW68, GW67, Hal68]. Aquarium [Har65, OMK67, RPS68, Sti67]. Aquatic [Gas62, Gil66, HD62, OS62a, Smi67, ST63, TJ62, Ver69]. Arctic [EW65]. Area [BSR69, Cok68, McC61b, SS62c]. Areas [Man64, Wat64a]. Arizona [Joh68b]. Arkansas [BHH66, Col66, Mil67]. Arnica [Ben60]. Aroma [BST61, BSR69]. Array [MPS67]. Arsenite [Cow65, Gil66]. Artificial [BH68, Het68, Kle66, MB62, Sch62, SN68, SH62b, Web62]. Ascend [CGE62]. Ascorbic [HAS69]. Aspects [HH69, II60, SM69, Tar69, Wes64]. asper [McL68]. Assay [PYS68]. Assessing [GG65]. Assessment [Ghe66]. Associated [All65, Bro61b, JW62, ML60, Ree67b, TEB69]. Association [Ghe66]. Associative [PBK69]. atherinoides [Fuc67]. Atitlan [PB67b]. Atlantic [Ano65b, Bea69, Boy61, CS61, Dov68, Hav60, Lew65, Lew66, LH68, Log68, MP60, Mei62, Mur68, Ric67, Roe66, Sti67]. atlanticus [Hof68]. atratulus [Tar69, Hou64]. auriculatus [HD65]. Automatic [BSH69, BC67, SLN68]. Autopsy [Mou64b]. Availability [Joh62, LGLB61]. Available [Whi69b]. Avoidance [Han69, WWD60].

Bacteria [BS69]. Bacterial [PH69, RNBH60]. Bags [Ver69]. Bait [Chi60, She60, Sho64]. Balance [HC60]. Barb [ER60]. Barb-Type [ER60]. Barracuda [SM61, SM61]. Barriers [Sai62]. Base [SP69b]. Based [HG64]. Bass [AH69, AM67, BT62, Bro65, BC63b, Cha67a, CS61, Cla68, Clu64, Col66, CHA63, Dov68, Faj62, Gre62, Hun60, Jer69, Kel62, KS60, KS62, LD68, Lau69, LGLB61, Lew62, LH64, LB66, MSS62, Moo60, RT67, Str61]. Basses [App66]. Bathymetric [Dry66]. Bay [Ber61, Bud60, Chi60, Eds60, Gor61, Hof68, OG68, Poo69, Ren63, Roe65, Pyc61, Ric67, Sai62, Sha69b, SP69a, TJ62]. Bayfield [McC61b]. be [HG61]. Beach [Rob67]. Beam [CDF68, MD68]. Beaver [AC60, BHH66, KB69]. before [Rah67]. Behavior [BF68a, Bro61a, BJ67, Cro62b, Cum63, EG65, Faj62, FSC66, HSH62, Hel66, Lam65, OWW69, OS62b, PR69, PBK69, Sie68, Whi69a]. Benthos [Hil65, MB62]. Benzene [LM63]. Benzine [ST63]. Benzocaine [MK68]. Bering [Lar64]. Bertalanffy [Abr65, All66b]. Between [Gri66, GFR63, Lux60, Nil63, Pod61]. Beverton [PG64]. Bight [Cla68]. Bigmouth [McC67]. Binomial [Ghe66]. Biochemical [OW63]. Biochemistry [Das61]. Biological [LGS^{+}62, WV63, Wes64, ZEK65]. Biology [Daw67a, Daw68, Gra61, HW69, II60, Mar66, SM66]. Biotope [Mac66]. Birth [Hit62]. Biscayne [Roe65]. Black [Alb69, DM65a, HP65, Isa65, II66, Sie69]. Black-Spot [HP65]. Blacknose [Hou64, Tar69, Wag69a]. Bleeding [ACP66]. Bloater [DB68, WB63]. Bloaters [Wel66]. Block [RT67]. Blood [BB66, BS69, Cro64, Das61, EE66, FB69, LS61, Swa69]. Blue [Erd62, Erd68, FW62a, Fis65, Min62, Tag69, Wer67]. Blue-gill [Wer67]. Blueback [EW65]. Bluefish [DCM66]. Bluegill [Gro68, Lam63, LM63, Pat68a, Sch67]. Bluegills [ASV66, Gil66, Reg62b]. Board [Woo69]. boating [AS67]. Body [BSR69, Hor66, Rei63, WH61]. Body-Fat [Hor66]. Body-Scale [WH61]. Bog [MZC66]. Bonito [Bos66]. Bordering [Sch68c]. borealis [HW69]. Borgne [SS62c]. Bottle [Tal64]. Bottom [Joh61, Log63, Tet60, WK61, Wat62, Wat64a, Sha69a]. botulinum [SG67]. Bowfin [HW67]. Box [LHK68]. Brains [OW63]. Branch [SA66]. Branchiostegal [Bul60]. Branding [GJ69]. Breeding [DM61]. Brevoortia [Cha67b, CG60, Coo65b, Dah69, Het68, Lew65, Lew66, LH68, MH68, RW63, TR67]. Bridger [Log63]. Brine [BSH69, VC62]. Bristol [Nor69a]. Brood [All69]. Broods [CI68]. Brook [AC60, Bro61a, Cho64, Coo61, Coo67a, Eco63, EW69, FW62b, FW64, Fra68, FB69, Gre64, Hal68, HH69, Hat61, Hun65a, Lat62, Lat63, MBD67, Mei62, Mom65, Mom67, MAR68, Nel68a, Pod61, RVIS68, SS62a, SS62b, Sch65b, She61, Vin60, Web62]. Brooks [HSH62, Hea62]. Brown [BV69, BC60, BB67, Eip63, Jen69, Joy61, KS65, MC62, MCA62, Swi63]. Brownlee [MMPS69]. Bubble [Ren63, Wes64]. Budworm [Cop61]. Buffalo [McC67]. bulboglossa [Hou64]. Bull [KB69, HD67]. Bulldozed [WP60]. Bullhead [DM65a, FB66, Joy61]. Bullheads [Alb69]. Burbot [BM60]. Burkenroad [Cum61]. Burning [CO63]. Bush [Cok68]. Bush-Cleared [Cok68]. Butoxyethanol [MS67]. Buzzards [Hof68].

C [GW67, MZC66, GW67, RD65]. Caddis [Hoo60]. Cadmium [HS64]. Caeca [App66]. Cage [EM69]. Calculated [WLWY60]. Calculi [Tal67]. California [CM69b, RT67, Rei63, She68a]. Callinectes [Tag69]. calva [HW67]. Campostoma [LP60]. camtschatica [Gra64, Pow64a]. Canada [Log68]. canadense [Nel68c]. Canadian [BC62, PC67]. canadum [Ric67]. Cane [LNCT66]. Cantor [Sch60]. Capacity [HA64]. caprodes [MAR68]. Captive [Cum63]. Capture [BB66, Chu63, HK67, MD68]. Carapace [Per65, RS63]. Carolina [Daw62, FW62a, Lew62, RH68, RH69]. carolinus [Fin69b]. Carp [Das61, Joh68b, Kev66, NS60, Rob62, SM66, Sta62]. carpio [Beh69a, Joh68b, SM66]. Carpiodes [Beh69a]. Carpsuckers [Beh69a, Beh69b]. Cartridge [Del68]. Cartridge-Fed [Del68]. Cases [Daw69]. Casoron(R) [WB69]. Cataract [All63]. Catch [All69, CDF68, Fis65, LS60, PB67a, RC61, Sil69, SF63, SH67, Wag69b]. Catch-Sampling [Fis65]. Catch-to-Escapement [All69]. Catchability [AH69]. Catchable [LSL63, Von61]. Catchable-Sized [Von61, LSL63]. Catches [HD67, ML60]. Catching [Fab68]. Catfish [Alb69, BF67, HR68, Hub63, May69, Min62, Rol68, Sch68b, Sch68c, She63, Wel67]. Catostomid [MGC64]. Catostomids [GN68]. Catostomus [Hau69, PH69]. Caught [Kle66]. caurinus [HD65]. caused [Mou64b]. Cavefish [Hil68]. Cayuga [WLWY60]. CDC [MC69, SH67]. Cellulose [HS67]. Census [JW62, NW63].

Censuses [Reg62a]. Central [Cok68]. Centrarchid [HP65, WWD60]. Centrarchids [WH66]. Centrarchus [CW66]. Certain [Bar69, Ghe66, Gri66]. Cestode [NB65]. Chain [Mil62]. Chalfonte [Ano65b]. Chalfonte-Haddon [Ano65b]. Chambers [BM67]. Changed [She63]. Changes [Bee61, BH68, Dav68, DB68, HH67, Jen69, KM64, LS60, Par69, Pyc61]. Channel [Alb69, BF67, Els68, Hub63, Sch68b, She63, Wel67, Wes64]. Channelization [MLRW68]. Channels [SP66a]. Chapman [Pau62]. Char [Aca68, Nil63, EW65]. Characteristic [Coo65a]. Characteristics [Bar69, BB66, Cru69, Nel68b, OPA68, Ryd64, Sie69, TEB69]. Chelated [Cup69]. Chemical [Cru69, Hun69, KTW61, MR69, Ryd64, TEB69, WYHW60]. Chemistry [EE66, HA64]. Chemosterilant [SC69]. Chesapeake [Ric67, Sha69b]. Chicot [Mil67]. Chignik [ND65, Roo60]. Chinook [Bon68, DM61, DB64, EB60, Gun67, Nig69, Nor69b, Par69, Ray68, Ray69, Rob67, RPR60, SWD63, TBG69, TEB69, Wes64]. chittendeni [Daw68]. Chlorinated [JM63]. Choice [LGLB61]. Cholesterol [Gri66]. Chologaster [Hil68]. Chromatographically [PYS68]. Chronic [Bru69, DB64, Low64]. Chronology [DM62]. Chrosomus [Phi69]. Chub [Gra61, VK69]. Chum [GF66, Lan63, SN68]. Cichlid [Hub68, SF63]. Cisco [Smi64]. City [Ano65b, Ano67c]. Cladocerans [SC66]. Clark [Fuc67, McC67, Nel68c, Swe68]. clarki [Low65, Low65, MO63, Sto69, UD66]. Class [For61, Pyc61]. Classes [FS63, KS62]. Clayton [KCN62]. Clear [CWSM60, CW61]. Cleared [Cok68]. Clearing [HD65]. Clearwater [Ano60a]. Clinostomum [UD66]. Clipped [SH69]. Clostridium [SG67]. clupeaformis [Boo65, Eds60, Fen64, MF60]. Coast [Ket68, All67, BK65, CS61, Goo67, Log68]. Coastal [Dav67, FW62a, Hit62, Low65, Sum62]. Coastrange [McL68]. Coasts [Dah69, Joh62]. Cobia [Ric67]. Cod [FR63, Jen67]. Coefficients [Reg62a]. Coelomic [FG66]. Coho [All65, All67, All69, Cha65, DB64, GJ69, HAS69, HM69, HWD62, Laa69, Roo60, SWD64, UPH65]. Cold [FH68]. Cold-Water [FH68]. Collapsable [EM69]. Collect [Jer69, LM61, Man64]. Collected [For61]. Collecting [PC67]. Collection [BR68, BA65]. Color [Jen69]. Colorado [Ano61a, DM65b, FR69, HT64a, HK67, VK69]. Coloration [Daw69, PJS^{+}66]. Columbia [Par69, BR68, Ray68, Ray69]. Combined [MPS67]. Comment [Ano67a]. Commercial [All67, Roe66, SP61, SF63, SP66b, Van69]. Commercially [Sil69]. commersoni [PH69]. Committees [Ano62a, Ano63a]. Common [Bos66]. Comparative [AK62, Das61, FW64, KC68, LED67, MBD67, Swi60]. Compared [RD65]. Comparison [And62, Gre64, Hec69, MC62, MK68, RM66, Rol68]. Comparisons [SH62b]. Complement [HM67, LED67]. Complex [DLS66]. Components [PYS68]. Composition [Bai64, Hun69, SH67]. Compounds [Dah63]. Computer [Abr63, MC69, Pau69, Sil69, Wal69]. Concentration [DFWD63, HWD62, Hor66, SWD64, Swi63]. Concentrations [AE66, BF68b, DLS66, FB69, Gil66, Low64, RT67, WWD60]. Condition [Lux60, Rei63, Sak66a]. Conditioned [Tar64]. Conditions [CO63, Eip63]. Confidence [Dav64, ML60]. Conifer [SK65, SKO66]. Conifer-Groundwood [SKO66]. Connecticut [Mar69, Ric62, Wal63]. Connecting [Joh61]. Consequences [Tan60]. Conservation [Lin63]. Constituents [OW63]. Consumption [Bea64, Bla61, EW69, Hun60, MS66, Mag69, MNB65]. Contagion [GG65, Ghe66, Sea66]. Containing [DLS66]. Contamination [LGS^{+}62, LNCT66]. Content [EP65, Mou61, Mou64a]. Contents [Dox63, JF60, Jer69]. Continental [Gra68]. Contingents [Cla68]. Continuous [BM67, SLN68]. Contribution [All67, War62]. Contributions [Daw67a, Daw68, NW62]. Control [BM66, JM63, WL63, VR67]. Controlling [Mou61, Mou64a]. Conversion [Gam63]. Convict [Rei63]. Cope [SM69]. Copepod [BHH66, Dah69]. Copper [WCH62]. Coregonine [GC62]. Coregonus [Boo65, DB68, Eds60, Fen64, MF60, WB63, Wel66]. Corner [Ghe66]. Corner-Association [Ghe66]. Correlation [Fas67]. Correlation-Regression [Fas67]. Coryphaena [RH68, RH69]. Cottid [Hea65, Pat62]. Cottus [McL68, PC66]. Counties [SA66]. Counting [BC67, Mur68]. Counts [App66, Bro65]. County [MSS62, MGC64, SA62, Sum62, Tar69]. Courtship [BF68a]. Cove [OG68, Mei62]. Cover [BH68]. Covered [PC67]. Crab [FW62a, Gra64, Per65, Pow65, RS63, Rop61, Tag69]. Crabs [Fis65, Pow64a]. Cranberry [MZC66]. Crane [Wer67]. Crappie [Han65, Sie65, Sie68, Sie69]. Crassiphiala [Hou64]. Crassostrea [Sak66a, Sak66b]. Crawfish [MO63]. Crawford [SA66]. Crayfish [Mom67]. Creek [Ben60, Log63, Rei63, She68a, SA62]. Creel [JW62]. Crepidostomum [KOB69]. Criteria [Glu66]. Criticisms [Ghe66]. Critique [Rot63, Ver69]. Croaker [Dov68, Isa64]. Croakers [MP60]. Crooked [Nel68a]. Crops [IB63, RD65, Tur60]. Crosses [SN68]. Crow [MSS62]. Cryptobia [BK65, BK66]. crysoleucas [LFB68, Sum64, Sum66]. Culture [CM69a, HPS61, Sha69a]. Current [LM61, WI69]. Curve [Abr65]. Curves [All66b, RC61]. Cutlassfish [Daw67a]. Cutthroat [Ben60, DM65b, Low65, Sto69, Sum62, UD66]. Cutting [Gas62]. Cuvier [SS62c]. cyanellus [GFR63]. Cyanides [DLS66]. Cycle [Erd68, Sak66b]. Cyclopsetta [Daw68]. cylindraceum [Nor69a]. Cynoscion [SS62c]. Cyprinidae [Phi69]. Cyprinodon [MI67]. Cyprinodontidae [RM66]. Cyprinus [Joh68b, SM66].

D [Fas67, MS67]. Dace [Hou64, Tar69, Wag69a]. Dakota [Fuc67, Swe68]. Dam [BR69, SPJ61]. Dams [BR68]. Danish [Dah63]. Daphnia [Gal67]. Dart [AB66b]. Data [Beh69a, Fis65, PBK69, VR67, Whi69a]. Day [Ray69]. DDT [AE66, BKA63, BHD^{+}64, Cop61, CW65, GW67]. DDT-C [GW67]. DDVP [Kon69]. Debris [MMPS69]. Decalcifying [Sch68b]. Decimated [GB69]. Deepwater [Smi64]. Deflecting [BJ61]. DeLury [Von61]. Densities [Rob62]. Densitometric [Hor66]. Denuded [Wat64a]. Denver [Ano61a]. Deposits [CS67]. Depression [Nor69b]. Depth [Cok68, HT64a, Lac68, Mur62, Wel66]. Description [DCM66, Jen69]. Desert [McC63a]. Destruction [ES61]. Detect [Cum63]. Detergent [Eis65]. Determinant [Ste66]. Determination [All66a, DM62, Fli69]. Determinations [Mar69]. Determine [Bul60, Kee65, Kem68]. Determined [Lew62]. Determining [Hor66, Rot63, Von61, Wat64b]. Detoxication [KCN62]. Developing [OS62a, SWD63]. Development [Cha67b, CW66, FS60, GH60, Kni63, Lew66, Mou61, Pri64, Pri66, Pur61, Rid60, Sie65, Wat60]. Device [BC67, BM67, KM65, Man64, SLN68, VC62]. Diameters [Lew62]. Diatoms [Man67]. Diazinon [MZC66]. Diazinon-C [MZC66]. Dichlobenil [WB69]. Dichromatism [BF68a]. Diel [FR69]. Dieldrin [GW67]. Dieldrin-C [GW67]. Dietary [Pod61]. Diets [Cup69, She63]. Differences [Cro64]. Different [PHL62, Rob62, She62, SWD63]. Differential [Bro61b, Fra68, Mac66]. Digestion [AB66a, Hun60, KW68, Mag69]. Digestive [SM64]. Dimensions [HA64]. Dinosaur [VK69]. Diphyllobothrium [NB65]. Diquat(R) [WB69]. Direct [LM61, MP61, PYS68]. Direct-Current [LM61]. Direction [CDF68]. Disappearance [Bri61, MP60]. Discussion [Wil69]. Disease [AYM69a, RNBH60, RPR60, Wes64]. Disk [Jon66]. Dispersal [Hun65a, Pri63a, Sch60]. Dispersion [ES60]. Displaced [Hel66]. Dissolved [Cob61, MS61, Mou61, Mou64a, RT67, SWD63]. Dissolved-Oxygen [MS61, Mou61]. Dissolving [Zur63]. Distribution [All66a, CCF67, CG60, Cok68, Cus63, Dry66, For61, For63, GG65, HT64a, KCN62, Lac68, Lam63, Lam65, MMPS69, Qad61, Wel66]. Distributions [All66a]. Ditch [CW65]. Diurnal [BB65, DM65a, KTW61]. Divers [HV68]. Documentation [Fre68]. Doe [MGC64, Tar69]. Dolly [Aca68, AB66a, AB66b, BA65, LW62, ND65, Ree67a, TT67]. Dolphin [RH68, RH69]. Domestic [FW62b, FW64, Gre64, MBD67]. Domestication [Vin60]. Dominant [BH68]. Donna [TJ62]. Doppler [Cum63]. Dorosoma [KB69, Mil67]. Dorsal [BA65, WKH63]. Double [Woo69]. Downstream [BJ61, MS62, MMPS69, Par69, SPJ61]. Downstream-Migrating [MMPS69]. Drainage [Pri63a, SP69b]. Drainages [She68b]. Drawdown [HCR69]. Drift [She68a, Tal64, Wat64a]. Drop [Fra68, FB69, LHK68]. Drops [Tal64]. Drugs [Cho64]. Drum [But65, KC68, Pri66, Pri67, Pri69a, Swe68]. Drums [Zur63]. Dry [BC67]. Dumeril [SF63]. duorarum [Cum61, II60]. Duration [LGS^{+}62, Sta62]. during [Gra64, Hun65a, Joh61, Lat62, Lau69, She62]. Dwarfed [Fen64]. Dyes [Kel67]. Dynamics [FLM66].

E. [McC60, Sch62]. Early [DCM66, FS63, HH69, Hel69, Mar67, Nel68c, Pri64, Pri66, Pur61, Sch60, Sie65]. Earthworms [She60]. East [BC65a, ZE61]. Eastern [RVIS68, Gra68, Tur69]. Ecdysis [Gra64]. Ecological [KS60, MT68, MC62, Pat69]. Ecology [DL69, Fin69b, MI67, Nor69a, RB60]. Economic [Ste69, Van69]. Economical [Spe69]. Economist [Ste69]. Editorial [Ano67a]. Editor’s [Edi63]. Effect [Bla61, BB67, BHD^{+}64, Chu63, Cum63, Eip63, FG66, Hil65, Hir66, HT64b, KDO63, Lew65, LH68, MS66, MR62, Mur62, NB65, Ray69, Smi67, TJ62, TBG69, Wag68, Wag69b, Wei61, WCH62]. Effectiveness [She60, Sho64]. Effects [ASV66, Bar69, BH61, BF67, CDF68, Car66, CM69b, Cop61, Cow65, CW65, CO63, DB64, Eis65, EE66, EW69, FSC66, FS65, Fri65, Gas62, Gil66, Hil67, Hub68, JG64a, JG64b, Jen67, KTW61, KOB69, KS65, Laa69, LD68, LGS^{+}62, LM63, Lew66, MB62, MI67, Mom67, MLRW68, NS60, Sch68a, She63, SA62, She67, SKO66, Sok69, SC69, ST63, VAE68, WB69, ZEK65]. Efficiency [BJ61, Ber61, MPS67]. Effort [Fis65]. Egg [Joh61, SP66a, She61]. Eggs [BM66, BC67, CS67, DB64, Hil67, HD65, KS65, Man64, Mar67, PYS68, RP69, Rot61, Rup65, She60, Sie68, SK63, WP67]. Eight [MM60]. Eighty [Ano60a]. Eighty-Ninth [Ano60a]. Elasmobranch [HS67]. Electric [JK69, LM61, MD68, Vib63]. Electrode [MPS67]. Electrofishing [BC63a, MB65]. Electrophoresis [HS67]. Electrophoretic [Rol68]. elegans [WL63]. Elements [Hos67, Sok69]. Eleven [AK62]. Elimination [GW67, PB67b]. Elk [CCF67]. Elongation [Bon68]. Embiotoca [Isa65, II66]. Embryo [Nel68b]. Embryological [Hec69]. Embryonic [DCM66, Kni63]. Embryos [CP68, Cob61, PC66, SWD64, SWD63]. Emerald [Fuc67]. Emergence [Fre68]. Emetics [Jer69]. Emphasis [RB60]. Employing [SL68]. Endrin [Bri61, EE66, FLM66, FB66, KC61, LFB68]. Ends [FR63]. Endurance [MS66]. Energy [Lau69]. England [HJ62, JW61, Lux69]. Environment [Bri61, Gil66]. Environmental [Bee61, Joy61]. Environments [Ver69, WS62]. Ephemeridae [Fre67]. Ephemeroptera [Fre67]. Epinepheline [RB60]. Epipelagic [Lar64]. Epizootics [WL63]. Equation [PG64]. Equilibrium [Pat68a]. Eradicant [HGC67]. Eradicator [Mey65]. Erie [Bee61, Man64, Wol63]. Errata [Ano67b]. Error [All65]. Errors [JW62]. erythrogaster [Phi69]. Escape [Pat68b]. Escapement [All69]. Esox [BC65a, FS60, FS63, JM69, McC60, McC61a, McC62, Sch62]. Established [LSL63]. Establishing [MS67]. Establishment [Sch60]. Ester [MS67]. Estimate [Beh69b, Fis65, Pau62, Wat60, Wat62]. Estimated [Kev66]. Estimates [Dav64, Joh65, Lat63, MC69, Phi69, RR66, RD65]. Estimating [All65, Beh69a, HG61, HD67, Reg62a, Reg62b, Ryd65, Sch68a]. Estimation [CJ68, Coo66, IM62, SH67, Wal63]. Estimations [Col66]. Estuarine [Eis65, Mur68, ND65]. Estuary [LW62, Tho62]. Ethyl [AYM69b]. Eugenol [Tar66]. Euthynnus [BJ67, NY61]. Evaluation [AB66b, CWSM60, Cup69, Sok69, Swi60, Van69]. Everglades [Cro62a]. Evidence [Fen64]. Evidenced [CS61]. Exchange [Cob61, Joy61]. Excision [Bon68, She67]. Exercise [WI64]. Expected [HG61]. Expenditure [Lau69]. Experimental [BV69, Ber61, Eco63, GC62, HG64, MB62, Mey65, OMK67, UPH65, WP60]. Experimentally [GB69]. Experiments [Pau62, RR64, SLN68]. Exploitation [BC62, Coo67a, Hir66, Mar66, She67]. Exploited [Reg62a]. Exposed [Ell69]. Exposure [GW67, Low64]. extensa [DL69]. Extension [McC61a]. Extinction [Tar64].

Facilities [Web62]. Factor [AH69]. Factors [AYM69b, Ben60, FS63, KS60, Lew69, LS60, SP61, Sok69, SH67]. Failure [BC63b, Glu66]. falcatus [Fin69b]. Fall [And62, Mil62, SA62, TBG69]. Fall- [And62]. False [Hat61]. farionis [KOB69]. Farm [BKA63, Cow65, IB63, LGS^{+}62, Tur60]. Farming [LNCT66]. Fat [EP65, Hor66]. Fathead [Bru69, Fli69, MS66, MT68]. Fauna [TJ62, Tet60, WK61, CM69b]. Faunas [McD68]. Features [Hir66]. Fecundity [Beh69a, IW69, IT68, Kel62, KB69, LB66, Nor67, Phi69, Rah65, Ric67, SFH69, SS62c]. Fed [Del68]. Feeder [BSH69]. Feeding [AW68, BJ67, DM62, Dav68, FSC66, Goo67, Gre62, Kev66, Nak65, SM64, VC62, Cus63]. felis [HR68]. Female [Per65]. Fence [Mur68]. Fertilization [Het68, HB64a, Kem68, MB62]. Fertilized [Hil67, RP69]. Fertilizer [Tan60]. Fiber [CS67, KS65, MS66, SK65, SKO66]. Fibers [SK63]. Field [Cou64]. fifth [Ano66a]. Filter [Cus63]. Filter-feeding [Cus63]. Fin [All65, All69, BB67, Chu63, She67, SH69]. Fin-Clipped [SH69]. Fin-Marked [All69, All65]. Fingerling [Bur69b, Cho64, Pod61, She61]. Fingerlings [App66, Bla63, Bro61a, Chu63, EB60, She67, SK65, TEB69]. Fins [BA65, WKH63]. First [FW64, Hun65a, KS60, Lat62, Ano62c]. First-Year [KS60]. Fish [Arn66, BST61, BSR69, BR69, Bjo61, BB65, BC67, Bri61, BL60, Bro61c, BM67, CHB63, Cok68, Cop61, Cou64, CO64, Cru69, Cum63, Del68, DLS66, Dox63, Dry66, EE66, Els68, EM69, ER60, FS65, FCKP60, GW67, Gri66, GB69, HA64, HV68, HCR69, Her68, HGC67, Hil67, Hna69, HC60, HPS61, HT64a, Hor66, HT64b, Hub61, IB63, JK69, Joh65, Jon66, Kee65, Kel67, KM65, KM64, Lac68, LSL63, Lew69, Mac66, McC63a, McD68, Mey65, MZC66, MS61, Mou64b, MS67, NW63, OW63, Pat68b, PYS68, PKW63, PB67b, Reg62a, Rid60, RR66, Roe65, Ryd65, SF69, Sch68a, Spe69, SF63, SP66b, Sto69, Wal69, WI69, Wei61, Woo69, Yas63, War62, WF63]. Fished [Sil69]. Fisheries [Abr63, Ano60a, Ano61a, Ano62a, Ano62c, Ano63a, Ano63b, Ano64a, Ano65b, Ano66a, Ano67c, CS61, Chr69, Dah63, De 69, Gra68, Ket68, Log68, McF69, Pat69, Pau69, Van69, WI69]. Fishermen [Kle66, MRC64]. Fishermen’s [Whi69a]. Fishery [All67, Car63, Chi60, Erd62, GH60, Har62, Pur63, Pyc61, RH69, Sok69, Ste66, War62, WF63]. Fishes [ACP66, AK62, Bap66, BH66, DM62, Dav67, Daw65, Eis65, GC62, HHC66, HW67, HP65, Hoo60, Joh68a, KC61, Lar64, LED67, LSL63, LHK68, MM60, MS62, MGC64, Moe67, OMK67, PBK69, PHL62, RB60, Ree67b, Ree69, Ren63, RD65, SM64, Tur60, VAE68, VC62, WP67, WWD60, Whi69b, WI64].

Fishing [AS67, BC65b, HK67, Kem68, Pow64a, Rup61, Sch68a, SA65, Sho64, Ste69, Vib63, Wal60]. Fishways [CGE62]. Fishwheel [Mee61]. Fission [Bap66]. Fitness [HW67]. Fitting [All66b]. Five [Dah69, For61, HB64b, SH62b, Tan60]. Five-year [HB64b]. Fixed [BF68b]. Flag [HV68]. Flagfish [FSC66]. Flatfishes [Daw69, Lux69]. Flathead [May69, Sch68c]. flavescens [RNBH60]. Flavor [BST61, BSR69]. Flesh [EP65]. Flier [CW66]. Flies [Hoo60, SA62]. Flies-Only [SA62]. Floods [EW69]. Florida [Ano60a, TJ62, Clu64, Dah69, Fin69b, Hun60, Moo60, Roe65, SF63]. floridae [FSC66]. floridanus [Clu64]. Flounder [BSH65, Daw68, Hos64, Lux60, OWW69, Pea62, Poo69, Sai62]. Flounders [Hos67]. Flow [BF68b, Har65, SLN68]. Flow-Through [BF68b]. Flowing [Zur63]. Flows [Ray68]. Fluid [FG66]. Fluke [KOB69]. Flume [BV69]. Fluorescent [PMD67]. Fluoride [NS60]. Fluvial [IT68]. following [Der63, GW67, Hun69, KM64, Nor69b]. fontinalis [Cho64, Coo67a, FW62b, FW64, FB69, Hal68, HH69, RVIS68, SS62b, Sch65b, Vin60, Web62]. Food [AC60, Alt65, AM67, BW68, BM60, Bor63, Bur69a, CCF67, Cro62a, Cus63, DET65, ES60, EW69, Gam63, Gor61, Hea65, Hou67, Hun60, Hun65b, Laa69, Lac69, LGLB61, Mag69, Man67, Mar66, McC67, Mil67, Mom65, MAR68, Nak65, ND65, Pri63b, Pri67, Pri69b, Qad61, RB60, SD63, Sie68, SP69b, SP66b, Swe68, WB63, WT65]. Food-Habits [Bor63]. Foods [Hel69, Min62]. Forage [HW67, LH64]. Forecasting [Sil69]. Foreword [Edi63]. Form [Fen64, BC63b]. Format [Swi64a, Swi64b, SS66]. Formation [KS62, SP69a]. formula [Cup69]. Formulation [AYM69b]. Fort [Col66, Hou65]. Fortran [All66a, All66b, VR67, Abr65, Cha66, Dah66, Fas67, PG65, PB67a, PD67, SH67, Swi64a, Swi64b, SS66, VR67]. Fortran/Format [Swi64a, Swi64b, SS66]. Four [BST61, BL60, FR63, Hat61, MB62, Sch67]. Fourteen [BC65b]. fourth [Ano65b]. Free [SP66b]. Frequencies [Sto69]. Frequency [Dah66]. Fresh [BST61, Gro60, Mur62, WYHW60]. Fresh-Water [BST61, Mur62]. Freshwater [BK65, But65, Car63, Kel67, KDO63, KC68, LED67, McF69, Pri66, Pri67, Pri69a, Swe68]. Freshwaters [PKW63]. Fright [Ree69]. Fructose [Gre68]. Fry [App66, CWSM60, Cho64, CS67, DM65b, Fab68, HSH62, Hil67, LD68, Pat62, Roo60, Str61, Tho62, WV63, Wer67]. Function [LGLB61]. Fundulus [CS64, Coo65a, MK69, RM66]. Further [Han65, Hod66]. Furunculosis [Eco60, Her68]. Future [Dav69].

gairdneri [BM66, BB66, HM67, HK67, Jen69, MPS67, Ree67a, SG67, UPH65, UD66, WWC63]. gairdnerii [KOB69, Wag68, Wag69b]. Galeichthys [HR68]. gallon [Hal68]. Galveston [Chi60, Ren63, ZE61]. Game [Bjo61, BL60, Bro61c]. Gar [Hun60, NW62, Pri63a]. Garcin [NY61]. Garman [MK69]. Gars [Goo67]. Gas [MI67, Ren63, Wes64]. Gas-Bubble [Ren63, Wes64]. Gastric [KW68]. Gatewells [BR68, BR69]. Gear [For61]. Gel [PYS68]. Generalized [SF69, Wal69]. Generation [SH62a]. Genus [CG60, Hos64, MI67]. Geographic [PYB63]. Georgian [Ber61]. Giant [NY61, PB67b]. gibbosus [Hec69, KW68]. Gibson [Hou65]. Gila [VK69]. Gill [Ber61, Har62, Hea62, HT64a, HG64, Lac68, MF60, McC61b, Wer67]. Gill-Net [Har62, McC61b]. Girard [MO63]. Girth [Lan63]. Girth-Length [Lan63]. Given [Reg62a]. Glacial [Ryd64]. Gladstone [MSS62]. Globulin [Boo65]. Gmelin [SP69a]. Goals [Chr69]. Goderich [McC61b]. Golden [LFB68, Sum64, Sum66, WL63]. Goldfish [Fri65]. gorbuscha [SE60, SH62a]. Gradient [Mac66]. Granby [FR69]. Grand [Ano63b]. Graphic [Dav64]. Gray [Cro62a]. Great [SM61, LP60, Poo69, Pri63a, Van61]. Grebe [PB67b]. Green [GFR63, Pyc61, VK69]. griseus [Cro62a]. Ground [WT65]. Groundwater [Lat65]. Groundwood [SK65, SKO66]. Group [Par69]. Groups [Cro64]. Growth [Abr65, Alb69, All66b, Bai64, BSH65, BC60, Bjo61, Bon68, BL60, BT62, BB67, Bud60, Bur69a, CW61, Cha66, Chu63, CS64, Clu64, Coo61, CHA63, Coo67b, Coo67a, Cro62a, Der63, DB68, Eds60, GFR63, Hav60, HWD62, Hir66, HJ62, Hou65, Jen67, Kee65, KS60, Laa69, May69, MM60, Mra64, Nor67, Nor69b, PG64, Pri69a, Pri69b, Rah65, Rah67, Ree67a, Reg62b, Ric67, RH68, SD63, SM64, She68a, She63, She67, SWD64, SH69, Str61, Sum62, Swe68, Swi63, Van61, VR67]. Growth-Curves [All66b]. grunniens [But65, Pri66, Pri67]. Guatemala [PB67b]. Guide [LP66]. Guiding [MPS67]. Gulf [Bos66, Daw63, Daw67a, Daw68, Daw69, Goo67, RW63, Tur69, Cha67b, CG60, Daw65, HW69, Het68, Lar64, Moe67]. Gull [Bur69a]. Guppy [SC69]. Guthion(R) [Mey65]. Gyrodactylus [WL63].

Habitat [Els68, Fin69a, Nav62, OWW69, SS62a]. Habitats [Jen69]. Habits [Alt65, AB66b, BM60, Bor63, Bur69a, Goo67, Mar66, McC67, McL68, Mom65, Nak65, RB60, SM64, WT65]. Haddock [WT65]. Haddon [Ano65b]. Hake [JF60, Mar67]. Halibut [JW61]. Hampshire [Nor69a]. Hand [Fra68]. Handling [Bro61b]. Harbor [BC65a, Sak66a, Sak66b]. Harvest [Bjo61, MSS62, McC63a, MR62, Sch65a, Von61]. Hatcheries [Dav69]. Hatchery [All63, And62, BK66, Bro61c, Bue61, BS69, Coo61, Eip63, Har67, Hav60, HM69, Lat63, Rei63, She67, Spe69, TT67, WWC63, War62, WYHW60]. Hatchery-Reared [And62, Bro61c, Bue61, Hav60, TT67, WWC63, War62, Rei63]. Hawaii [Sak66a, Sak66b]. Hawaiian [NY61]. Head [Coo65a]. Headwaters [GB69, HB64b]. Healthy [BS69]. Hebgen [Gra61]. Helminth [BHH66]. Hematological [Bar69, TEB69]. Hematopoietic [AYM69a]. Hemoflagellate [BK65]. Hemosiderin [Gro68]. Hepatoma [AH63]. Heptachlor [ASV66]. Herbicides [Hil67, WB69]. Hermaphrodite [Aca68, MS64]. Hermaphroditism [RYW63]. Herring [Boy61, RP69, Sti67, Tho62, Wat64b, Wat65]. heteroclitus [Coo65a, RM66]. Hetersomata [Daw67b]. heudeloti [SF63]. Hexagenia [Car66, Fre67, Fre68]. Hexamitiasis [UPH65]. High [AN64, CGE62, Fre65, RT67]. High-Seas [Fre65]. Highlands [WT65]. Hill [MK69, SM69]. Hilsa [IT68]. Hilton [Ano66a]. Hiodon [Van61]. hippurus [RH68, RH69]. Hirudinella [NY61]. Histological [WYHW60]. History [DL69, EW65, FS63, Fuc67, HH69, Hau69, II66, Mey62, MK69, Nel68a, Nel68c, NW62, Nor69a, Pyc61, Ryd64, SM69, Tur69, VK69]. Hogchokers [Daw62]. Holding [BM67, EM69, KM65]. Holt [PG64]. Homarus [SFH69]. Homing [Cro62b, Faj62, OS62b, OS63, Sto69]. Homology [Sum66]. Horseshoe [Rop61]. Horsetooth [HT64a]. Host [BK66, Dah69, Hna69, PYB63]. Hotel [Ano64a, Ano65b, Ano66a]. hoyi [DB68, WB63, Wel66]. Human [Rol68]. Humpbacked [MH68]. Humper [Rah65]. Hunt [SA62]. Huron [Ber61, Bud60, BC62, Gor61, McC61b, Tet60]. Hurricane [Sai62, TJ62]. Hybrid [Dah69, Etn68, Het68]. Hybridization [GC62, SN68, WH66]. Hybrids [McC60, Nel68b, Sch62, SH62b]. Hydra [ES61]. Hydraulic [CO64]. Hydrobiology [HPS61]. Hydrocarbons [JM63]. Hydrocyanic [DLS66]. Hydrogen [BF67]. Hydrolytic [SMB67]. Hydropsyche [Car66]. Hydrothol [Ell69]. Hyperplasia [RVIS68]. Hypoxic [Hun69].

IBM [Abr65, All66a, All66b, Dah66, Fas67, PG65, PB67a, PD67, Swi64a, Swi64b, SS66]. Ice [Dox63, PC67]. Ictalurus [Alb69, BF67, DM65a, FB66]. Idaho [BM64, BCC68, Cup69, OPA68]. Identification [Coo65a]. II [Abr65, Dah66, Fas67, JW61, LGS^{+}62, PG65, PD67, Swi64b]. II-D [Fas67]. ilisha [IT68]. Immature [Ben60, Boy61, Car66]. Immunology [LED67]. Impounded [Wes64]. Impoundment [McC63a, BHH66]. Impoundments [But65, MLRW68, Mur62, Ray68]. Impressions [Cou64]. Improve [SS62a]. Improved [WK61]. Improvement [Nav62, WP60]. Incidence [AW68, Dah69, Hub61, MGC64, NY61, Wag69a]. Incomplete [BA65]. Inconnu [Alt65]. Increase [Boo65]. Increased [HCB66]. Increases [Rob62]. Increasing [Yas63]. Incubation [Joh61, SP66a, Sie68]. Index [CJ60]. Indian [Das61, GH60]. Indiana [Nel68a, Wer67]. Individual [Alb69, HV68]. Indus [IT68]. Infection [PH69]. Infections [BK65]. Infectious [Eco63, PYB63]. Inferences [Whi69a]. Infestation [Wag69a]. Influence [Bea64, BB65, BB66, Bur69b, Cob61, DFWD63, Dov68, Faj62, HWD62, HK67, Mac66, SWD64, Swi63]. Influences [Vin60]. Influencing [AYM69b, AH69, Ben60, FS63, Lew69, LS60]. Information [Mou64a, OS62a]. Inheritance [SN68]. Inhibition [Bon68]. Injected [Kel67, ZEK65]. Injection [ACP66, Arn66, GN68]. Inland [Dah63]. Insect [CM69b, Cus63, SC69]. Insecticide [LGS^{+}62, SMB67]. Insecticides [Car66, JG64a, JG64b, Kat61, Kon69, MO63, PHL62, Wei61]. Instantaneous [PD67]. Institute [TYII69]. Intake [BR68]. Intakes [BR69]. Interaction [Nil63]. Interactions [McD68]. Intergeneric [WH66]. Interrelationships [Lac69]. Intervals [Dav64]. Intestinal [KOB69, MI67]. Intragravel [PC66]. Intralacustrine [Wer67]. Intravascular [ACP66]. Introduction [Lea63, MM60, Sch66]. Introductions [McD68]. Invertebrate [Wat62]. Invertebrates [Smi67, ST63, WB69].

Investigations [Har62]. Involving [Fen64, Ver69]. Iodine [Eco63]. Iowa [CWSM60, CW61, May69, Mey62, Sch68c, Wel67]. Iron [Dah63]. irradians [Hof68]. Irradiation [DB64]. Island [Pow64a, BSH65, Cla68, Coo65b, Coo66, Coo67b, OG68]. Islands [Dry66, DB68, Erd68, Nak65, RB60]. Isle [Rah65]. Isopod [Dah69, TR67]. Items [LGLB61]. IV [Abr65, All66a, All66b, Cha66, Dah66, PB67a, SH67, VR67].

Jackson [Ano63b]. jacksoni [Isa65, II66]. Jaw [Arn66, Bon68, She67, SH69]. Jaw-Injection [Arn66]. Jaw-Tagged [SH69]. Jellyfishes [PBK69]. Jersey [Ano65b]. Joaquin [Cha67a, RT67]. John [Ray69]. Johns [Moo60]. Jordanella [FSC66]. Journal [Win66]. Jr. [Ano62d]. Juvenile [BR68, BR69, Bla61, Cha65, DFWD63, Gro60, GN65, Hea65, HWD62, Min62, MPS67, OG68, Ray69, She68a, Tar64, Tar66, UPH65, Wag69b]. Juveniles [RM66].

kansae [MK69]. Kansas [Ano67c, MK69, SM69]. Kariba [Cok68]. Katadidymus [Gun67]. Katsuwonus [Mag69, Nak65]. Katz [BK65]. Kentucky [MGC64, Tar69, Tur60]. Keynote [McF69]. Keys [All66a]. Kidney [BS69]. Killifish [CS64, MK69]. Kinetics [CW65]. King [Gra64, Pow64a, Pow65, RS63]. kisutch [All67, GJ69, UPH65]. Kodiak [Pow64a]. Kokanee [CCF67, FR69]. Kuhlia [MNB65].

L [FS60, Rop61, Swi63]. L. [FS63, SM66]. Labidesthes [MAR68, Nel68a]. Laboratory [Cou64, HH69, Kon69, ST63]. Lac [Qad61]. Lacépède [Clu64, CW66, Kel62, KS60]. Lactic [FB69]. Lagoon [ZE61]. Lake [All63, And62, AE66, Bai64, Ber61, BCC68, Boo65, Bud60, BC62, Bue61, Bul60, Bur69a, BHD^{+}64, CWSM60, CW61, CCF67, Dav67, Der63, DM65b, DET65, Dry66, DB68, Eds60, EP65, Fen64, For63, Fuc67, Gor61, Gra61, HSH62, Har67, HA64, Hea62, Hil65, HB64a, Hou67, Joh68b, JM69, KCN62, Kem68, Lac69, LSL63, LHK68, LS60, MSS62, Man67, Man64, Mar66, McC60, MF60, McC61b, McC63b, MR62, NB65, Nel68a, Nor69a, OS63, PB67b, Qad61, Rah65, Rah67, Rah68, Rid60, Sch66, SP69a, SK64, Sta62, Swe68, Tau63, UD66, WLWY60, Wel66, Wer67, Bee61, BT62, Bue61, Cok68, Col66, DET65, EP65, FH68, Hav60, HH67, Joh61, Man68, McC67, McC63b, Mil67, Mom65, Mra64, Nel68c, Nor67, Pri63b, Pri67, Pri69a, Pri69b, Pur63, Rah67]. Lake [Rah68, Rob67, SE60, SH62a, Smi64, Sto69, SS62c, SM66, Tet60, TT67, WB63, Wol63]. Lakes [BGHK61, BC65b, Etn68, HG61, HA64, HC60, KDO63, LM61, Lat63, McC62, Pat68a, RD65, Ryd64, Ryd65, SP61, Tan60, War62, WF63, WI69, Wat60, Bjo61, Pri63a, Van61]. Lamprey [HS64, HT64b, Hub61, Man68, Smi67, WS62]. Lampreys [AK62, Dav67, Man67, MD68, Rah67, Sta62]. Landed [Lux60]. Landlocked [AE66, Bro61b, Dav67, IW69, WH61, War63]. Lands [GH60]. Large [CS61, Coo65b, GN68, JM69, OMK67]. Largemouth [AH69, App66, AM67, Clu64, Col66, CHA63, Hun60, Jer69, Kel62, KS60, KS62, Lau69, LGLB61, LH64, MSS62, Moo60, Str61]. Larvae [ES61, HS64, Hea65, JM63, Lau69, Man68, Sie69, WS62]. Larval [AK62, CM69a, DCM66, Kni63, Lew65, Lew66, Man67, Mar67, MD68, Nel68b, Nor69b, Sta62, Wel66, Hos67]. Larvicide [Smi67]. Latex [GN68]. Latin [HG64]. Latrobe [Coo65b]. Lawrence [SM66]. Lead [FCKP60]. Leech [Wag69a]. Legal [She62, Sok69]. Legal-sized [She62]. Leiostomus [Low64]. Length [All66a, Bon68, BC65a, Bur69a, Dah66, Daw62, Daw63, Daw65, Isa64, Lan63, Lux60, Lux69, Per65, RS63, Swi64a, Swi64b]. Length-Frequency [Dah66]. Length-Width [Per65, RS63]. Lengths [WLWY60]. Lentic [WS62]. Lepisosteus [Goo67, NW62, Pri63a]. Lepomis [ASV66, GFR63, Hec69, KW68, LM63, Pat68a, SH62b]. lepturus [Daw67a]. Lernaeenicus [Dah69]. less [Pea62]. LeSueur [Clu64]. Lethal [MR69]. Level [Cro64, DB64, HA64]. Levels [Boo65, Gri66]. Lewis [Fuc67, McC67, Nel68c, Swe68]. Life [DL69, EW65, FS63, Fuc67, HH69, Hau69, Hun65a, II66, Lat62, Mey62, MK69, Nel68a, Nel68c, NW62, Nor69a, Sta62, SM69, Tur69, VK69]. Light [Coo67a, Whi69b]. like [RPR60]. Limestone [Zur63]. Limited [Laa69]. Limits [MS67, ML60]. Limnetic [Fab68, Hea65]. Limnology [MLRW68]. Limulus [Rop61]. Lindane [GW67]. Lindane-C [GW67]. Linear [PG65]. Linnaeus [BF68a, Coo66, Coo67b, EW65, NY61, RH68]. Liparis [Hof68]. Literature [Joh68a]. Little [BJ67, Wel67]. Live [BSH69, Lux60, Sho64]. Loach [Sch60]. Lobster [SFH69]. Lobsters [Kha64]. Localities [KC68]. Location [Wag69b]. Loch [Sak66a, Sak66b]. Locomotor [Mef68]. Logbooks [Whi69a]. Logperch [MAR68]. Long [Cla68, GN68, JG64b, SKO66, Hav60, Hav61]. Long-Term [JG64b, SKO66]. Longevity [Coo67a, Rop61]. Longnose [NW62]. Lookdowns [OG68]. Loss [Gra64, LHK68, RR66]. Lost [Mom65]. Louisiana [LNCT66, SS62c]. Louver [BJ61, MPS67]. Loven [Sch66]. Low [DB64, Sto69, WWD60]. Low-Level [DB64]. Lower [SK64]. lucius [BC65a, FS60, FS63, JM69, McC60, McC61a, McC62, Sch62, VK69]. Lures [Kle66]. Lutjanid [RB60]. Lutjanus [Cro62a].

M. [Clu64]. M.S. [EB60]. macrochirus [Hec69, ASV66, Hec69, LM63, Pat68a]. Macroinvertebrates [HD62]. macropterus [CW66]. maculatus [Daw62]. Main [MLRW68]. Maine [And62, AE66, Dav67, EW65, Hav60, Hav61, HW69, Kel62, Lac69, Mei62, Rot61, WP60, War62, WF63, WI69]. Maintaining [BF68b]. majalis [CS64, Coo65a, RM66]. Major [BR68, Els68]. Making [Cou64]. Malathion [MS67]. Male [Bon68, Per65, Pow64a, TEB69]. malma [AB66b, LW62, Ree67a]. Man [Sum66]. Management [BGHK61, BC65b, Glu66, JM69, Mar66, Mee61, Pau69, PKW63, Van69, Wal60]. Manipulation [HCR69, LP66]. Many [OS63]. marginatum [UD66]. marina [NY61]. Marine [Bap66, De 69, EE66, Fin69a, Gra68, Kel67, Ket68, OMK67, SMB67, Swa69, TYII69]. marinus [KTW61]. Mark [Cho64, FCKP60, MC69, RR64, SH69]. Marked [All69, For63, RR66, Tau63, Wel67, All65]. Markers [WV63]. Marking [Arn66, GN68, GN65, HS64, Hos67, JK69, Kel67, Nig69, PMD67]. Marl [HB64a]. Marlin [Erd62, Erd68, Tal67]. Marquesan [NY61]. Marquesas [Nak65]. Marsh [CW65]. Martin [BT62]. Maryland [CS64, SD63]. Mass [Fre67, Fre68, PMD67]. Mass-marking [PMD67]. Mass-Rearing [Fre67]. Massachusetts [Hof68, MM60]. Mastigophora [BK66]. Materials [Bri61, WP67]. Mats [Gas62]. Maturation [Cum61, PD67]. Mature [Lat65]. Maturity [Bai64, Kel62, Lew62, Rah65]. Maxillary [Bon68]. Maximum [HG61, Rah67]. Mayflies [Fre67, Fre68, Hoo60]. McNary [BR69, SPJ61, Wes64]. Meade [MGC64, Tar69]. Means [ML60, PC67]. Measured [HH67]. Measurement [Rup61, Ste69]. Measuring [MNB65, Woo69]. Mechanism [SL68]. Medians [ML60]. Medusae [PBK69]. Meeting [Ano60a, Ano61a, Ano62c, Ano63b, Ano64a, Ano65b, Ano66a, Ano67c]. melas [Alb69, DM65a]. meleagris [Tar69]. Memoriam [Ano60b, Ano62b, Ano65a, Ano69a]. Memphis [Ano62c]. Mendota [HH67]. Menhaden [Cha67b, CG60, Coo65b, Dah69, Het68, Lew65, Lew66, LH68, MH68, TR67]. Menhadens [Tur69]. Menidia [DL69]. Mercuric [HS64]. Mercury [AYM69b]. Meristic [Bro65, SH62b]. Merluccius [Mar67]. Meshed [Hea62]. Metabolism [Pod61]. Metabolites [BKA63]. Metal [DLS66]. Metastasis [AH63]. Methanesulfonate [Bla61, EB60]. Method [ACP66, App66, Beh69a, Bor63, Cou64, GN68, GG65, Ghe66, HG61, Hor66, Joh65, Kee65, Kem68, Kev66, MS67, MNB65, Reg62b, Rot63, Ryd65, Sch68b, SP66b, Von61, Wat62]. Methods [BB66, Fre67, HK67, RP69, Sch68a, Ver69]. Methyl [CI68]. Mexican [Daw68]. Mexico [Bos66, CG60, Daw63, Daw65, Daw67a, Daw68, Daw69, KCN62, Moe67, Nav62, RW63, Tur69]. Miami [TYII69]. Michigan [AS67, Bur69a, Man67, MRC64, Mra64, Nor67, Pat68a, Sch60, She60, She61, SA62, She62, SA65, SA66, She67, She68b, Smi64, Sta62, Tan60, WB63, Wel66, Win66]. Microhabitats [BV69]. Microhematocrit [LS61]. Microinjecting [WP67]. microlepis [MS64]. Micropterus [AM67, BHH66, Clu64, Kel62, KS60, Lau69]. Microsporidian [Sum64]. Mid [Ric67]. Mid-Atlantic [Ric67]. Midwater [HD67]. Midwestern [KC68]. Migrant [BJ61, LP66, SPJ61]. Migrating [HSH62, MMPS69]. Migration [AB66b, BCC68, CS61, Coo66, IT68, Moe67, MR62, Par69, RT67, Ray68, Ray69, SF69, Sum62, WWC63]. Migrations [Bea69, Cha67a, MS62]. Minerals [Cup69]. Minimum [Lew65]. Minneapolis [Ano64a]. Minnesota [Ano64a, CS67, Etn68, Joh61, JM69, JM63, LS60, MSS62, OS63, Sch66, SM64, SP61, SK63, SK64]. Minnow [Bru69, Fli69, MS66, MT68, Phi69]. Minnows [Han69, HCB66]. Mirex [VAE68]. Misgurnus [Sch60]. Mississippi [But65, Car66, Fre68, Hub61, Hub63, Sch65c, Goo67, Hoo60, PBK69, Sch68c]. mississippiensis [BC63b]. Missouri [Ano67c, FH68, Hel69, McC67, MLRW68, NW62, Pur63]. Mitchill [Bai64, Bur69a, EG65, HH69, Mar69, Rot61, SS62b, Sch65b, Vin60]. Mixed [Bap66]. Mixing [Mur62]. Model [MZC66, SF69, Wal69]. Models [Pau69]. Modification [LS61]. Modifications [Tal64]. Moines [May69, Mey62]. Molecular [DLS66]. Montana [BL60, Gra61, Hau69, Log63]. Montmorency [SA62]. Montreal [Cus63]. Monument [VK69]. Mooneye [Van61]. mordax [Bai64, Bur69a, Rot61]. Morone [SP69a]. Morphological [McL68]. Morphometry [KC68]. Mortalities [SPJ61]. Mortality [BB66, Bro61b, HK67, IM62, Kle66, Lat62, Lat63, Mac66, Mou64b, PD67, Poo69, Pow64a, Reg62a, Ren63, RC61, She68a, Sie68, Wal61]. Mosquito [JM63]. Mosquitofish [FLM66]. mossambica [AN64, CI68]. Motel [Ano67c]. Mountain [Hau69]. Mountains [LP60]. Mounting [Wat65]. Movement [Bro61c, FW62a, For63, Hav60, Hub63, JW61, Low65, PC66, Sch65c, SWD64]. Movements [BM64, Cla68, Cum63, FR69, HHC66, Pow64a, Rah68, Sai62, She68b, Wel67, Wer67, Wol63]. Moxostoma [Mey62]. MS [MK68]. MS-222 [MK68]. Muehlebach [Ano67c]. Multi [Pau62]. Multi- [Pau62]. Multiple [AH63, HD62]. Multiple-Plate [HD62]. Municipal [BH61]. Munising [Eds60]. Muscle [CJ60]. Muskellunge [Gam63]. Mussels [MZC66]. Mycteroperca [MS64]. Mysis [Sch66].

Naiads [JG64a, JG64b]. namaycush [Bul60, McC63b, WLWY60]. Naphthol [SMB67]. Narcotization [EB60]. Narragansett [OG68, Sai62]. natalis [FB66]. natans [SK63]. National [Ano63b, Cro62a, LP60, VK69, Wal60]. Native [Bro61c, McD68]. Natural [BGHK61, Etn68, Lat63, OWW69, PJS^{+}66, Wal61, War63]. Naturally [CJ60]. Navigational [But65]. near [CDF68, CS64, SD63]. Nebraska [McC60, McC62, MT68, MLRW68]. nebulosus [SS62c]. Necrosis [Eco63, PYB63]. Nematode [NB65]. Nephropsids [Roe66]. nerka [BCC68, FR69]. Nesting [Han65]. Nests [Man68]. Net [Cha65, Fab68, Har62, HG64, McC61b, ML60, Mur68, RD65]. Nets [Ber61, Hea62, HT64a, Lac68, MF60, Pat68b]. Neurophysiology [Vib63]. Newfound [Nor69a]. Newfoundland [Hod66]. Newly [Dav67]. Newly-Transformed [Dav67]. Nicollet [Ano64a]. Night [LM61]. Ninetieth [Ano61a]. Ninety [Ano62c, Ano63b, Ano64a, Ano65b, Ano66a, Ano67c]. Ninety-fifth [Ano66a]. Ninety-First [Ano62c]. Ninety-fourth [Ano65b]. Ninety-Second [Ano63b]. Ninety-sixth [Ano67c]. Ninety-third [Ano64a]. Ninth [Ano60a]. nitrophenol [AK62, Smi67]. Noise [Sto69]. Normal [Aca68, SA65, SA66]. Norris [CHB63]. North [Ket68, McF69, Ryd65, HA64, Hir66, Hub68, Lar64, Lew62, Pow65, RH68, RH69, Sha69a, SA66, TJ62]. North-temperate [Ryd65]. Northern [BW68, BC65a, Clu64, Eco60, FS60, FS63, HW69, Hin61, Hun65b, McC61a, McC62, PR69, TT67, Bos66, Daw63, Daw67a, Daw68, Daw69, Man67, Pyc61, RW63, WP60]. Northwest [BHH66, KC61].

Northwestern [Hir66]. Note [FS60]. Notemigonus [LFB68, Sum64, Sum66]. Notes [II66]. Notropis [Fuc67, SM69]. Numerical [FS63].

Objects [KM65, MP61]. Observation [BF68a, Ree67b]. Observations [BW68, Hal68, Han65, Jen69, JF60, Lam65, Log63, MT68, PBK69, RB60, Sch68c, She68b, Wea65]. Observed [WLWY60]. Occupied [BV69]. Occurrence [Bos66, Gre68, Gun67, Hat61, Hof68, Mil62, MZC66, Rob67, RW63, TR67]. Occurring [CJ60]. Ocean [All65, Bea69, Lar64]. Off [Sha69a, Erd62, Erd68, JW61, Joh62, MZC66]. Off-bottom [Sha69a]. Offshore [Man64, Mar67]. oglinum [RP69]. Ohio [KM64, BC65a, Min62]. Oklahoma [Hou65]. Old [AM67, MAR68]. Olencira [Dah69, TR67]. oligactis [ES61]. Oncorhynchus [All67, BCC68, FR69, GJ69, LW62, Nig69, SE60, SH62a, SN68, UPH65]. One [BA65, Hna69]. Oneida [For63, Hou67]. onitis [BF68a, Coo66, Coo67b]. Only [SA62]. Ontario [Mar66, McC63b, Ryd64, SP69a, SM66]. Open [Cup69]. Open-formula [Cup69]. Opinion [Hub68]. Opinions [MRC64]. Opisthonema [RP69]. Optimum [Wal63]. Oregon [Ano66a, Cha65, CCF67, Hit62, IB63, Joh62, Low65, Sum62, WWC63, Wag68]. oregonensis [MR69, PR69]. Organic [JG64a, JG64b, Kat61, PHL62, Wei61]. Organisms [LH64, Log63, SMB67]. Organophosphorus [Car66, Kon69]. Orientation [BJ67, Sto69]. Orifices [BR69]. Origin [CW66, LGS^{+}62]. Osage [Pur63]. Osmerus [Bai64, Bur69a, Rot61]. osseus [NW62]. Other [Dav67, BA65, Bri61]. Otoliths [Wat64b, Wat65]. Otsego [SA66]. Otter [Roe65]. Ova [FG66]. Over-winter [Rei63]. Overcrowded [HC60]. Overhead [BH68]. Ovum [Lew62]. owned [BC65b]. Oxygen [Bea64, Bla61, BB65, Cob61, DFWD63, HWD62, Hil68, MS66, MS61, Mou61, Mou64a, MNB65, SWD64, SWD63, Swi63, WWD60, WI64, ZEK65]. Oxygenation [Spe69]. Oyster [Sak66a, Sak66b, SD63, Sha69b, Sha69a]. Ozark [MAR68]. Ozarks [Pur63]. Ozone [BM66].

P [MR69]. Pacific [Lar64, All67, BK65, DFWD63, Hir66, KC61, Ket68, Pow65]. Pacifica [CP68]. Paddlefish [Hou65, Pur61, Pur63]. Pallas [Nor69a]. Pancreatic [Eco63, PYB63]. Pandalus [HW69]. Paper [CS67, SKO66]. Paralichthid [Hos67]. Paralichthys [Hos64]. Paralithodes [Gra64, Pow64a]. Parameters [BV69]. Parasite [Sum64]. Parasites [BHH66, NB65]. Parasitic [TR67]. Parasitism [Dav67]. Parathion [MZC66]. Parathion-S [MZC66]. Parent [Wal63]. Parents [Gre64]. Park [Ano63b, Mar66, Ben60, Cro62a, LP60]. Parks [Wal60]. Part [LGS^{+}62, WB69]. Partial [Daw67b]. Particular [MB65]. Passage [BR69, Bla63]. Passamaquoddy [Boy61]. Passing [CO64]. Past [Sha69a]. Pathology [Ell69]. patronus [Cha67b, RW63]. Pattern [Coo65a, Hir66, Jen69, LP66]. Patterns [FS60]. Pearl [Sak66a, Sak66b]. Pectoral [BB67, Sch68b]. Pelagic [Hou67, OMK67]. pelamis [Mag69, NY61, Nak65]. Pelleted [ES60]. Pellets [SP66b]. Penaeid [CM69a]. Penaeus [CDF68, Cum61, II60]. Pennsylvania [BC60]. Perception [Tar66]. Perch [Gal67, HH67, Isa65, II66, RNBH60, SP69a]. Percina [MAR68]. Percy [Ano62d]. Peres [RNBH60]. Performance [Cup69, DFWD63, Gro60]. Periodicity [DM65a, KTW61, Lat62]. Permanent [GJ69]. Permanently [RR66]. Permit [Fin69b]. perplexus [PC66]. Persistence [BKA63]. Pesticide [LNCT66, SP66b]. Pesticide-Free [SP66b]. Pesticides [Han69, HCB66, Joh68a, SC66]. petenense [KB69, Mil67]. Peters [AN64]. Petersen [Dav64, RR64]. Petromyzon [KTW61]. Phalanx [BJ67]. Philonema [NB65]. Phosphamidon [Kon69]. Phosphatase [Cro64]. Phosphate [AYM69b]. Phosphorus [PHL62, Pod61, Wei61]. Photography [MP61]. Physical [BV69, Cru69, HW67, Lew69, SS62a, Wes64]. Physiological [Wei61]. Pick [Ano64a]. Pick-Nicollet [Ano64a]. Pickerel [Mil62]. Pigment [PMD67]. Pigmentation [BA65]. Pigments [ZE61]. Pike [BW68, BC65a, Eco60, FS60, FS63, Hin61, Hun65b, JM69, McC60, McC61a, McC62, Sch62]. Pimephales [Bru69, MS66, MT68]. Pink [Cum61, GF66, Hel66, Hir66, II60, SE60, SH62a, SN68, Tho62]. Pisces [RM66]. Place [FCKP60]. Plains [MK69]. Plankton [Cow65, JM63, MB62, ZE61]. Planktonic [Cus63]. Planted [BM64, Fra68, FB69]. Planting [CWSM60]. Plasma [Gre68, Rol68]. Plastic [ER60, Ver69]. Plate [HD62]. Plates [PYS68]. platostomus [Pri63a]. platyrhynchus [Hau69]. Poecilia [Dav68, SC69]. Point [OS63]. Pollock [AW68, HJ62]. Pollution [KM64, PH69]. polyphemus [Rop61]. Polyvinylpyrrolidone [Eco63]. Polyvinylpyrrolidone-Iodine [Eco63]. Pomatomus [DCM66]. Pomoxis [Han65]. Pompano [Fin69b]. Pond [AC60, Bri61, BKA63, CHA63, Hav60, Hav61, LGS^{+}62, WB69]. Pondfishes [Swi60]. Ponds [BL60, Cow65, Gas62, HG61, IB63, JM63, MB62, Tur60, Yas63]. Pools [Lew69]. Population [Beh69b, Cha67a, Col66, Coo66, Dav64, Dov68, EW69, Fis65, HD67, IM62, KM64, LSL63, MSS62, MC69, Mom67, Smi64, SS66, Wal69, Wat60]. Populations [Bjo61, CJ68, Coo67a, Cow65, Els68, Etn68, Faj62, Hat61, HCR69, HC60, Joh65, JM63, KDO63, Lew69, LFB68, MC62, MCA62, NW63, Reg62a, Rob62, Roe65, Sch68a, Sch67, Sea66, SA62, SA66, Sil69, WL63]. portions [AS67]. Portland [Ano66a]. Possible [BF68a, Dov68]. Post [Hos67]. Post-larval [Hos67]. Postlarvae [RM66]. Postlarval [CDF68, Sie68]. Potassium [CJ60, Fri65]. Potential [CP68, MR62, PC66, Rup61, Ryd65, Sch65a]. Potentially [Roe66]. praegustator [Dah69, TR67]. Pre [BHH66]. Pre-impoundment [BHH66]. Precaution [Har67]. Precocious [Bon68, TEB69]. Predation [Dov68, Gal67, HR68, HW67, LW62, Mom67, Pat62, Roo60, Sti67, TT67, Tho62]. Predator [CP68, PC66]. Predicting [DM65b]. Preference [Hil68].

Preparation [SP66b]. Preparing [Pow64b]. Present [Dav69, Sha69a]. Pressure [FS65]. Prevent [Dox63]. Prickly [McL68]. Priest [Bjo61, Bjo61]. Primary [McC63a, RD65]. Prince [Ree67a]. Principle [HG64]. Problem [McD68]. Problems [Abr63, Car63, FW62b]. Procambarus [MO63]. Procedures [Wat60]. Product [SMB67]. Production [Ben60, Cha65, CI68, CHA63, EW69, FW64, JM69, Man68, Mei62, Pat68a, Rot61, Ryd65, SS62a, SP61, Wat62, Yas63]. Productive [HA64]. Productivity [McC63a, Mur62, RD65]. Products [Bap66, BH61]. Progeny [Wal63]. Program [Cha66, MC69, PB67a, Rid60, VR67]. Programs [Abr63, PKW63, SS66]. Project [Boy61]. Prolonged [Moe67]. promelas [Bru69, MS66, MT68]. Propagation [Hin61, SH62b]. Properties [MK68]. Property [Chr69]. Proposed [Bea69, Sai62]. Propriety [Ste69]. Prosopium [Nor69a]. Protein [Gri66]. Proteins [BB65, HS67, Rol68, Sum66]. Protogynous [MS64]. Protozoa [BK66]. pseudoharengus [Mar69, Nor67, Rot63]. Pseudopleuronectes [BSH65]. Ptychocheilus [MR69, PR69, VK69]. Public [BC65b]. Publications [Ano60c, Ano60d, Ano60e, Ano60f, Ano61b, Ano61c, Ano61d, Ano61e, Ano62e, Ano62f, Ano62g, Ano62h, Ano63c, Ano63d, Ano64b, Ano64c, Ano65c, Ano65d, Ano65e, Ano65f, Ano66b, Ano66c, Ano66d, Ano67d, Ano67e, Ano67f, Ano67g, Ano68a, Ano68b, Ano68c]. Puerto [Erd62, Erd68]. Puget [All67, All69]. Pugheaded [GV68]. Pugheadedness [Che65, Isa65]. Pulpwood [MS66]. Pulse [BB65]. Pumping [Man64]. Pumpkinseed [KW68]. punctatus [Alb69, BF67]. Pup [CJ68]. Pupfish [MI67]. Purse [Gre67]. Pyloric [App66].

Quabbin [MM60]. Quality [Bur69b, Rup61, Sch65b, Ste66]. Quantitative [Dav68, PBK69]. Quantity [HA64]. Quinte [SP69a].

Rachycentron [Ric67]. radiatus [Dah69]. Radioactive [Hos67]. Radioactivity [CJ60]. Radiography [BH66]. Radioisotopes [HPS61]. Radioisotopic [Kev66]. Radiothyroidectomy [Nor69b]. Radius [CDF68]. Rafinesque [Bru69, DM65a, GFR63, LP60, Pri66]. Rainbow [AYM69a, AYM69b, AH63, Bar69, BM66, BB66, Gal67, HAS69, HK67, Hun69, Jen69, KOB69, Kni63, KS65, MB65, MR62, NB65, NS60, Ree67a, RYW63, She67, SG67, SH69]. Rainy [CS67, SK63]. Raising [SH67]. Randomized [CJ68]. Range [McC61a]. Rapid [Bor63, Del68, Sch68b]. Rate [AB66a, HJ62, Hun60, Kev66, KW68, Pat68b, Pri69a, Ray69, Tar64, Wat62]. Rates [EW69, HH67, Laa69, Poo69, Ray68, RC61, SM64]. Ratio [Bur69a, Erd68, MCA62]. Ratios [All69, Lam63, Sch67]. Rays [Bul60]. Reaction [Ree69]. Reactions [BH68, WWD60]. Rear [RP69]. Reared [And62, Bro61c, Bue61, Hav60, TT67, WWC63, War62, Rei63]. Rearing [Eip63, Fre67]. Recapture [RR64]. Recaptured [Pow65]. Recaptures [Moo60]. Reciprocal [Nel68b]. Recognition [HV68]. Recolonization [Wat64a]. Record [PYB63]. Recorder [Mef68]. Records [Dah69, Daw67b, Hod66, Van61]. Recoveries [All65, Bue61, RR66, She62]. Recovery [Fra68, Hav60, MC69, McC63b, Tau63]. Recreational [HM69, Ste66]. Recruitment [DM65b]. Red [Daw63, MO63, SK64, SP61]. Redds [Cob61, OPA68]. Redear [Ell69, GG65, Lam63]. Redfish [BCC68]. Redhorse [Mey62]. Reef [Moe67]. Reference [Boy61, FS63, MB65, OW63, Sto69]. Refuge [HB64b, PB67b]. Region [Dry66, DB68]. Regression [Bon68, Fas67, PG65]. Regular [Hat61]. Regulated [JK69]. Regulated-Temperature [JK69]. Regulations [SA65, SA66]. Regurgitation [Dox63]. Related [HA64, KS60, Ryd64, Sai62, SP61]. Relation [Dah63, DLS66, Els68, Gri66, HW67, MM60, Pod61, Ray68, Sch65b, SK63, Wal63, WI64]. Relations [Tag69]. Relationship [BSR69, BC65a, Bur69a, Gre67, GFR63, Isa64, Lat65, Per65, PH69, RS63]. Relationships [BK66, CS64, Daw62, Daw63, Daw65, Lan63, LFB68, Lux69, Swi64a, Swi64b, WH61]. Relative [She60, Whi69b]. Release [FLM66, Pau62, Pow65]. Released [Kle66, Pow64a]. Releases [Hub68]. relicta [Sch66]. Removal [BB67, CHB63, Chu63, Joh65]. Renfro [CDF68]. Repopulation [GB69]. Reports [Ano68d, Ano69b]. Representation [Dav64]. Reproduction [BHD^{+}64, Nel68c, Pur61, Tar69]. Reproductive [Etn68, PR69, Sak66b, Sie68, SM66]. Requirements [HAS69, Kem68, MS61, SWD63, WI64]. Rereading [Car61]. Research [De 69, Mee61, Pau69, RPS68, TYII69, Ver69]. Reservoir [AM67, BHH66, CHB63, FR69, FH68, HCR69, HT64a, Hou65, MM60, MAR68, HD67, MMPS69, Ray69]. Reservoirs [KB69]. Residual [JM63]. Resistance [FB66]. Resistant [FLM66, LFB68]. Resources [GH60, Sok69]. Response [Ghe66, Sto69, Tar64]. Responses [HB64a]. Restoration [Hav61]. Restoring [HC60]. Restriction [SA62]. Resulting [Gas62]. Results [LSL63, SA65]. Retention [SH69]. reticulata [Dav68, SC69]. Reticulate [PC66]. Retrieval [OS62a]. Return [WWC63]. Returns [And62]. reversed [CI68]. Review [Joh68a]. Reviews [Ano60g, Ano60h, Ano60i, Ano60j, Ano61f, Ano61g, Ano61h, Ano61i, Ano62i, Ano62j, Ano62k, Ano62l, Ano63e, Ano63f, Ano63g, Ano63h, Ano64d, Ano64e, Ano64f, Ano64g, Ano65g, Ano65h, Ano65i, Ano65j, Ano66e, Ano66f, Ano66g, Ano66h, Ano67h, Ano67i, Ano67j, Ano67k, Ano68e, Ano68f, Ano68g]. Revolving [Zur63]. Rhinichthys [Hou64, Tar69]. Rhode [OG68, BSH65, Coo65b, Coo66, Coo67b]. Richardson [Low65, Ree67a, SG67, WWC63]. Rico [Erd62, Erd68]. Rights [Chr69]. Ripe [Mil62]. River [Beh69a, Beh69b, BM64, CS67, MP60, May69, MS62, Mey62, MK69, Moo60, MLRW68, PH69, RT67, She60, SA66, SK63, Sta62, SM69, WWC63, War62, WF63, WI69, Wel67, AS67, BR68, But65, Car66, Cha67a, Cop61, Cus63, Fre68, Hel69, Hoo60, Hub61, Hub63, IT68, KM64, McC67, Min62, MMPS69, Par69, Pur63, Ray69, Ric62, Sch65c, Sch68c, SP69b, VK69, Wal63]. Rivers [Ray68]. Robson [Pau62]. robusta [VK69]. Roccus [BC63b, LB66, RT67]. Rockfish [Hit62]. Rockfishes [HD65]. Role [HW67]. Rollers [Har62]. Romania [RVIS68]. Ronge [Qad61]. Roosevelt [Joh68b]. Rotenone [BC63a, BH61, CM69b, HC60, KDO63]. Rotenoning [PC67]. Rough [CHB63]. round [Mra64, Nor69a]. Routine [Bea64]. Royale [Rah65]. Run [MGC64, Tar69].

S [HD65, MZC66, Clu64]. S. [Jen69]. Sable [AS67, SA66]. Sac [TBG69]. Sacramento [Cha67a]. Sagehen [She68a]. Saginaw [Gor61]. salar [AE66, IW69, PH69, WH61, War63, WI69]. Salinity [Lew66, LH68, Tag69]. Salmo [AE66, BM66, BB66, HM67, HK67, IW69, Jen69, KOB69, Low65, MC62, MPS67, PH69, Ree67a, SG67, Sto69, Swi63, UPH65, UD66, WWC63, Wag68, Wag69b, WH61, War63, WI69]. salmoides [Clu64, AM67, Clu64, Kel62, KS60, Lau69]. Salmon [All65, All67, All69, AYM69a, AE66, AW68, BCC68, Bla61, Bla63, Bon68, Bur69b, Cha65, Cro64, DFWD63, Dav67, DM61, DB64, EB60, ES61, FR69, Fre65, GF66, Gro60, GJ69, Gun67, HAS69, HSH62, Hav60, HM69, Hea65, Hel66, HWD62, Hir66, IW69, Laa69, Lac69, Lan63, LP66, Lin63, Mee61, Mei62, MMPS69, Mur68, Nig69, Nor69b, Par69, Pat62, PH69, Ray68, Ray69, Ree67b, Rob67, Roo60, RPR60, SPJ61, SE60, SH62a, She60, SWD64, SWD63, SN68, Tar64, Tar66, TBG69, TEB69, TT67, Tho62, UPH65, WV63, WH61, War62, War63, WF63, WI69, Wes64, WYHW60]. Salmonid [CP68, PC66, WWD60]. Salmonidae [SN68]. Salmonids [BR68, CGE62, GN65, Kat61, PMD67]. Salmons [LW62]. salmositica [BK65, BK66]. Salt [Gro60]. saltatrix [DCM66]. Salvelinus [AB66b, Bul60, Cho64, Coo67a, EW65, FW62b, FW64, FB69, Hal68, HH69, LW62, McC63b, RVIS68, Ree67a, SS62b, Sch65b, Vin60, WLWY60, Web62]. Sample [Rid60, RR64, SH67]. Sampler [HD62, SL68, WK61]. Sampling [CJ68, Fis65, FW62b, JW62]. San [Cha67a, RT67]. Sand [SM69]. Sandhill [McC60]. sandvicensis [MNB65]. sapidissima [GV68, Wea65]. sapidus [Tag69]. Saprolegnia [BM66]. Sarda [Bos66]. Saskatchewan [Qad61]. Saturation [Spe69]. Sauger [Nel68b, Nel68c, Pri63b, Sch65c]. saxatilis [LB66, RT67]. Scalation [MGC64]. Scale [Cha67b, Coo65a, Cou64, FS60, Pow64b, Pri64, Pri66, Reg62b, Rot63, Sie65, WH61]. Scales [Car61, CW66, FCKP60, Kee65, Mar69, SP69a]. Scandinavia [Nil63]. Scarring [Hub61]. Scattergram [Fas67]. Scent [Gre62]. Schedules [PD67]. Schooling [Beh69b, Whi69b]. Science [Pat69]. Sciences [OS62a, TYII69]. SCUBA [PKW63]. Sculpin [Mac66, McL68, PC66]. Sculpins [She68a]. Sea [Dav67, HS64, HR68, Joh62, Man67, Man68, Rah67, Sch65a, Sta62, WS62, Lar64]. Seal [CJ68]. Seas [Fre65]. Seasonal [Cla68, HH67, LS60, Par69, RW63, Wel66, WT65]. Seasons [CHB63, Hit62, She62]. Seatrout [IM62, SS62c]. Seaward [Pow64a, WWC63]. Seawater [Low64]. Sebago [AE66]. Sebastodes [Hit62, HD65]. Second [Ano63b]. Sectioning [Sch68b]. Segments [GB69].

Seining [Gre67]. Selected [BC60, Joh68a, Smi67]. Selection [Sie68]. Selective [DM61, Kon69, Mey65, Gal67]. Selectively [MR69]. Selectivity [AC60, Ber61, BW68, FR63, Hea62, MF60, McC61b]. Selene [OG68]. Semiannual [Lat63]. Semicircular [CDF68]. Seminal [Gre68]. Sensitivity [CGZ68, HCB66]. Separated [PYS68]. Separating [App66]. Separation [Sie69]. September [Ano60a, Ano61a, Ano62c, Ano63b, Ano64a, Ano65b, Ano66a, Ano67c]. Series [ML60]. Serum [Boo65, BB65, Cro64, Gri66, HS67, LED67, Sum66]. Session [Chr69]. Setting [Sha69b]. Several [Joh61, SC66, Wei61]. Sevin(R) [SMB67]. Sex [Bai64, Bro61b, Bur69a, CI68, Erd68, Lam63, MCA62, Sch67]. Sex-reversed [CI68]. Sexes [Fli69]. Sexual [BF68a, Kel62, Lew62, MS64]. Shad [Che65, GV68, Hod66, HD67, KB69, Lam65, Mil67, Ric62, Wal61, Wal63, Wea65]. Shallow [JM69]. Sheepshead [Han69, HCB66]. Shelf [Gra68]. Shellfish [Dav69]. Shellfisheries [Glu66]. Shelter [Kha64]. Shiner [Fuc67, SK64, Sum64, Sum66, SM69, WL63]. Shiners [LFB68]. Shipboard [RPS68]. Shoals [HD67, KB69]. Shocker [LM61]. Shore [Rup65]. Shore-Spawning [Rup65]. Shoreline [CDF68]. Short [SLN68]. Short-Term [SLN68]. Shortnose [Pri63a]. Shovelnose [Hel69]. Shrimp [BSH69, CDF68, Chi60, CM69a, Cum61, HR68, HW69, II60, VC62, ZEK65]. Shrinkage [Lux60]. sicculus [MAR68, Nel68a]. Sight [Gre62]. Significance [Mar66, PBK69]. Silica [PYS68]. Siltation [SP66a]. Silver [JF60]. Silverside [MAR68, Nel68a]. Silvex [Cow65]. Simcoe [McC63b]. Simplified [BF68b]. Simulated [Har65]. Simulation [Pau69, Sil69, Wal69]. Single [GW67, Pau62]. Single- [Pau62]. Sioux [Wel67]. Siscowets [EP65]. Six [FS65, JM63, Lux69, MC62, Pow65]. sixth [Ano67c]. Size [BSR69, Bon68, Dov68, Gal67, HD67, Lam65, RR64]. Size-selective [Gal67]. Sized [Von61, LSL63, She62]. Skipjack [Mag69, NY61, Nak65, Sho64, Yue66]. Slowed [Der63]. Sludge [CS67]. Small [Bla63, BC63a, CHA63, HG61, Hea62, Kem68, Lac68, Man67, McC63a, PBK69, RD65, SS62b, Wat60]. Small-Meshed [Hea62]. Smallmouth [Faj62, LD68, McC67]. Smears [Pow64b]. Smelt [Bai64, Bur69a, FR63, Gor61, Lac69, MM60, Rot61, Rup65]. Smith [Col66]. smithi [TR67]. Smoky [LP60, MK69, SM69]. Snake [Ray68, MMPS69]. Snapper [Cro62a]. Snappers [Daw63]. Snow [PC67]. Society [Ano60a, Ano61a, Ano62a, Ano62c, Ano63a, Ano63b, Ano64a, Ano65b, Ano66a, Ano67c, RB60]. Sockeye [AYM69a, BCC68, Bla61, GF66, Gro60, HSH62, Hea65, Lan63, Roo60, Tar64, Tar66, TT67, WV63]. Sodium [Cow65, Gil66, NS60]. Soil [Dah63]. Soils [JM63]. Solenoid [SL68]. Solids [RT67]. Solomons [CS64, SD63]. Solutions [DLS66, Joy61]. Some [AYM69b, ASV66, BW68, BH61, Cru69, Daw65, Eis65, Gil66, HH69, Hel69, Hil67, Hoo60, Kat61, KC61, KS60, Lar64, LM63, LFB68, LS60, MT68, MRC64, MK68, OW63, Ree69, Sch68c, SM64, Sok69, SMB67, SC69, Swa69, Tag69, Tan60, Tar69, VAE68, Vin60]. Sonar [HH67]. Sound [All67, All69, Cla68, PBK69]. Sounds [Sto69]. Sources [Alb69, BST61].

South [Fuc67, Poo69, Swe68, Ber61, Bud60, Daw62, FW62a]. Southeastern [LW62, Ree69, Tho62, Wel66]. Southwest [GH60]. Spawning [Cum61, EG65, Erd68, Hal68, HH69, IT68, Lam65, Man64, McL68, MR62, OS62b, OS63, RT67, Ree67b, Rob67, Rup65, War63, Web62, Wes64]. Spearfishing [PKW63]. Special [FS63, OW63, SA65, SA66]. Species [AK62, BL60, Dah69, Das61, FS65, Goo67, JG64a, JG64b, Kat61, MM60, Mey62, MS61, PHL62, SC66, SH62b, Wei61]. Specimens [MH68]. Spectra [Sto69]. Speed [Boy61, MNB65]. Speeds [Yue66]. Sperm [Cru69]. Spermatozoan [Fri65]. Sphaerotilus [SK63]. Sphyraena [SM61]. Spill [LP66]. Spill-Pattern [LP66]. Spillway [LHK68]. Spines [Sch68b]. Spiny [Kha64]. Splake [Bro61a]. Spleens [Gro68]. Sponge [Kha64]. Sport [De 69, Erd62, GH60, Kem68, MRC64, McF69, Ste66, Ste69, Wal60]. Spot [HP65, Low64]. Spottail [SK64]. Spotted [App66, BT62, IM62, SP69b, SS62c]. spp [BHH66, CDF68, Hit62, LW62]. Spraying [Cop61]. Spring [And62, Hil68, Min62]. Spring-Stocked [And62]. Spruce [Cop61]. Square [HG64]. Squawfish [MR69, PR69, TT67, VK69]. St [SM66]. St. [Moo60]. Stability [Faj62]. Stage [She61]. Stages [DCM66, Mar67]. Stains [WV63, ZEK65]. Stamina [Gre64, HK67]. Standard [Bea64, Daw62]. Standing [IB63, RD65, Tur60]. Start [Bon68]. Starvation [Bea64]. State [BC65b]. State-owned [BC65b]. States [Gra68, PKW63, Ree69]. Status [BGHK61, PKW63, Sha69a, Smi64, WI69]. Steelhead [BJ61, Cob61, Laa69, MPS67, Nor69b, OPA68, SWD64, SWD63, UPH65, UD66, WWC63, Wag68, Wag69b]. Stem [MLRW68]. Stepwise [SH67]. Stickleback [Kat61]. Stimulation [KTW61]. Stizostedion [EG65, Hou67, Nel68c]. Stocked [And62, Chu63, She62, Von61]. Stocking [CWSM60, LSL63, MR62, Wag68, Wag69b]. Stocks [FW62b, Vin60]. Stoeltzner’s [Kee65]. Stomach [Dox63, JF60, Jer69]. Stomachs [KM65]. Stonefly [CP68, JG64a, JG64b]. Stoneroller [LP60]. Storing [Wat65]. Strains [Coo61, FLM66, FW64, MBD67]. stramineus [SM69]. Stratified [Hil65]. Stream [BK66, Bro61c, BB67, CM69b, CO63, Eip63, Els68, EW69, Faj62, Har65, Hun65b, Jen69, Lew69, Mac66, Nav62, NW63, Pat62, PC67, SS62a, SS62b, SA65, She68b, WP60, WK61, Wat62, Wat64a]. Streams [BC60, BC63a, Cha65, GB69, Joh65, Low65, Man68, MBD67, OPA68, She61, She62, SP69b]. Strength [FS63]. Streptomycin [CGZ68]. Stress [Hun69]. Strigeid [Hou64]. Strigeoidea [HP65]. Striped [Bro65, Cha67a, CS61, Cla68, CS64, Dov68, Lew62, LB66, RT67, Tal67]. Structure [Bjo61, Nav62]. Studies [BSH65, Bor63, Cum63, HHC66, HGC67, Kon69, Mef68, MNB65, ST63, Swa69, Wal69, WYHW60]. Study [AB66b, Das61, GN68, HB64b, MS62, Poo69]. Studying [HT64a, Lac68, PC67]. Sturgeon [Hel69, She60]. Styrylpyridine [HT64b]. Subjected [Coo67a]. Sublethal [GW67, Gil66, Low64]. Subpopulations [Fen64]. Subsequent [MZC66]. Substances [Ree69]. Subtropical [Clu64]. Success [BC65b, Etn68, Glu66, Gre67, HK67, Ste66, War63, WCH62]. Successful [Sch66]. Succession [MS64]. Sucker [Hau69]. Suckers [OS63, PH69]. Sugar [LNCT66]. Suitable [RPS68]. Sulfamerazine [ES60]. Sulfate [WCH62]. Sulfide [BF67, HS64]. Sulfonate [LM63, ST63]. Sulphur [Dah63]. Summarization [PB67a]. Summary [Chr69]. Summer [CCF67, Hel69, Hun65a, Lat62]. Sunfish [Ell69, Etn68, GG65, GFR63, KW68, Lam63]. Sunfishes [SH62b]. Superior [Bai64, Bue61, DET65, Dry66, DB68, Eds60, EP65, Man68, Rah65, Rah67, Rah68, SE60, SH62a]. Supersaturation [Har67]. Supplies [BH61]. Supply [Laa69]. Surface [LNCT66]. Surplus [HM69]. Survey [BHH66, MRC64, RH69]. Surveying [BC63a]. Surveys [SH67]. Survival [BCC68, BT62, BB67, Bud60, Chu63, Cob61, CS67, Eip63, FW64, FS65, HSH62, Joh61, Lew65, Lew66, LH68, MBD67, Moe67, Pau62, Rei63, Rot61, Rup65, Sch65b, SP66a, She61, She67, SK63, SK65, SH69, Wag68]. Survivals [Bur69b]. Susceptibility [AYM69b, SG67, Sti67]. Susceptible [FLM66, LFB68]. Suspended [KS65]. Sweden [Lin63]. Swimming [Boy61, DFWD63, Gro60, HH67, LD68, MS66, MNB65, TBG69, Wea65, Yue66]. Symposium [McF69, Car63]. Synchronous [Aca68]. Synthetic [Eis65]. System [BR68, BM64, Cop61, Mou61, Mou64a, OS62a, OMK67, SS62b]. Systematic [JW62]. Systems [Pat69, Van69].

T. [Fin69b]. Tag [Del68, ER60, Gra64, HV68, RR66]. Tagged [Bud60, Bue61, Hub63, JW61, McC63b, Moe67, Pow65, RR66, SH69, Sti67, Wol63]. Tagging [CS61, Der63, Fis65, Jen67, Pau62, SM61]. Tags [AB66b, Jon66, She67]. Tail [Pea62]. Tail-less [Pea62]. Taken [BST61, Roe65]. Taku [MS62]. Taneycomo [FH68]. Tanner [Per65]. Tautog [BF68a, Coo66, Coo67b]. Tautoga [BF68a, Coo66, Coo67b]. Teaching [Pau69]. Tebay [NB65]. Technique [Arn66, GN65, LS61, MP61, Mou64b, Pow64b, Wat65, WP67]. Techniques [BH66, RH69]. Teleosts [BK65, Swa69]. Tellurite [Fri65]. temperate [Ryd65]. Temperature [AN64, JK69, LD68, Lew65, Lew66, LH68, Pod61, Rei63, Tag69]. Temperatures [DFWD63, HG61, Joh62, Str61]. Temporary [KM64]. Ten [JG64a, MO63]. Tendency [OS63]. Tennessee [Ano62c, CHB63]. TEPA [SC69]. tergisus [Van61]. Term [GN68, JG64b, SLN68, SKO66]. Testing [Sea66]. Testosterone [CI68]. Teton [Ano63b]. Tetracycline [Cho64]. Texas [Chi60, Ren63, ZE61]. TFM [AK62]. Their [Hea65, Het68, Hun65a, Nel68b, SH62b, HW67, Wat64b]. Theoretical [Pat68a]. Thermal [GN65, GJ69, Nig69, Tag69]. Thermal-Marking [GN65]. Thermally [Hil65]. Thermally-Stratified [Hil65]. Thermocline [Gre67]. Third [SH62a, Ano64a]. Thread [RP69]. Threadfin [HD67, KB69, Lam65, Mil67]. Three [BL60, Car66, CHB63, Cha65, Dah69, Das61, Daw67b, Daw69, Etn68, GC62, Goo67, Kat61, Low65, Mey62, Moe67, MS61, RD65, SH62b, Vin60]. Threespine [Kat61]. Threshold [Tar66]. Thunnus [Bea69]. Thyroid [GFR63, RVIS68]. Tidal [Boy61, CW65]. Tilapia [Sho64, AN64, CI68, SF63]. Tilapias [Swi60]. Tilesius [Gra64, Pow64a]. Tillamook [Sum62]. Time [FCKP60, Wag68]. Timsan [AYM69b]. Tissue [EE66, OW63]. Toadfish [SD63]. Tolerance [AN64, Tag69]. Tool [JK69]. Tools [Jon66]. Tortugas [II60]. Total [Daw62, Gri66, RT67]. Tow [CDF68]. Toxaphene [Hil65, KCN62, Low64]. Toxicant [Fin69a, MS67, PB67b]. Toxicants [BF68b, Kon69]. Toxicities [SC66]. Toxicity [AYM69b, AK62, Bru69, DLS66, Kat61, KC61, MO63, PHL62, SLN68, SMB67, WB69]. Toxicology [HGC67]. Toxins [SG67]. Trachinotus [Fin69b]. Transfer [Hna69]. Transformation [Dah63]. Transformed [Dav67]. Translocation [MZC66]. Translucent [MP61]. Transmitter [HHC66]. Transplantation [SS62b]. Trap [Pat68b]. Trappers [DM65b]. Trapping [Mur68]. Trawl [CDF68, FR63, HD67, MD68, PB67a, Roe65]. Treated [JM63]. Treatment [BM67, CM69b, Eco63, Gas62]. Trematoda [HP65]. Trematode [Hou64, NY61]. Trends [De 69, Gra68, Ket68, Log68, McF69]. Tributaries [Sha69b]. Tributary [Man67]. Tricaine [Bla61, EB60]. Trichiurus [Daw67a]. Trifluormethyl [AK62]. trifluoromethyl [Smi67]. Trinectes [Daw62]. Tripping [SL68]. Troll [All67]. Trophic [HA64, PBK69]. Trout [AC60, All63, AYM69a, AYM69b, And62, AH63, BV69, Bar69, BM66, Ben60, BC60, BM64, BB66, Bro61a, BB67, Bud60, Bue61, Bul60, BS69, BHD^{+}64, BH68, CCF67, Cho64, Cob61, Coo61, Coo67a, Der63, DM65b, DET65, ES60, Eco63, Eip63, Els68, EW69, EP65, EW65, FG66, FW62b, FW64, Fra68, FB69, Gal67, Gre64, Gre68, Hal68, HH69, HAS69, Hat61, HK67, Hun69, HB64b, Hun65a, Hun65b, Jen69, Kle66, KOB69, Kni63, KS65, Laa69, Lac69, LS61, LM61, Lat62, Lat63, Lat65, Lew69, Log63, Low65, Mac66, Mar66, MBD67, McC63b, MB65, MC62, MCA62, Mom65, Mom67, MPS67, MR62, NB65, NS60, Nil63, Nor69b, OPA68, PJS^{+}66, Pod61, PC67, RVIS68, Rah65, Rah67, Rah68, Ree67a, Rei63, RYW63, SS62a, SS62b, Sch65b, She60, She61, SA62]. Trout [She62, SA65, SA66, She67, She68b, SWD64, SWD63, SG67, SH69, Sto69, Sum62, Swi63, Tan60, Tau63, UPH65, UD66, Vin60, Von61, WWC63, Wag68, Wag69b, WP60, WF63, Wat60, WLWY60, Web62, WKH63, Win66]. Trout-Sculpin [Mac66]. trutta [HM67, Jen69, MC62, Swi63]. tshawytscha [Nig69]. Tuamotu [Nak65]. Tube [Spe69]. Tubificid [CGZ68]. Tuna [BJ67, CJ60, Gre67, Mag69, NY61, Nak65, Sho64, Whi69a, Yue66]. Tunnels [Bla63]. Turbidity [Mur62]. Turbine [BR68, BR69]. Turbines [CO64]. Two [BA65, GB69, Hod66, JG64a, JG64b, Jon66, Kon69, KC68, Man68, Pat68a, PG65, Reg62a, SC66, Sha69b, She61, She68b, Swi60, VAE68, WV63, WI69, WKH63, Faj62]. Type [ER60]. Types [For61, Joh61]. tyrannus [Coo65b, Lew65, Lew66, LH68, MH68].

U [Spe69]. U-Tube [Spe69]. Ultrasonic [HHC66, Mef68]. umpquae [MR69]. Underwater [HV68, NW63, Sto69]. Uniformity [ES60]. Unisexual [CI68]. United [Gra68, Ree69, PKW63]. UNIVAC [Abr65, Cha66]. University [TYII69]. Unspawned [FG66]. Untagged [Sti67]. Untrained [Han69]. Unusual [Daw69, Jen69]. Unusually [Coo65b]. upon [MLRW68, Pat62, She67, Vin60]. Upper [Bjo61, But65, Car66, Fre68, Hub61, Hub63, KM64, Par69, Sch65c]. Upstream [LP66]. Uptake [Bap66, FLM66, ZEK65]. Urinary [Tal67]. Urine [Hun69]. Usage [Har67]. Use [AB66b, Arn66, BC63a, Bul60, Cho64, Dox63, Fis65, FCKP60, Har62, HM69, Hea62, HHC66, HC60, HPS61, HT64a, Jer69, LM61, Mee61, Mey65, Pau62, PYS68, PKW63, RPS68]. Used [Man64, RP69]. Usefulness [Ste69]. Using [Cum63, PJS^{+}66, PB67b]. Utah [Gra61]. Utilization [Hea65, Hoo60, Sok69]. Uvulifer [HP65].

Validation [Reg62b]. Validity [Ste69]. Values [Ste66, Ste69]. Varden [Aca68, AB66a, AB66b, BA65, LW62, ND65, Ree67a, TT67]. Variability [Roe65]. Variables [Bar69, PG65]. Variation [Alb69, McL68]. Variations [Car61]. Various [DFWD63, Str61]. Vegetation [Gas62]. Velocities [SWD63]. Ventral [BB67]. vermiculatus [McC60, Sch62]. Vernier [Woo69]. Versenate [FCKP60]. Vertebrae [MB65]. Vertebral [GF66, MH68]. Vertical [FR69, HT64a, Lac68]. Viability [SN68]. Views [Ste69]. Viosca [Ano62d]. Virgin [Erd68]. Virginia [MP60]. virginica [Sak66a, Sak66b]. Virtual [Col66]. Virus [AYM69a, RPR60]. Virus-like [RPR60]. Visceral [Fre65]. vitreum [Hou67, EG65, Hou67]. vomer [OG68]. Vulnerability [Chu63, HW67, LGLB61, LH64].

Wabash [SP69b]. Walbaum [AB66b, BSH65, Bul60, LB66, Ree67a, SE60, SH62a, SM61]. Wales [Ree67a]. Walleye [AW68, CWSM60, Car61, Chu63, CS67, EG65, Hou67, Jer69, Joh61, Man64, Nel68b, Pri63b, Pri64, Pyc61, Sch65c, SP61, SK63, SK65]. Walleyes [CW61, Cro62b, For61, For63, OS62b, Pri69b, SKO66, Wol63]. Warm [BL60, Bro61c, VAE68]. Warm-Water [BL60, Bro61c, VAE68]. Warmouth [Hun60]. Warmwater [Arn66, Joh65, LSL63, PHL62, SM64]. Washington [BK66, Joh62, Pat62, Rob67, TT67, UD66]. Water [BST61, BH61, BL60, Bro61c, BF68b, Cob61, Dox63, FH68, Gro60, Har67, HA64, MZC66, Mou61, Mou64a, Mur62, Pod61, Rei63, Sch65b, SWD64, SWD63, Spe69, SL68, WYHW60, Zur63, VAE68]. Waters [Hit62, Hod66, Joh61, LNCT66, Mar69, MT68, NY61, Ric67, RH68]. Weed [CO63]. Weight [Bon68, BC65a, Bur69a, CS64, Dav68, Daw62, Daw63, Daw65, Isa64, Lux69, Rob62, Swi64a, Swi64b]. Weighted [PG65]. Weights [Erd68]. Wenatchee [TT67]. West [Mom65, Sak66a, Sak66b]. Western [Tar69, Bai64, BK66, Man64, Roe66, UD66]. White [Gre62, Han65, Isa64, OS63, SP69a, Sie65, Sie68, Sie69]. Whitefish [Boo65, BC62, Eds60, Fen64, MF60, McC61b, Mra64, Nor69a, Qad61]. Wickford [OG68]. Width [Per65, RS63]. Wild [BM64, BB67, Coo61, FW62b, FW64, Gre64, Hun65a, MBD67, MCA62, She68b, WYHW60]. Wilson [Mar69, Nor67, Rot63]. Wing [MSS62]. Winnebago [Pri63b, Pri67, Pri69a, Pri69b]. Winnibigoshish [Joh61]. Winter [BSH65, Log63, OWW69, Poo69, PC67, Sai62, Rei63]. Winterkill [JM69]. Wisconsin [BB67, Hun65b, Pri63b, Pri67, Pri69a, Pri69b]. within [SS62b].

Wood [KS65, SK63]. Wyoming [Ano63b, AC60, MR62].

X [Hec69]. Xanthophylls [PJS^{+}66]. xanthurus [Low64].

Year [CWSM60, FW64, For61, FS63, KS60, KS62, Lat65, Pyc61, She62, HB64b]. Year-Class [For61]. Yearling [Ray68]. Yearlings [Wol63]. Years [BC65b, Pow65, Tau63]. Yellow [BC63b, FB66, Gal67, HH67, RNBH60]. Yellowfin [Het68, TR67, Yue66]. Yellowstone [Ben60, Cop61, Sto69]. Yellowtail [Lux60]. Yield [PG64, Wal63]. Yields [Hir66, Pat68a]. Yolk [HD65, Lau69, TBG69]. York [Cla68, For63, Hou67, MP60, Poo69, WLWY60]. Young [AM67, AW68, BT62, Daw63, Gam63, Hou67, Kha64, LW62, Lat65, LH68, MP60, PMD67, Pri69b, Rid60, Roo60, Wat64b]. Young-of-the-Year [Lat65].

Zealand [McD68]. Zinc [Bru69, Hos64, Joy61, Mou64b]. Zinc-65 [Hos64]. Zinc-caused [Mou64b]. Zones [Els68]. Zooplankton [KDO63].

References

Armstrong:1966:DRD

[AB66a] Robert H. Armstrong and Roger F. Blackett. Digestion rate of the Dolly Varden. Transactions of the American Fisheries Society, 95(4):429–430, October 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Armstrong:1966:UED

[AB66b] Robert H. Armstrong and Roger F. Blackett. Use and evaluation of dart tags to study the migration habits of Dolly Varden, Salvelinus malma (Walbaum). Transactions of the American Fisheries Society, 95(3):320–323, July 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Abramson:1963:CPF

[Abr63] Norman J. Abramson. Computer programs for fisheries problems. Transactions of the American Fisheries Society, 92(3):310, July 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Abramson:1965:BGC

[Abr65] Norman J. Abramson. Von Bertalanffy growth curve II, IBM 7094, UNIVAC 1107, Fortran IV. Transactions of the American Fisheries Society, 94(2):195–196, April 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Allen:1960:SFB

[AC60] George H. Allen and Leighton G. Claussen. Selectivity of food by brook trout in a Wyoming beaver pond. Transactions of the American Fisheries Society, 89(1):80–81, January 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Acara:1968:NSH

[Aca68] A. H. Acara. A normal, synchronous hermaphrodite Dolly Varden char. Transactions of the American Fisheries Society, 97(2):201– 202, April 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ames:1966:MII

[ACP66] Warren E. Ames, R. M. Condie, and B. Pollara. A method for intravascular injection and bleeding of fishes. Transactions of the American Fisheries Society, 95(3):317–318, July 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anderson:1966:CDL

[AE66] Richard B. Anderson and W. Harry Everhart. Concentrations of DDT in landlocked salmon (Salmo salar) at Sebago Lake, Maine. Transactions of the American Fisheries Society, 95(2):160–164, April 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ashley:1963:MMR

[AH63] Laurence M. Ashley and John E. Halver. Multiple metastasis of rainbow trout hepatoma. Transactions of the American Fisheries Society, 92(4):365–371, October 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anderson:1969:AFI

[AH69] Richard O. Anderson and M. LeRoy Heman. Angling as a factor influencing catchability of largemouth bass. Transactions of the American Fisheries Society, 98(2):317–320, April 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Applegate:1962:CTT

[AK62] Vernon C. Applegate and Everett L. King, Jr. Comparative toxicity of 3-trifluormethyl-4-nitrophenol (TFM) to larval lampreys and eleven species of fishes. Transactions of the American Fisheries Society, 91(4):342–345, October 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Albaugh:1969:SGV

[Alb69] Douglas W. Albaugh. Sources of growth variation among individual black bullheads, Ictalurus melas, and channel catfish, Ictalurus punctatus. Transactions of the American Fisheries Society, 98(1): 35–44, January 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Allison:1963:CHL

[All63] Leonard N. Allison. Cataract in hatchery lake trout. Transactions of the American Fisheries Society, 92(1):34–38, January 1963. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Allen:1965:EEA

[All65] George H. Allen. Estimating error associated with ocean recoveries of fin-marked Coho salmon. Transactions of the American Fisheries Society, 94(4):319–326, October 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Allen:1966:DAD

[All66a] K. Radway Allen. Determination of age distribution from age– length keys and length distributions, IBM 7090, 7094, Fortran IV. Transactions of the American Fisheries Society, 95(2):230–231, April 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Allen:1966:FBG

[All66b] K. Radway Allen. Fitting of von Bertalanffy growth-curves, IBM 709, 7094, Fortran IV. Transactions of the American Fisheries Society, 95(2):231–232, April 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Allen:1967:CPS

[All67] George H. Allen. Contribution of Puget Sound Coho salmon, Oncorhynchus kisutch, to the 1953 Pacific Coast commercial troll fishery. Transactions of the American Fisheries Society, 96(1):42–54, January 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Allen:1969:CER

[All69] George H. Allen. Catch-to-escapement ratios of fin-marked 1950brood Puget Sound Coho salmon. Transactions of the American Fisheries Society, 98(4):599–610, October 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Alt:1965:FHI

[Alt65] Kenneth T. Alt. Food habits of inconnu in Alaska. Transactions of the American Fisheries Society, 94(3):272–274, July 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Applegate:1967:FYL

[AM67] Richard L. Applegate and James W. Mullan. Food of young largemouth bass, Micropterus salmoides, in a new and old reservoir. Transactions of the American Fisheries Society, 96(1):74–77, January 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Allanson:1964:TTM

[AN64] B. R. Allanson and R. G. Noble. The tolerance of Tilapia mossambica (Peters) to high temperature. Transactions of the American Fisheries Society, 93(4):323–332, October 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anderson:1962:CRF

[And62] Richard B. Anderson. A comparison of returns from fall- and spring-stocked hatchery-reared lake trout in Maine. Transactions of the American Fisheries Society, 91(4):425–427, October 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1960:ENA

[Ano60a] Anonymous. Eighty-Ninth Annual Meeting of the American Fisheries Society: Clearwater, Florida September 16–18, 1959. Transactions of the American Fisheries Society, 89(1):87–132, January 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1960:M

[Ano60b] Anonymous. In memoriam. Transactions of the American Fisheries Society, 89(1):133, January 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1960:RPa

[Ano60c] Anonymous. Recent publications. Transactions of the American Fisheries Society, 89(1):86, January 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1960:RPb

[Ano60d] Anonymous. Recent publications. Transactions of the American Fisheries Society, 89(2):246, April 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1960:RPc

[Ano60e] Anonymous. Recent publications. Transactions of the American Fisheries Society, 89(3):318–319, July 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1960:RPd

[Ano60f] Anonymous. Recent publications. Transactions of the American Fisheries Society, 89(4):385–386, October 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1960:Ra

[Ano60g] Anonymous. Reviews. Transactions of the American Fisheries Society, 89(1):84–86, January 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1960:Rb

[Ano60h] Anonymous. Reviews. Transactions of the American Fisheries Society, 89(2):243–246, April 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1960:Rc

[Ano60i] Anonymous. Reviews. Transactions of the American Fisheries Society, 89(3):315–318, July 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1960:Rd

[Ano60j] Anonymous. Reviews. Transactions of the American Fisheries Society, 89(4):383–385, October 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1961:NAM

[Ano61a] Anonymous. Ninetieth Annual Meeting of the American Fisheries Society: Denver, Colorado: September 12–14, 1960. Transactions of the American Fisheries Society, 90(1):91–120, January 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1961:RPa

[Ano61b] Anonymous. Recent publications. Transactions of the American Fisheries Society, 90(1):88–89, January 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1961:RPb

[Ano61c] Anonymous. Recent publications. Transactions of the American Fisheries Society, 90(2):235, April 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1961:RPc

[Ano61d] Anonymous. Recent publications. Transactions of the American Fisheries Society, 90(3):342–343, July 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1961:RPd

[Ano61e] Anonymous. Recent publications. Transactions of the American Fisheries Society, 90(4):503–504, October 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1961:Ra

[Ano61f] Anonymous. Reviews. Transactions of the American Fisheries Society, 90(1):86–88, January 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1961:Rb

[Ano61g] Anonymous. Reviews. Transactions of the American Fisheries Society, 90(2):232–235, April 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1961:Rc

[Ano61h] Anonymous. Reviews. Transactions of the American Fisheries Society, 90(3):341–342, July 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1961:Rd

[Ano61i] Anonymous. Reviews. Transactions of the American Fisheries Society, 90(4):502–503, October 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1962:CAF

[Ano62a] Anonymous. Committees: American fisheries society 1961–62. Transactions of the American Fisheries Society, 91(2):241, April 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1962:M

[Ano62b] Anonymous. In memoriam. Transactions of the American Fisheries Society, 91(4):433, October 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1962:NFA

[Ano62c] Anonymous. Ninety-First Annual Meeting of the American Fisheries Society: Memphis, Tennessee: September 13–15, 1961. Transactions of the American Fisheries Society, 91(1):103–128, January 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1962:PVJ

[Ano62d] Anonymous. Percy Viosca, Jr. 1892–1961. Transactions of the American Fisheries Society, 91(1):132, January 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1962:RPa

[Ano62e] Anonymous. Recent publications. Transactions of the American Fisheries Society, 91(1):130–131, January 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1962:RPb

[Ano62f] Anonymous. Recent publications. Transactions of the American Fisheries Society, 91(2):240, April 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1962:RPc

[Ano62g] Anonymous. Recent publications. Transactions of the American Fisheries Society, 91(3):331, July 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1962:RPd

[Ano62h] Anonymous. Recent publications. Transactions of the American Fisheries Society, 91(4):432, October 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1962:Ra

[Ano62i] Anonymous. Reviews. Transactions of the American Fisheries Society, 91(1):129–130, January 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1962:Rb

[Ano62j] Anonymous. Reviews. Transactions of the American Fisheries Society, 91(2):237–240, April 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1962:Rc

[Ano62k] Anonymous. Reviews. Transactions of the American Fisheries Society, 91(3):330–331, July 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1962:Rd

[Ano62l] Anonymous. Reviews. Transactions of the American Fisheries Society, 91(4):430–432, October 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1963:CAF

[Ano63a] Anonymous. Committees: American fisheries society 1962–63. Transactions of the American Fisheries Society, 92(2):190–191, April 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1963:NSA

[Ano63b] Anonymous. Ninety-Second Annual Meeting of the American Fisheries Society: Grand Teton National Park, Jackson, Wyoming: September 9–12, 1962. Transactions of the American Fisheries Society, 92(1):71–89, January 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1963:RPa

[Ano63c] Anonymous. Recent publications. Transactions of the American Fisheries Society, 92(1):70, January 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1963:RPb

[Ano63d] Anonymous. Recent publications. Transactions of the American Fisheries Society, 92(2):189, April 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1963:Ra

[Ano63e] Anonymous. Reviews. Transactions of the American Fisheries Society, 92(1):69–70, January 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1963:Rb

[Ano63f] Anonymous. Reviews. Transactions of the American Fisheries Society, 92(2):187–189, April 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1963:Rc

[Ano63g] Anonymous. Reviews. Transactions of the American Fisheries Society, 92(3):323–326, July 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1963:Rd

[Ano63h] Anonymous. Reviews. Transactions of the American Fisheries Society, 92(4):440, October 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1964:NTA

[Ano64a] Anonymous. Ninety-third Annual Meeting of the American Fisheries Society: Pick-Nicollet Hotel, Minneapolis, Minnesota, September 11–13, 1963. Transactions of the American Fisheries Society, 93(1):109–125, January 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1964:RPa

[Ano64b] Anonymous. Recent publications. Transactions of the American Fisheries Society, 93(3):321–322, July 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1964:RPb

[Ano64c] Anonymous. Recent publications. Transactions of the American Fisheries Society, 93(4):404, October 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1964:Ra

[Ano64d] Anonymous. Reviews. Transactions of the American Fisheries Society, 93(1):108, January 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1964:Rb

[Ano64e] Anonymous. Reviews. Transactions of the American Fisheries Society, 93(2):213, April 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1964:Rc

[Ano64f] Anonymous. Reviews. Transactions of the American Fisheries Society, 93(3):320–321, July 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1964:Rd

[Ano64g] Anonymous. Reviews. Transactions of the American Fisheries Society, 93(4):402–404, October 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1965:M

[Ano65a] Anonymous. In memoriam. Transactions of the American Fisheries Society, 94(4):416, October 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1965:NFA

[Ano65b] Anonymous. Ninety-fourth Annual Meeting of the American Fisheries Society: Chalfonte-Haddon Hotel, Atlantic City, New Jersey, 13–15 September 1964. Transactions of the American Fisheries Society, 94(1):105–122, January 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1965:RPa

[Ano65c] Anonymous. Recent publications. Transactions of the American Fisheries Society, 94(1):103, January 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1965:RPb

[Ano65d] Anonymous. Recent publications. Transactions of the American Fisheries Society, 94(2):201, April 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1965:RPc

[Ano65e] Anonymous. Recent publications. Transactions of the American Fisheries Society, 94(3):285, July 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1965:RPd

[Ano65f] Anonymous. Recent publications. Transactions of the American Fisheries Society, 94(4):415–416, October 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1965:Ra

[Ano65g] Anonymous. Reviews. Transactions of the American Fisheries Society, 94(1):99–103, January 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1965:Rb

[Ano65h] Anonymous. Reviews. Transactions of the American Fisheries Society, 94(2):197–200, April 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1965:Rc

[Ano65i] Anonymous. Reviews. Transactions of the American Fisheries Society, 94(3):281–285, July 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1965:Rd

[Ano65j] Anonymous. Reviews. Transactions of the American Fisheries Society, 94(4):413–415, October 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1966:NFA

[Ano66a] Anonymous. Ninety-fifth Annual Meeting of the American Fisheries Society: Hilton Hotel, Portland, Oregon, 22–24 September 1965. Transactions of the American Fisheries Society, 95(1):119– 144, January 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1966:RPa

[Ano66b] Anonymous. Recent publications. Transactions of the American Fisheries Society, 95(1):118, January 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1966:RPb

[Ano66c] Anonymous. Recent publications. Transactions of the American Fisheries Society, 95(2):236–237, April 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1966:RPc

[Ano66d] Anonymous. Recent publications. Transactions of the American Fisheries Society, 95(4):446–447, October 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1966:Ra

[Ano66e] Anonymous. Reviews. Transactions of the American Fisheries Society, 95(1):116–118, January 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1966:Rb

[Ano66f] Anonymous. Reviews. Transactions of the American Fisheries Society, 95(2):234–236, April 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1966:Rc

[Ano66g] Anonymous. Reviews. Transactions of the American Fisheries Society, 95(3):332–334, July 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1966:Rd

[Ano66h] Anonymous. Reviews. Transactions of the American Fisheries Society, 95(4):442–446, October 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1967:EC

[Ano67a] Anonymous. Editorial comment. Transactions of the American Fisheries Society, 96(3):372, July 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1967:E

[Ano67b] Anonymous. Errata. Transactions of the American Fisheries Society, 96(4):435, October 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1967:NSA

[Ano67c] Anonymous. Ninety-sixth Annual Meeting of the American Fisheries Society: Muehlebach Motel, Kansas City, Missouri, 11–14 September 1966. Transactions of the American Fisheries Society, 96(1):85–115, January 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1967:RPa

[Ano67d] Anonymous. Recent publications. Transactions of the American Fisheries Society, 96(1):83, January 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1967:RPb

[Ano67e] Anonymous. Recent publications. Transactions of the American Fisheries Society, 96(2):235–236, April 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1967:RPc

[Ano67f] Anonymous. Recent publications. Transactions of the American Fisheries Society, 96(3):371–372, July 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1967:RPd

[Ano67g] Anonymous. Recent publications. Transactions of the American Fisheries Society, 96(4):435, October 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1967:Ra

[Ano67h] Anonymous. Reviews. Transactions of the American Fisheries Society, 96(1):78–83, January 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1967:Rb

[Ano67i] Anonymous. Reviews. Transactions of the American Fisheries Society, 96(2):231–235, April 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1967:Rc

[Ano67j] Anonymous. Reviews. Transactions of the American Fisheries Society, 96(3):367–371, July 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1967:Rd

[Ano67k] Anonymous. Reviews. Transactions of the American Fisheries Society, 96(4):431–435, October 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1968:RPa

[Ano68a] Anonymous. Recent publications. Transactions of the American Fisheries Society, 97(2):217–218, April 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1968:RPb

[Ano68b] Anonymous. Recent publications. Transactions of the American Fisheries Society, 97(3):310–311, July 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1968:RPc

[Ano68c] Anonymous. Recent publications. Transactions of the American Fisheries Society, 97(4):514, October 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1968:Ra

[Ano68d] Anonymous. Reports. Transactions of the American Fisheries Society, 97(1):93–108, January 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1968:Rb

[Ano68e] Anonymous. Reviews. Transactions of the American Fisheries Society, 97(2):212–217, April 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1968:Rc

[Ano68f] Anonymous. Reviews. Transactions of the American Fisheries Society, 97(3):306–310, July 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1968:Rd

[Ano68g] Anonymous. Reviews. Transactions of the American Fisheries Society, 97(4):508–514, October 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1969:M

[Ano69a] Anonymous. In memoriam. Transactions of the American Fisheries Society, 98(4):727, October 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Anonymous:1969:R

[Ano69b] Anonymous. Reports. Transactions of the American Fisheries Society, 98(1):170–192, January 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Applegate:1966:PCC

[App66] Richard L. Applegate. Pyloric caeca counts as a method for separating the advanced fry and fingerlings of largemouth and spotted basses. Transactions of the American Fisheries Society, 95(2):226, April 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Arnold:1966:UJI

[Arn66] Dean E. Arnold. Use of the jaw-injection technique for marking warmwater fish. Transactions of the American Fisheries Society, 95(4):432–433, October 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Alexander:1967:FBP

[AS67] Gaylord R. Alexander and David S. Shetter. Fishing and boating on portions of the Au Sable River in Michigan, 1960–63. Transactions of the American Fisheries Society, 96(3):257–267, July 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Andrews:1966:SEH

[ASV66] Austin K. Andrews, Bruce E. Stebbings, and Charles C. Van Valin. Some effects of heptachlor on bluegills (Lepomis macrochirus). Transactions of the American Fisheries Society, 95(3):297–309, July 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Armstrong:1968:IWP

[AW68] Robert H. Armstrong and Peter C. Winslow. An incidence of walleye Pollock feeding on salmon young. Transactions of the American Fisheries Society, 97(2):202–203, April 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Amend:1969:HVD

[AYM69a] Donald F. Amend, William T. Yasutake, and Robert W. Mead. A hematopoietic virus disease of rainbow trout and sockeye salmon. Transactions of the American Fisheries Society, 98(4):796–804, October 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Amend:1969:SFI

[AYM69b] Donald F. Amend, William T. Yasutake, and Reginald Morgan. Some factors influencing susceptibility of rainbow trout to the acute toxicity of an ethyl mercury phosphate formulation (Timsan). Transactions of the American Fisheries Society, 98(3):419– 425, July 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Blackett:1965:CTA

[BA65] Roger F. Blackett and Robert H. Armstrong. Collection of two abnormal Dolly Varden: One with two dorsal fins, the other with incomplete pigmentation. Transactions of the American Fisheries Society, 94(4):408–409, October 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bailey:1964:AGM

[Bai64] Merryll M. Bailey. Age, growth, maturity, and sex composition of the American smelt, Osmerus mordax (Mitchill), of Western Lake Superior. Transactions of the American Fisheries Society, 93(4): 382–395, October 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baptist:1966:UMF

[Bap66] John P. Baptist. Uptake of mixed fission products by marine fishes. Transactions of the American Fisheries Society, 95(2):145–152, April 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Barnhart:1969:ECV

[Bar69] Roger A. Barnhart. Effects of certain variables on hematological characteristics of rainbow trout. Transactions of the American Fisheries Society, 98(3):411–418, July 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bouck:1965:IDO

[BB65] Gerald R. Bouck and Robert C. Ball. Influence of a diurnal oxygen pulse on fish serum proteins. Transactions of the American Fisheries Society, 94(4):363–370, October 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bouck:1966:ICM

[BB66] Gerald R. Bouck and Robert C. Ball. Influence of capture methods on blood characteristics and mortality in the rainbow trout (Salmo gairdneri). Transactions of the American Fisheries Society, 95(2): 170–176, April 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brynildson:1967:EPV

[BB67] Oscar M. Brynildson and Clifford L. Brynildson. The effect of pectoral and ventral fin removal on survival and growth of wild brown trout in a Wisconsin stream. Transactions of the American Fisheries Society, 96(3):353–355, July 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beyerle:1960:GBT

[BC60] George B. Beyerle and Edwin L. Cooper. Growth of brown trout in selected Pennsylvania streams. Transactions of the American Fisheries Society, 89(3):255–262, July 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Budd:1962:ECL

[BC62] John C. Budd and D. Cucin. Exploitation of Canadian Lake Huron whitefish. Transactions of the American Fisheries Society, 91(2): 223–224, April 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boccardy:1963:URE

[BC63a] Joseph A. Boccardy and Edwin L. Cooper. The use of rotenone and electrofishing in surveying small streams. Transactions of the American Fisheries Society, 92(3):307–310, July 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Buchholz:1963:FYB

[BC63b] Marvin M. Buchholz and Kenneth D. Carlander. Failure of yellow bass, Roccus mississippiensis, to form annuli. Transactions of the American Fisheries Society, 92(4):384–390, October 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:1965:LWR

[BC65a] Edward H. Brown, Jr. and Clarence F. Clark. Length–weight relationship of northern pike, Esox lucius, from East Harbor, Ohio. Transactions of the American Fisheries Society, 94(4):404–405, October 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Byrd:1965:FYM

[BC65b] I. B. Byrd and Johnie H. Crance. Fourteen years of management and fishing success in Alabama’s state-owned public fishing lakes. Transactions of the American Fisheries Society, 94(2):129– 134, April 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boyar:1967:ADC

[BC67] H. C. Boyar and R. A. Clifford. An automatic device for counting dry fish eggs. Transactions of the American Fisheries Society, 96 (3):361–363, July 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bjornn:1968:MSR

[BCC68] T. C. Bjornn, D. R. Craddock, and D. R. Corley. Migration and survival of Redfish Lake, Idaho, sockeye salmon, Oncorhynchus nerka. Transactions of the American Fisheries Society, 97(4):360– 373, October 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beamish:1964:ISS

[Bea64] F. W. H. Beamish. Influence of starvation on standard and routine oxygen consumption. Transactions of the American Fisheries Society, 93(1):103–107, January 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beardsley:1969:PMA

[Bea69] Grant L. Beardsley, Jr. Proposed migrations of albacore, Thunnus alalunga, in the Atlantic Ocean. Transactions of the American Fisheries Society, 98(4):589–598, October 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beeton:1961:ECL

[Bee61] Alfred M. Beeton. Environmental changes in Lake Erie. Transactions of the American Fisheries Society, 90(2):153–159, April 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Behmer:1969:MEF

[Beh69a] David J. Behmer. A method of estimating fecundity; with data on river carpsuckers, Carpiodes carpio. Transactions of the American Fisheries Society, 98(3):523–524, July 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Behmer:1969:SRC

[Beh69b] David J. Behmer. Schooling of river carpsuckers and a population estimate. Transactions of the American Fisheries Society, 98(3): 520–523, July 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Benson:1960:FIP

[Ben60] Norman G. Benson. Factors influencing production of immature cutthroat trout in Arnica Creek, Yellowstone Park. Transactions of the American Fisheries Society, 89(2):168–175, April 1960. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Berst:1961:SEE

[Ber61] A. H. Berst. Selectivity and efficiency of experimental gill nets in South Bay and Georgian Bay of Lake Huron. Transactions of the American Fisheries Society, 90(4):413–418, October 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bonn:1967:EHS

[BF67] Edward W. Bonn and Billy J. Follis. Effects of hydrogen sulfide on channel catfish, Ictalurus punctatus. Transactions of the American Fisheries Society, 96(1):31–36, January 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bridges:1968:SDT

[BF68a] David W. Bridges and Michael P. Fahay. Sexual dichromatism in the tautog, Tautoga onitis (Linnaeus), with an observation of possible courtship behavior. Transactions of the American Fisheries Society, 97(2):208–209, April 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Burke:1968:SFT

[BF68b] W. David Burke and Denzel E. Ferguson. A simplified flow-through apparatus for maintaining fixed concentrations of toxicants in water. Transactions of the American Fisheries Society, 97(4):498–501, October 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Benson:1961:SMN

[BGHK61] Norman G. Benson, John R. Greeley, Melvin I. Huish, and Jerome H. Kuehn. Status of management of natural lakes. Transactions of the American Fisheries Society, 90(2):218–224, April 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bonn:1961:SER

[BH61] Edward W. Bonn and Leon R. Holbert. Some effects of rotenone products on municipal water supplies. Transactions of the American Fisheries Society, 90(3):287–297, July 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bartlett:1966:TRF

[BH66] Martin R. Bartlett and Richard L. Haedrich. Techniques in the radiography of fishes. Transactions of the American Fisheries Society, 95(1):99–101, January 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Butler:1968:RDT

[BH68] Robert L. Butler and Vernon M. Hawthorne. The reactions of dominant trout to changes in overhead artificial cover. Transactions of the American Fisheries Society, 97(1):37–41, January 1968. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Burdick:1964:ADL

[BHD^{+}64] G. E. Burdick, E. J. Harris, H. J. Dean, T. M. Walker, Jack Skea, and David Colby. The accumulation of DDT in lake trout and the effect on reproduction. Transactions of the American Fisheries Society, 93(2):127–136, April 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Becker:1966:PIS

[BHH66] David A. Becker, Robert G. Heard, and Perry Don Holmes. A pre-impoundment survey of the helminth and copepod parasites of Micropterus spp. of Beaver Reservoir in Northwest Arkansas.

Transactions of the American Fisheries Society, 95(1):23–34, January 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bates:1961:LED

[BJ61] Daniel W. Bates and Stanley G. Jewett, Jr. Louver efficiency in deflecting downstream migrant steelhead. Transactions of the American Fisheries Society, 90(3):336–337, July 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bullis:1967:POF

[BJ67] Harvey R. Bullis, Jr. and Rolf Juhl. Phalanx orientation in feeding behavior of the little tuna, Euthynnus alletteratus. Transactions of the American Fisheries Society, 96(2):122–125, April 1967. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bjornn:1961:HAS

[Bjo61] Ted C. Bjornn. Harvest, age structure, and growth of game fish populations from Priest and Upper Priest Lakes. Transactions of the American Fisheries Society, 90(1):27–31, January 1961. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Becker:1965:IHC

[BK65] Clarence D. Becker and Max Katz. Infections of the hemoflagellate, Cryptobia salmositica Katz, 1951, in freshwater teleosts of the Pacific Coast. Transactions of the American Fisheries Society, 94 (4):327–333, October 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Becker:1966:HRC

[BK66] Clarence D. Becker and Max Katz. Host relationships of Cryptobia salmositica (Protozoa: Mastigophora) in a Western Washington hatchery stream. Transactions of the American Fisheries Society, 95(2):196–202, April 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bridges:1963:PDM

[BKA63] W. R. Bridges, B. J. Kallman, and A. K. Andrews. Persistence of DDT and its metabolites in a farm pond. Transactions of the American Fisheries Society, 92(4):421–427, October 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:1960:AGF

[BL60] C. J. D. Brown and Sidney M. Logan. Age and growth of four species of warm-water game fish from three Montana ponds. Transactions of the American Fisheries Society, 89(4):379–382, October 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Blahm:1961:ETM

[Bla61] T. H. Blahm. Effect of tricaine methanesulfonate on oxygen consumption of juvenile sockeye salmon. Transactions of the American Fisheries Society, 90(2):226–227, April 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Blahm:1963:PSF

[Bla63] Theodore H. Blahm. Passage of salmon fingerlings through small tunnels. Transactions of the American Fisheries Society, 92(3): 302–303, July 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bonde:1960:FHB

[BM60] Thomas Bonde and J. E. Maloney. Food habits of burbot. Transactions of the American Fisheries Society, 89(4):374–376, October 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bjornn:1964:MPW

[BM64] Ted C. Bjornn and Jerry Mallet. Movements of planted and wild trout in an Idaho river system. Transactions of the American Fisheries Society, 93(1):70–76, January 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Benoit:1966:CSE

[BM66] Raymond F. Benoit and Nathaniel A. Matlin. Control of Saprolegnia on eggs of rainbow trout (Salmo gairdneri) with ozone. Transactions of the American Fisheries Society, 95(4):430–432, October 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brungs:1967:DCT

[BM67] William A. Brungs and Donald I. Mount. A device for continuous treatment of fish in holding chambers. Transactions of the American Fisheries Society, 96(1):55–57, January 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bonham:1968:GIM

[Bon68] Kelshaw Bonham. Growth inhibition by maxillary excision, size at start of jaw elongation, and regression of weight on length of precocious male Chinook salmon. Transactions of the American Fisheries Society, 97(2):198–201, April 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Booke:1965:ISG

[Boo65] Henry E. Booke. Increase of serum globulin levels with age in lake whitefish, Coregonus clupeaformis. Transactions of the American Fisheries Society, 94(4):397–398, October 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Borgeson:1963:RMF

[Bor63] David P. Borgeson. A rapid method for food-habits studies. Transactions of the American Fisheries Society, 92(4):434–435, October 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boschung:1966:OCB

[Bos66] Herbert T. Boschung, Jr. The occurrence of the common bonito, Sarda sarda, in the Northern Gulf of Mexico. Transactions of the American Fisheries Society, 95(2):227–228, April 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Boyar:1961:SSI

[Boy61] H. C. Boyar. Swimming speed of immature Atlantic herring with reference to the Passamaquoddy Tidal Project. Transactions of the American Fisheries Society, 90(1):21–26, January 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bentley:1968:CJS

[BR68] Wallace W. Bentley and Howard L. Raymond. Collection of juvenile salmonids from turbine intake gatewells of major dams in the Columbia River system. Transactions of the American Fisheries Society, 97(2):124–126, April 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bentley:1969:PJF

[BR69] Wallace W. Bentley and Howard L. Raymond. Passage of juvenile fish through orifices in gatewells of turbine intakes at McNary Dam. Transactions of the American Fisheries Society, 98(4):723– 727, October 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bridges:1961:DEF

[Bri61] W. R. Bridges. Disappearance of endrin from fish and other materials of a pond environment. Transactions of the American Fisheries Society, 90(3):332–334, July 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:1961:BSB

[Bro61a] Bradford E. Brown. Behavior of splake and brook trout fingerlings. Transactions of the American Fisheries Society, 90(3):328– 329, July 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:1961:DSM

[Bro61b] Bradford E. Brown. Differential sex mortality associated with handling landlocked alewives. Transactions of the American Fisheries Society, 90(1):80–81, January 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:1961:MNH

[Bro61c] Edward H. Brown, Jr. Movement of native and hatchery-reared game fish in a warm-water stream. Transactions of the American Fisheries Society, 90(4):449–456, October 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:1965:MCS

[Bro65] Bradford E. Brown. Meristic counts of striped bass from Alabama. Transactions of the American Fisheries Society, 94(3):278–279, July 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brungs:1969:CTZ

[Bru69] William A. Brungs. Chronic toxicity of zinc to the fathead minnow, Pimephales promelas Rafinesque. Transactions of the American Fisheries Society, 98(2):272–279, April 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bullock:1969:BBK

[BS69] G. L. Bullock and S. F. Snieszko. Bacteria in blood and kidney of apparently healthy hatchery trout. Transactions of the American Fisheries Society, 98(2):268–271, April 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Berry:1965:GSW

[BSH65] Richard J. Berry, Saul B. Saila, and Donald B. Horton. Growth studies of winter flounder, Pseudopleuronectes americanus (Walbaum), in Rhode Island. Transactions of the American Fisheries Society, 94(3):259–264, July 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Benoit:1969:ALB

[BSH69] Duane Benoit, Roll Syrett, and John Hale. Automatic live brine shrimp feeder. Transactions of the American Fisheries Society, 98 (3):532–533, July 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baldwin:1969:RSB

[BSR69] Ruth E. Baldwin, Kay Gonnerman Sides, and John W. Robinson. Relationship of size and body area to the flavor and aroma in fish. Transactions of the American Fisheries Society, 98(3): 533–537, July 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baldwin:1961:FAF

[BST61] Ruth E. Baldwin, Dorothy H. Strong, and James H. Torrie. Flavor and aroma of fish taken from four fresh-water sources. Transactions of the American Fisheries Society, 90(2):175–180, April 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Brown:1962:GSY

[BT62] Bradford E. Brown and Walter M. Tatum. Growth and survival of young spotted bass in Lake Martin, Alabama. Transactions of the American Fisheries Society, 91(3):324–326, July 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Budd:1960:SGT

[Bud60] John C. Budd. Survival and growth of tagged lake trout in South Bay, Lake Huron. Transactions of the American Fisheries Society, 89(3):308–309, July 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Buettner:1961:RTH

[Bue61] Howard J. Buettner. Recoveries of tagged, hatchery-reared lake trout from Lake Superior. Transactions of the American Fisheries Society, 90(4):404–412, October 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Bulkley:1960:UBR

[Bul60] Ross V. Bulkley. Use of branchiostegal rays to determine age of lake trout, Salvelinus namaycush (Walbaum). Transactions of the American Fisheries Society, 89(4):344–350, October 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Burbidge:1969:AGL

[Bur69a] Richard G. Burbidge. Age, growth, length–weight relationship, sex ratio, and food habits of American smelt, Osmerus mordax (Mitchill), from Gull Lake, Michigan. Transactions of the American Fisheries Society, 98(4):631–640, October 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Burrows:1969:IFQ

[Bur69b] Roger E. Burrows. The influence of fingerling quality on adult salmon survivals. Transactions of the American Fisheries Society, 98(4):777–784, October 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Butler:1965:FDA

[But65] Robert L. Butler. Freshwater drum, Aplodinotus grunniens, in the navigational impoundments of the Upper Mississippi River. Transactions of the American Fisheries Society, 94(4):339–349, October 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Baldes:1969:PPM

[BV69] Richard J. Baldes and Robert E. Vincent. Physical parameters of microhabitats occupied by brown trout in an experimental flume. Transactions of the American Fisheries Society, 98(2):230–238, April 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Beyerle:1968:SOF

[BW68] George B. Beyerle and John E. Williams. Some observations of food selectivity by northern pike in aquaria. Transactions of the American Fisheries Society, 97(1):28–31, January 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carlander:1961:VRW

[Car61] Kenneth D. Carlander. Variations on rereading walleye scales. Transactions of the American Fisheries Society, 90(2):230–231, April 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carlander:1963:SFF

[Car63] Kenneth D. Carlander. Symposium on Freshwater Fishery Problems in ??? Transactions of the American Fisheries Society, 92 (3):257, July 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carlson:1966:ETO

[Car66] Clarence A. Carlson. Effects of three organophosphorus insecticides on immature Hexagenia and hydropsyche of the Upper Mississippi River. Transactions of the American Fisheries Society, 95(1):1–5, January 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chapman:1967:SDF

[CCF67] D. W. Chapman, H. J. Campbell, and J. D. Fortune, Jr. Summer distribution and food of Kokanee and trout in Elk Lake, Oregon. Transactions of the American Fisheries Society, 96(3):308– 312, July 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Caillouet:1968:ERD

[CDF68] Charles W. Caillouet, Jr., Ronald J. Dugas, and Bennie J. Fontenot, Jr. Effects of radius and direction of semicircular tow near the shoreline on catch of postlarval shrimp (Penaeus spp.) with the Renfro beam trawl. Transactions of the American Fisheries Society, 97(2):127–130, April 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Christmas:1960:DMG

[CG60] J. Y. Christmas and Gordon Gunter. Distribution of menhaden, genus Brevoortia, in the Gulf of Mexico. Transactions of the American Fisheries Society, 89(4):338–343, October 1960. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Collins:1962:ASA

[CGE62] Gerald B. Collins, Joseph R. Gauley, and Carl H. Elling. Ability of salmonids to ascend high fishways. Transactions of the American Fisheries Society, 91(1):1–7, January 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Coler:1968:TSS

[CGZ68] Robert A. Coler, Haim B. Gunner, and Bert M. Zuckerman. Tubificid sensitivity to streptomycin. Transactions of the American Fisheries Society, 97(4):502–503, October 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cooper:1963:GPL

[CHA63] Edwin L. Cooper, Herbert Hidu, and John K. Andersen. Growth and production of largemouth bass in a small pond. Transactions of the American Fisheries Society, 92(4):391–400, October 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chapman:1965:NPJ

[Cha65] D. W. Chapman. Net production of juvenile Coho salmon in three Oregon streams. Transactions of the American Fisheries Society, 94(1):40–52, January 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chadwick:1966:AGP

[Cha66] Harold K. Chadwick. Age and growth program, Univac 1107, Fortran IV. Transactions of the American Fisheries Society, 95(2): 229–230, April 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chadwick:1967:RMS

[Cha67a] Harold K. Chadwick. Recent migrations of the Sacramento–San Joaquin River striped bass population. Transactions of the American Fisheries Society, 96(3):327–342, July 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chapoton:1967:SDG

[Cha67b] Robert B. Chapoton. Scale development in the Gulf menhaden, Brevoortia patronus. Transactions of the American Fisheries Society, 96(1):60–62, January 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carroll:1963:TSR

[CHB63] Billy B. Carroll, Gordon E. Hall, and Robert D. Bishop. Three seasons of rough fish removal at Norris Reservoir, Tennessee. Transactions of the American Fisheries Society, 92(4):356–364, October 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cheek:1965:PAS

[Che65] Randall P. Cheek. Pugheadedness in an American shad. Transactions of the American Fisheries Society, 94(1):97–98, January 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chin:1960:BSF

[Chi60] Edward Chin. The bait shrimp fishery of Galveston Bay, Texas. Transactions of the American Fisheries Society, 89(2):135–141, April 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Choate:1964:UTD

[Cho64] Jerry Choate. Use of tetracycline drugs to mark advanced fry and fingerling brook trout (Salvelinus fontinalis). Transactions of the American Fisheries Society, 93(3):309–311, July 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Christy:1969:SSF

[Chr69] Francis T. Christy, Jr. Session summary: Fisheries goals and the rights of property. Transactions of the American Fisheries Society, 98(2):369–378, April 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Churchill:1963:EFR

[Chu63] Warren S. Churchill. The effect of fin removal on survival, growth, and vulnerability to capture of stocked walleye fingerlings. Transactions of the American Fisheries Society, 92(3):298–300, July 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clemens:1968:PUB

[CI68] Howard P. Clemens and Theophilus Inslee. The production of unisexual broods by Tilapia mossambica sex-reversed with methyl testosterone. Transactions of the American Fisheries Society, 97 (1):18–21, January 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chakravarti:1960:PIN

[CJ60] Diptiman Chakravarti and Timothy Joyner. Potassium as an index of naturally occurring radioactivity in tuna muscle. Transactions of the American Fisheries Society, 89(3):312–314, July 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chapman:1968:ESP

[CJ68] Douglas G. Chapman and Ancel M. Johnson. Estimation of fur seal pup populations by randomized sampling. Transactions of the American Fisheries Society, 97(3):264–270, July 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clark:1968:SMS

[Cla68] John Clark. Seasonal movements of striped bass contingents of long Island Sound and the New York bight. Transactions of the American Fisheries Society, 97(4):320–343, October 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clugston:1964:GFL

[Clu64] James P. Clugston. Growth of the Florida largemouth bass, Micropterus salmoides floridanus (LeSueur), and the northern largemouth bass, M. s. salmoides (Lacépède), in subtropical Florida. Transactions of the American Fisheries Society, 93(2):146–154, April 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cook:1969:CLP

[CM69a] Harry L. Cook and M. Alice Murphy. The culture of larval penaeid shrimp. Transactions of the American Fisheries Society, 98(4):751– 754, October 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cook:1969:ERT

[CM69b] S. F. Cook, Jr. and R. L. Moore. The effects of a rotenone treatment on the insect fauna of a California stream. Transactions of the American Fisheries Society, 98(3):539–544, July 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cushing:1963:EWB

[CO63] Colbert E. Cushing, Jr. and P. A. Olson. Effects of weed burning on stream conditions. Transactions of the American Fisheries Society, 92(3):303–305, July 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cramer:1964:PFT

[CO64] Frederick K. Cramer and Raymond C. Oligher. Passing fish through hydraulic turbines. Transactions of the American Fisheries Society, 93(3):243–259, July 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Coble:1961:IWE

[Cob61] Daniel W. Coble. Influence of water exchange and dissolved oxygen in redds on survival of steelhead trout embryos. Transactions of the American Fisheries Society, 90(4):469–474, October 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Coke:1968:DDF

[Cok68] Michael Coke. Depth distribution of fish on a bush-cleared area of Lake Kariba, Central Africa. Transactions of the American Fisheries Society, 97(4):460–465, October 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cole:1966:VPE

[Col66] Charles F. Cole. Virtual population estimations of largemouth bass in Lake Fort Smith, Arkansas, 1957–60. Transactions of the American Fisheries Society, 95(1):52–55, January 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cooper:1961:GWH

[Coo61] Edwin L. Cooper. Growth of wild and hatchery strains of brook trout. Transactions of the American Fisheries Society, 90(4):424– 438, October 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cooke:1965:HSP

[Coo65a] Peter Hayman Cooke. Head scale pattern as a characteristic for the identification of Fundulus heteroclitus and Fundulus majalis.

Transactions of the American Fisheries Society, 94(3):265–267, July 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cooper:1965:ULM

[Coo65b] Richard Arthur Cooper. An unusually large menhaden, Brevoortia tyrannus (Latrobe), from Rhode Island. Transactions of the American Fisheries Society, 94(4):412, October 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cooper:1966:MPE

[Coo66] Richard A. Cooper. Migration and population estimation of the tautog, Tautoga onitis (Linnaeus), from Rhode Island. Transactions of the American Fisheries Society, 95(3):239–247, July 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cooper:1967:GLB

[Coo67a] Edwin L. Cooper. Growth and longevity of brook trout (Salvelinus fontinalis) in populations subjected to light exploitation. Transactions of the American Fisheries Society, 96(4):383–386, October 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cooper:1967:AGT

[Coo67b] Richard A. Cooper. Age and growth of the tautog, Tautoga onitis (Linnaeus), from Rhode Island. Transactions of the American Fisheries Society, 96(2):134–142, April 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cope:1961:EDS

[Cop61] Oliver B. Cope. Effects of DDT spraying for spruce budworm on fish in the Yellowstone River system. Transactions of the American Fisheries Society, 90(3):239–251, July 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Courtright:1964:MMF

[Cou64] Robert C. Courtright. A method for making fish scale impressions in the field or laboratory. Transactions of the American Fisheries Society, 93(2):200–201, April 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cowell:1965:ESA

[Cow65] Bruce C. Cowell. The effects of sodium arsenite and Silvex on the plankton populations in farm ponds. Transactions of the American Fisheries Society, 94(4):371–377, October 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Claire:1968:SAP

[CP68] Errol W. Claire and Robert W. Phillips. The stonefly acroneuria pacifica as a potential predator on salmonid embryos. Transactions of the American Fisheries Society, 97(1):50–52, January 1968. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Croker:1962:GFG

[Cro62a] Robert A. Croker. Growth and food of the gray snapper, Lutjanus griseus in Everglades National Park. Transactions of the American Fisheries Society, 91(4):379–383, October 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Crowe:1962:HBW

[Cro62b] Walter R. Crowe. Homing behavior in walleyes. Transactions of the American Fisheries Society, 91(4):350–354, October 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Croston:1964:DAG

[Cro64] C. Bradford Croston. Differences among groups of salmon in the blood serum level of alkaline phosphatase. Transactions of the American Fisheries Society, 93(1):95–96, January 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cruea:1969:SCP

[Cru69] Darrell D. Cruea. Some chemical and physical characteristics of fish sperm. Transactions of the American Fisheries Society, 98(4): 785–788, October 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Chapoton:1961:ACM

[CS61] Robert B. Chapoton and James E. Sykes. Atlantic Coast migration of large striped bass as evidenced by fisheries and tagging. Transactions of the American Fisheries Society, 90(1):13–20, January 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Clemmer:1964:AGW

[CS64] Glenn H. Clemmer and Frank J. Schwartz. Age, growth, and weight relationships of the striped killifish, Fundulus majalis, near solomons, Maryland. Transactions of the American Fisheries Society, 93(2):197–198, April 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Colby:1967:SWE

[CS67] Peter J. Colby and Lloyd L. Smith, Jr. Survival of walleye eggs and fry on paper fiber sludge deposits in Rainy River, Minnesota. Transactions of the American Fisheries Society, 96(3):278–296, July 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cummings:1961:MSP

[Cum61] William C. Cummings. Maturation and spawning of the pink shrimp, Penaeus duorarum Burkenroad. Transactions of the American Fisheries Society, 90(4):462–468, October 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cummings:1963:UDE

[Cum63] William C. Cummings. Using the Doppler effect to detect movements of captive fish in behavior studies. Transactions of the American Fisheries Society, 92(2):178–180, April 1963. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cuplin:1969:PEC

[Cup69] Paul Cuplin. Performance evaluation of chelated minerals in Idaho open-formula diets. Transactions of the American Fisheries Society, 98(4):772–776, October 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Cushing:1963:FFI

[Cus63] Colbert E. Cushing, Jr. Filter-feeding insect distribution and planktonic food in the Montreal River. Transactions of the American Fisheries Society, 92(3):216–219, July 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carlander:1961:AGW

[CW61] Kenneth D. Carlander and Richard R. Whitney. Age and growth of walleyes in Clear Lake, Iowa, 1935–1957. Transactions of the American Fisheries Society, 90(2):130–138, April 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Croker:1965:KED

[CW65] Robert A. Croker and Alfred J. Wilson. Kinetics and effects of DDT in a tidal Marsh ditch. Transactions of the American Fisheries Society, 94(2):152–159, April 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Conley:1966:ODS

[CW66] Jerry M. Conley and Arthur Witt, Jr. The origin and development of scales in the flier, Centrarchus macropterus (Lacépède). Transactions of the American Fisheries Society, 95(4):433–434, October 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Carlander:1960:EWF

[CWSM60] Kenneth D. Carlander, Richard R. Whitney, Everett B. Speaker, and Kenneth Madden. Evaluation of walleye fry stocking in Clear Lake, Iowa, by alternate-year planting. Transactions of the American Fisheries Society, 89(3):249–254, July 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dahl:1963:TIS

[Dah63] Jorgen Dahl. Transformation of iron and sulphur compounds in soil, and its relation to Danish inland fisheries. Transactions of the American Fisheries Society, 92(3):260–264, July 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dahlberg:1966:LFA

[Dah66] Michael L. Dahlberg. Length-frequency analysis, IBM 7090/7094, Fortran II or Fortran IV. Transactions of the American Fisheries Society, 95(3):331, July 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dahlberg:1969:IIO

[Dah69] Michael D. Dahlberg. Incidence of the isopod, Olencira praegustator, and copepod, Lernaeenicus radiatus, in three species and hybrid menhaden (Brevoortia) from the Florida coasts, with five new host records. Transactions of the American Fisheries Society, 98(1):111–115, January 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Das:1961:CSB

[Das61] B. C. Das. Comparative study of the blood biochemistry of three species of Indian carp. Transactions of the American Fisheries Society, 90(1):1–5, January 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1964:GRC

[Dav64] William Spencer Davis. Graphic representation of confidence intervals for Petersen population estimates. Transactions of the American Fisheries Society, 93(3):227–232, July 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1967:PNT

[Dav67] Robert M. Davis. Parasitism by newly-transformed anadromous sea lampreys on landlocked salmon and other fishes in a coastal Maine lake. Transactions of the American Fisheries Society, 96(1): 11–16, January 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1968:QFW

[Dav68] Charles C. Davis. Quantitative feeding and weight changes in Poecilia reticulata. Transactions of the American Fisheries Society, 97(1):22–27, January 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1969:SHP

[Dav69] Harry C. Davis. Shellfish hatcheries — present and future. Transactions of the American Fisheries Society, 98(4):743–750, October 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dawson:1962:LWS

[Daw62] C. E. Dawson. Length — weight and standard length — total length relationships of South Carolina hogchokers, Trinectes maculatus. Transactions of the American Fisheries Society, 91(1):89–90, January 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dawson:1963:LWR

[Daw63] C. E. Dawson. Length and weight relationships of young red snappers from the Northern Gulf of Mexico. Transactions of the American Fisheries Society, 92(3):310–311, July 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dawson:1965:LWR

[Daw65] C. E. Dawson. Length–weight relationships of some Gulf of Mexico fishes. Transactions of the American Fisheries Society, 94(3):279– 280, July 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dawson:1967:CBC

[Daw67a] C. E. Dawson. Contributions to the biology of the cutlassfish (Trichiurus lepturus) in the Northern Gulf of Mexico. Transactions of the American Fisheries Society, 96(2):117–121, April 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dawson:1967:TNR

[Daw67b] C. E. Dawson. Three new records of partial albinism in American hetersomata. Transactions of the American Fisheries Society, 96 (4):400–404, October 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dawson:1968:CBM

[Daw68] C. E. Dawson. Contributions to the biology of the Mexican flounder, cyclopsetta chittendeni, in the Northern Gulf of Mexico. Transactions of the American Fisheries Society, 97(4):504–507, October 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dawson:1969:TUC

[Daw69] C. E. Dawson. Three unusual cases of abnormal coloration in Northern Gulf of Mexico flatfishes. Transactions of the American Fisheries Society, 98(1):106–108, January 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Donaldson:1964:ELL

[DB64] Lauren R. Donaldson and Kelshaw Bonham. Effects of low-level chronic irradiation of Chinook and Coho salmon eggs and alevins. Transactions of the American Fisheries Society, 93(4):333–341, October 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dryer:1968:GCB

[DB68] William R. Dryer and Joseph Beil. Growth changes of the bloater (Coregonus hoyi) of the Apostle Islands Region of Lake Superior. Transactions of the American Fisheries Society, 97(2):146–158, April 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Deuel:1966:DEE

[DCM66] David G. Deuel, John R. Clark, and A. J. Mansueti. Description of embryonic and early larval stages of bluefish, Pomatomus saltatrix. Transactions of the American Fisheries Society, 95(3): 264–271, July 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

DeSylva:1969:TMS

[De 69] Donald P. De Sylva. Trends in marine sport fisheries research. Transactions of the American Fisheries Society, 98(1):151–169, January 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dell:1968:NFT

[Del68] Michael B. Dell. A new fish tag and rapid, cartridge-fed applicator. Transactions of the American Fisheries Society, 97(1):57–59, January 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Deroche:1963:SGL

[Der63] Stuart E. Deroche. Slowed growth of lake trout following tagging. Transactions of the American Fisheries Society, 92(2):185– 186, April 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dryer:1965:FLT

[DET65] William R. Dryer, Leo F. Erkkila, and Clifford L. Tetzloff. Food of lake trout in Lake Superior. Transactions of the American Fisheries Society, 94(2):169–176, April 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1963:IOC

[DFWD63] Gerald E. Davis, Jack Foster, Charles E. Warren, and Peter Doudoroff. The influence of oxygen concentration on the swimming performance of juvenile Pacific salmon at various temperatures. Transactions of the American Fisheries Society, 92(2):111–124, April 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Davis:1969:LHE

[DL69] James R. Davis and Darrell E. Louder. Life history and ecology of Menidia extensa. Transactions of the American Fisheries Society, 98(3):466–472, July 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Doudoroff:1966:ATF

[DLS66] Peter Doudoroff, Gérard Leduc, and Carl R. Schneider. Acute toxicity to fish of solutions containing complex metal cyanides, in relation to concentrations of molecular hydrocyanic acid. Transactions of the American Fisheries Society, 95(1):6–22, January 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Donaldson:1961:SBC

[DM61] Lauren R. Donaldson and Deb Menasveta. Selective breeding of Chinook salmon. Transactions of the American Fisheries Society, 90(2):160–164, April 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Darnell:1962:DFC

[DM62] Rezneat M. Darnell and Richard R. Meierotto. Determination of feeding chronology in fishes. Transactions of the American Fisheries Society, 91(3):313–320, July 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Darnell:1965:DPB

[DM65a] Rezneat M. Darnell and Richard R. Meierotto. Diurnal periodicity in the black bullhead, Ictalurus melas (Rafinesque). Transactions of the American Fisheries Society, 94(1):1–8, January 1965. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Drummond:1965:PRC

[DM65b] Robert A. Drummond and Ted D. McKinney. Predicting the recruitment of cutthroat trout fry in Trappers Lake, Colorado. Transactions of the American Fisheries Society, 94(4):389–393, October 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dovel:1968:PSB

[Dov68] William L. Dovel. Predation by striped bass as a possible influence on population size of the Atlantic croaker. Transactions of the American Fisheries Society, 97(4):313–319, October 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Doxtater:1963:UIW

[Dox63] Gary D. Doxtater. Use of ice water to prevent regurgitation of stomach contents of fish. Transactions of the American Fisheries Society, 92(1):68, January 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Dryer:1966:BDF

[Dry66] William R. Dryer. Bathymetric distribution of fish in the Apostle Islands Region, Lake Superior. Transactions of the American Fisheries Society, 95(3):248–259, July 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eisler:1960:NCS

[EB60] Ronald Eisler and Tadeusz Backiel. Narcotization of Chinook salmon fingerlings with tricaine methanesulfonate (M.S. 222). Transactions of the American Fisheries Society, 89(2):164–167, April 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Economon:1960:FNP

[Eco60] Philip Economon. Furunculosis in northern pike. Transactions of the American Fisheries Society, 89(2):240–241, April 1960. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Economon:1963:ETI

[Eco63] Philip P. Economon. Experimental treatment of infectious pancreatic necrosis of brook trout with polyvinylpyrrolidone-iodine. Transactions of the American Fisheries Society, 92(2):180–182, April 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Editor:1963:EF

[Edi63] Kenneth D. Carlander Symposium Editor. Editor’s foreword. Transactions of the American Fisheries Society, 92(3):258, July 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Edsall:1960:AGW

[Eds60] Thomas A. Edsall. Age and growth of the whitefish, Coregonus clupeaformis, of Munising Bay, Lake Superior. Transactions of the American Fisheries Society, 89(4):323–332, October 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eisler:1966:EEB

[EE66] Ronald Eisler and Philip H. Edmunds. Effects of endrin on blood and tissue chemistry of a marine fish. Transactions of the American Fisheries Society, 95(2):153–159, April 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ellis:1965:SBW

[EG65] D. V. Ellis and M. A. Giles. The spawning behavior of the walleye, Stizostedion vitreum (Mitchill). Transactions of the American Fisheries Society, 94(4):358–362, October 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eipper:1963:EHR

[Eip63] Alfred W. Eipper. Effect of hatchery rearing conditions on stream survival of brown trout. Transactions of the American Fisheries Society, 92(2):132–139, April 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eisler:1965:SES

[Eis65] Ronald Eisler. Some effects of a synthetic detergent on estuarine fishes. Transactions of the American Fisheries Society, 94(1):26– 31, January 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eller:1969:PRS

[Ell69] Lafayette L. Eller. Pathology in redear sunfish exposed to hydrothol 191. Transactions of the American Fisheries Society, 98(1): 52–59, January 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Elser:1968:FPT

[Els68] Allen A. Elser. Fish populations of a trout stream in relation to major habitat zones and channel alterations. Transactions of the American Fisheries Society, 97(4):389–397, October 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Essbach:1969:CFH

[EM69] Alban R. Essbach and Riney Maxwell. A collapsable fish holding cage. Transactions of the American Fisheries Society, 98(2):337– 338, April 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eschmeyer:1965:FCF

[EP65] Paul H. Eschmeyer and Arthur M. Phillips, Jr. Fat content of the flesh of Siscowets and lake trout from Lake Superior. Transactions of the American Fisheries Society, 94(1):62–74, January 1965. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Everhart:1960:BTP

[ER60] W. Harry Everhart and Robert S. Rupp. Barb-type plastic fish tag. Transactions of the American Fisheries Society, 89(2):241– 242, April 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Erdman:1962:SFB

[Erd62] Donald S. Erdman. The sport fishery for blue marlin off Puerto Rico. Transactions of the American Fisheries Society, 91(2):225– 227, April 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Erdman:1968:SCS

[Erd68] Donald S. Erdman. Spawning cycle, sex ratio, and weights of blue marlin off Puerto Rico and the Virgin Islands. Transactions of the American Fisheries Society, 97(2):131–137, April 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Economon:1960:UDS

[ES60] Philip P. Economon and Robert E. Schumacher. Uniformity of dispersion of sulfamerazine in pelleted trout food. Transactions of the American Fisheries Society, 89(1):64–66, January 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Eisler:1961:DSL

[ES61] Ronald Eisler and Raymond C. Simon. Destruction of salmon larvae by Hydra oligactis. Transactions of the American Fisheries Society, 90(3):329–332, July 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Etnier:1968:RSN

[Etn68] David A. Etnier. Reproductive success of natural populations of hybrid sunfish in three Minnesota lakes. Transactions of the American Fisheries Society, 97(4):466–471, October 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Everhart:1965:LHB

[EW65] W. Harry Everhart and Charles A. Waters. Life history of the blueback trout (Arctic char, Salvelinus alpinus (Linnaeus)) in Maine. Transactions of the American Fisheries Society, 94(4):393–397, October 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Elwood:1969:EFF

[EW69] Jerry W. Elwood and Thomas F. Waters. Effects of floods on food consumption and production rates of a stream brook trout population. Transactions of the American Fisheries Society, 98(2): 253–262, April 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Faber:1968:NCL

[Fab68] Daniel J. Faber. A net for catching limnetic fry. Transactions of the American Fisheries Society, 97(1):61–63, January 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fajen:1962:ISS

[Faj62] Otto F. Fajen. The influence of stream stability on homing behavior of two smallmouth bass populations. Transactions of the American Fisheries Society, 91(4):346–349, October 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fast:1967:CRS

[Fas67] Arlo W. Fast. Correlation-regression with scattergram, IBM 1620, Fortran II-D. Transactions of the American Fisheries Society, 96 (2):230, April 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ferguson:1966:ERY

[FB66] Denzel E. Ferguson and C. Rex Bingham. Endrin resistance in the yellow bullhead, Ictalurus natalis. Transactions of the American Fisheries Society, 95(3):325–326, July 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fraser:1969:BLA

[FB69] J. M. Fraser and F. W. H. Beamish. Blood lactic acid concentrations in brook trout, Salvelinus fontinalis, planted by air drop. Transactions of the American Fisheries Society, 98(2):263–267, April 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fry:1960:ULV

[FCKP60] F. E. J. Fry, D. Cucin, J. C. Kennedy, and A. Papson. The use of lead versenate to place a time mark on fish scales. Transactions of the American Fisheries Society, 89(2):149–153, April 1960. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fenderson:1964:ESL

[Fen64] Owen C. Fenderson. Evidence of subpopulations of lake whitefish, Coregonus clupeaformis, involving a dwarfed form. Transactions of the American Fisheries Society, 93(1):77–94, January 1964. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fish:1966:AEC

[FG66] G. R. Fish and G. D. Ginnelly. An adverse effect of coelomic fluid on unspawned ova in trout. Transactions of the American Fisheries Society, 95(1):104–107, January 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fry:1968:LTC

[FH68] James P. Fry and Willis D. Hanson. Lake Taneycomo: a coldwater reservoir in Missouri. Transactions of the American Fisheries Society, 97(2):138–145, April 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Finucane:1969:ATM

[Fin69a] John H. Finucane. Antimycin as a toxicant in a marine habitat. Transactions of the American Fisheries Society, 98(2):288–292, April 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Finucane:1969:EPT

[Fin69b] John H. Finucane. Ecology of the pompano (Trachinotus carolinus) and the permit (T. falcatus) in Florida. Transactions of the American Fisheries Society, 98(3):478–486, July 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fischler:1965:UCE

[Fis65] Kenneth J. Fischler. The use of catch–effort, catch-sampling, and tagging data to estimate a population of blue crabs. Transactions of the American Fisheries Society, 94(4):287–310, October 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Flickinger:1969:DSF

[Fli69] Stephen A. Flickinger. Determination of sexes in the fathead minnow. Transactions of the American Fisheries Society, 98(3):526– 527, July 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ferguson:1966:DEU

[FLM66] Denzel E. Ferguson, J. Larry Ludke, and George G. Murphy. Dynamics of endrin uptake and release by resistant and susceptible strains of mosquitofish. Transactions of the American Fisheries Society, 95(4):335–344, October 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Forney:1961:YCD

[For61] John L. Forney. Year-class distribution of walleyes collected by five types of gear. Transactions of the American Fisheries Society, 90 (3):308–311, July 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Forney:1963:DMM

[For63] John L. Forney. Distribution and movement of marked walleyes in Oneida Lake, New York. Transactions of the American Fisheries Society, 92(1):47–52, January 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ferguson:1963:SFT

[FR63] Robert G. Ferguson and Henry A. Regier. Selectivity of four trawl cod ends toward smelt. Transactions of the American Fisheries Society, 92(2):125–131, April 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Finnell:1969:DVM

[FR69] Larry M. Finnell and Edward B. Reed. The diel vertical movements of Kokanee salmon, Oncorhynchus nerka, in Granby Reservoir, Colorado. Transactions of the American Fisheries Society, 98(2):245–252, April 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fraser:1968:DRB

[Fra68] J. M. Fraser. Differential recovery of brook trout planted by hand and by air drop. Transactions of the American Fisheries Society, 97(1):32–36, January 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

French:1965:VAH

[Fre65] Robert R. French. Visceral adhesions in high-seas salmon. Transactions of the American Fisheries Society, 94(2):177–181, April 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fremling:1967:MMR

[Fre67] Calvin R. Fremling. Methods for mass-rearing Hexagenia mayflies (Ephemeroptera: Ephemeridae). Transactions of the American Fisheries Society, 96(4):407–410, October 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fremling:1968:DME

[Fre68] Calvin R. Fremling. Documentation of a mass emergence of Hexagenia mayflies from the Upper Mississippi River. Transactions of the American Fisheries Society, 97(3):278–281, July 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fribourgh:1965:EPT

[Fri65] James H. Fribourgh. The effects of potassium tellurite on goldfish spermatozoan activity. Transactions of the American Fisheries Society, 94(4):399–402, October 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Franklin:1960:NDS

[FS60] Donald R. Franklin and Lloyd L. Smith, Jr. Note on development of scale patterns in the northern pike, Esox lucius L. Transactions of the American Fisheries Society, 89(1):83, January 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Franklin:1963:ELH

[FS63] Donald R. Franklin and Lloyd L. Smith, Jr. Early life history of the northern pike, Esox lucius L., with special reference to the factors influencing the numerical strength of year classes. Transactions of the American Fisheries Society, 92(2):91–110, April 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Foye:1965:EPS

[FS65] Robert E. Foye and Matthew Scott. Effects of pressure on survival of six species of fish. Transactions of the American Fisheries Society, 94(1):88–91, January 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Foster:1966:EAF

[FSC66] Neal R. Foster, Arthur Scheier, and John Cairns, Jr. Effects of ABS on feeding behavior of flagfish, Jordanella floridae. Transactions of the American Fisheries Society, 95(1):109–110, January 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fuchs:1967:LHE

[Fuc67] Everett H. Fuchs. Life history of the emerald shiner, Notropis atherinoides, in Lewis and Clark Lake, South Dakota. Transactions of the American Fisheries Society, 96(3):247–256, July 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Fischler:1962:BCM

[FW62a] Kenneth J. Fischler and Charles H. Walburg. Blue crab movement in coastal South Carolina, 1958–59. Transactions of the American Fisheries Society, 91(3):275–278, July 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Flick:1962:PSW

[FW62b] William A. Flick and Dwight A. Webster. Problems in sampling wild and domestic stocks of brook trout (Salvelinus fontinalis). Transactions of the American Fisheries Society, 91(2):140–144, April 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Flick:1964:CFY

[FW64] William A. Flick and Dwight A. Webster. Comparative first year survival and production in wild and domestic strains of brook trout, Salvelinus fontinalis. Transactions of the American Fisheries Society, 93(1):58–69, January 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Galbraith:1967:SSP

[Gal67] Merle G. Galbraith, Jr. Size-selective predation on Daphnia by rainbow trout and yellow perch. Transactions of the American Fisheries Society, 96(1):1–10, January 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gammon:1963:CFY

[Gam63] James R. Gammon. Conversion of food in young muskellunge. Transactions of the American Fisheries Society, 92(2):183–184, April 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gasaway:1962:ECA

[Gas62] Charles R. Gasaway. Effects of cutting aquatic vegetation and treatment of the resulting mats in ponds. Transactions of the American Fisheries Society, 91(1):8–13, January 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gunning:1969:FRE

[GB69] Gerald E. Gunning and Tim M. Berra. Fish repopulation of experimentally decimated segments in the headwaters of two streams. Transactions of the American Fisheries Society, 98(2):305–308, April 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Garside:1962:EHA

[GC62] E. T. Garside and W. J. Christie. Experimental hybridization among three coregonine fishes. Transactions of the American Fisheries Society, 91(2):196–200, April 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gill:1966:VAS

[GF66] Charles D. Gill and Donald M. Fisk. Vertebral abnormalities in sockeye, pink, and chum salmon. Transactions of the American Fisheries Society, 95(2):177–182, April 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gross:1963:RBT

[GFR63] W. L. Gross, P. O. Fromm, and E. W. Roelofs. Relationship between thyroid and growth in green sunfish, Lepomis cyanellus (Rafinesque). Transactions of the American Fisheries Society, 92 (4):401–408, October 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ghent:1965:NMA

[GG65] A. W. Ghent and B. Grinstead. A new method of assessing contagion, applied to a distribution of redear sunfish. Transactions of the American Fisheries Society, 94(2):135–142, April 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Garlick:1960:DSF

[GH60] Lewis R. Garlick and Jack E. Hemphill. The development of sport fishery resources on Indian lands in the southwest. Transactions of the American Fisheries Society, 89(3):271–273, July 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ghent:1966:BCA

[Ghe66] Arthur W. Ghent. Binomial corner-association assessment of contagion: a response to certain criticisms of the method. Transactions of the American Fisheries Society, 95(4):437–441, October 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gilderhus:1966:SES

[Gil66] Philip A. Gilderhus. Some effects of sublethal concentrations of sodium arsenite on bluegills and the aquatic environment. Transactions of the American Fisheries Society, 95(3):289–296, July 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Groves:1969:PTB

[GJ69] Alan B. Groves and Irving W. Jones. Permanent thermal branding of Coho salmon, Oncorhynchus kisutch. Transactions of the American Fisheries Society, 98(2):334–335, April 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Glude:1966:CSF

[Glu66] John B. Glude. Criteria of success and failure in the management of shellfisheries. Transactions of the American Fisheries Society, 95 (3):260–263, July 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Groves:1965:TMT

[GN65] Alan B. Groves and Anthony J. Novotny. A thermal-marking technique for juvenile salmonids. Transactions of the American Fisheries Society, 94(4):386–389, October 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Geen:1968:LIM

[GN68] G. H. Geen and T. G. Northcote. Latex injection as a method of marking large catostomids for long term study. Transactions of the American Fisheries Society, 97(3):281–282, July 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goodyear:1967:FHT

[Goo67] C. Phillip Goodyear. Feeding habits of three species of gars, Lepisosteus, along the Mississippi Gulf Coast. Transactions of the American Fisheries Society, 96(3):297–300, July 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gordon:1961:FAS

[Gor61] William G. Gordon. Food of the American smelt in Saginaw Bay, Lake Huron. Transactions of the American Fisheries Society, 90 (4):439–443, October 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Graham:1961:BUC

[Gra61] Richard J. Graham. Biology of the Utah chub in Hebgen Lake, Montana. Transactions of the American Fisheries Society, 90(3): 269–276, July 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gray:1964:TLD

[Gra64] George W. Gray, Jr. Tag loss during ecdysis by the king crab, Paralithodes camtschatica (Tilesius). Transactions of the American Fisheries Society, 93(3):303–304, July 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Graham:1968:TMF

[Gra68] Herbert W. Graham. Trends in the marine fisheries of the Continental Shelf of the Eastern United States. Transactions of the American Fisheries Society, 97(1):77–82, January 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Greene:1962:WBF

[Gre62] George N. Greene. White bass feeding: Scent or sight. Transactions of the American Fisheries Society, 91(3):326, July 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Green:1964:CSB

[Gre64] David M. Green, Jr. A comparison of stamina of brook trout from wild and domestic parents. Transactions of the American Fisheries Society, 93(1):96–100, January 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Green:1967:RTS

[Gre67] Roger E. Green. Relationship of the thermocline to success of purse seining for tuna. Transactions of the American Fisheries Society, 96(2):126–130, April 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gregory:1968:OFT

[Gre68] Richard W. Gregory. Occurrence of fructose in trout seminal plasma. Transactions of the American Fisheries Society, 97(2): 203–204, April 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Griffith:1966:RBS

[Gri66] J. Q. Griffith, Jr. Relation between serum protein and total serum cholesterol levels in certain fish. Transactions of the American Fisheries Society, 95(2):214–216, April 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Groves:1960:SPJ

[Gro60] Alan B. Groves. Swimming performance of juvenile sockeye salmon in salt and fresh water. Transactions of the American Fisheries Society, 89(2):218–221, April 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Grover:1968:HBS

[Gro68] John H. Grover. Hemosiderin in bluegill spleens. Transactions of the American Fisheries Society, 97(1):48–50, January 1968. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gunstrom:1967:OKC

[Gun67] Gary K. Gunstrom. The occurrence of a katadidymus Chinook salmon alevin. Transactions of the American Fisheries Society, 96(2):214–216, April 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Goodwin:1968:APA

[GV68] Walter F. Goodwin and Thomas L. Vaughn. An adult pugheaded American shad Alosa sapidissima. Transactions of the American Fisheries Society, 97(1):50, January 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Gakstatter:1967:EDC

[GW67] Jack H. Gakstatter and Charles M. Weiss. The elimination of DDT-C^{14}, dieldrin-c^{14}, and lindane-c^{14} from fish following a single sublethal exposure in aquaria. Transactions of the American Fisheries Society, 96(3):301–307, July 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hayes:1964:PCN

[HA64] F. R. Hayes and E. H. Anthony. Productive capacity of North American lakes as related to the quantity and the trophic level of fish, the lake dimensions, and the water chemistry. Transactions of the American Fisheries Society, 93(1):53–57, January 1964. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hale:1968:OBT

[Hal68] John G. Hale. Observations on brook trout, Salvelinus fontinalis spawning in 10-gallon aquaria. Transactions of the American Fisheries Society, 97(3):299–300, July 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hansen:1965:FON

[Han65] Donald F. Hansen. Further observations on nesting of the white crappie, Pomoxis annularis. Transactions of the American Fisheries Society, 94(2):182–184, April 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hansen:1969:APU

[Han69] David J. Hansen. Avoidance of pesticides by untrained sheepshead minnows. Transactions of the American Fisheries Society, 98(3): 426–429, July 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hartman:1962:UGN

[Har62] Gordon F. Hartman. Use of gill-net rollers in fishery investigations. Transactions of the American Fisheries Society, 91(2):224– 225, April 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hartman:1965:ASS

[Har65] G. F. Hartman. An aquarium with simulated stream flow. Transactions of the American Fisheries Society, 94(3):274–276, July 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harvey:1967:SLW

[Har67] Harold H. Harvey. Supersaturation of lake water with a precaution to hatchery usage. Transactions of the American Fisheries Society, 96(2):194–201, April 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Halver:1969:AAR

[HAS69] John E. Halver, Laurence M. Ashley, and Robert R. Smith. Ascorbic acid requirements of Coho salmon and rainbow trout. Transactions of the American Fisheries Society, 98(4):762–771, October 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hatch:1961:ROF

[Hat61] Richard W. Hatch. Regular occurrence of false annuli in four brook trout populations. Transactions of the American Fisheries Society, 90(1):6–12, January 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hauser:1969:LHM

[Hau69] William J. Hauser. Life history of the mountain sucker, Catostomus platyrhynchus, in Montana. Transactions of the American Fisheries Society, 98(2):209–215, April 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Havey:1960:RGM

[Hav60] Keith A. Havey. Recovery, growth, and movement of hatcheryreared Lake Atlantic salmon at Long Pond, Maine. Transactions of the American Fisheries Society, 89(2):212–217, April 1960. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Havey:1961:RAA

[Hav61] Keith A. Havey. Restoration of anadromous alewives at Long Pond, Maine. Transactions of the American Fisheries Society, 90(3):281– 286, July 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hooper:1964:RML

[HB64a] F. F. Hooper and R. C. Ball. Responses of a marl lake to fertilization. Transactions of the American Fisheries Society, 93(2): 164–173, April 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hunt:1964:FYS

[HB64b] Robert L. Hunt and Oscar M. Brynildson. A five-year study of a headwaters trout refuge. Transactions of the American Fisheries Society, 93(2):194–197, April 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hooper:1960:URR

[HC60] Archie D. Hooper and Johnie H. Crance. Use of rotenone in restoring balance to overcrowded fish populations in Alabama lakes. Transactions of the American Fisheries Society, 89(4):351–357, October 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Holland:1966:ISP

[HCB66] H. T. Holland, David L. Coppage, and Philip A. Butler. Increased sensitivity to pesticides in sheepshead minnows. Transactions of the American Fisheries Society, 95(1):110–112, January 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Heman:1969:MFP

[HCR69] M. LeRoy Heman, Robert S. Campbell, and Lee C. Redmond. Manipulation of fish populations through reservoir drawdown. Transactions of the American Fisheries Society, 98(2):293–304, April 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hester:1962:MPS

[HD62] F. Eugene Hester and J. S. Dendy. A multiple-plate sampler for aquatic macroinvertebrates. Transactions of the American Fisheries Society, 91(4):420–421, October 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hitz:1965:CYE

[HD65] Charles R. Hitz and Allan C. Delacy. Clearing of yolk in eggs of the rockfishes, Sebastodes caurinus and s. auriculatus. Transactions of the American Fisheries Society, 94(2):194–195, April 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Houser:1967:EST

[HD67] Alfred Houser and James E. Dunn. Estimating the size of the threadfin shad population in Bull Shoals Reservoir from midwater trawl catches. Transactions of the American Fisheries Society, 96(2):176–184, April 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Heard:1962:USS

[Hea62] William R. Heard. The use and selectivity of small-meshed gill nets at Brooks Lake, Alaska. Transactions of the American Fisheries Society, 91(3):263–268, July 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Heard:1965:LCL

[Hea65] William R. Heard. Limnetic cottid larvae and their utilization as food by juvenile sockeye salmon. Transactions of the American Fisheries Society, 94(2):191–193, April 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Heckman:1969:ECL

[Hec69] James Robert Heckman. Embryological comparison of Lepomis macrochirus X macrochirus and Lepomis macrochirus X gibbosus. Transactions of the American Fisheries Society, 98(4):669–675, October 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Helle:1966:BDA

[Hel66] John H. Helle. Behavior of displaced adult pink salmon. Transactions of the American Fisheries Society, 95(2):188–195, April 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Held:1969:SES

[Hel69] John W. Held. Some early summer foods of the shovelnose sturgeon in the Missouri River. Transactions of the American Fisheries Society, 98(3):514–517, July 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Herman:1968:FF

[Her68] Roger Lee Herman. Fish furunculosis 1952–1966. Transactions of the American Fisheries Society, 97(3):221–230, July 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hettler:1968:AFA

[Het68] William F. Hettler, Jr. Artificial fertilization among yellowfin and Gulf menhaden (Brevoortia) and their hybrid. Transactions of the American Fisheries Society, 97(2):119–123, April 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harbeck:1961:MEM

[HG61] G. Earl Harbeck, Jr. and A. F. C. Greene. A method of estimating maximum temperatures to be expected in ponds and small lakes. Transactions of the American Fisheries Society, 90(4):494–497, October 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Houser:1964:EGN

[HG64] Alfred Houser and Arthur W. Ghent. An experimental gill net based on the principle of the Latin Square. Transactions of the American Fisheries Society, 93(3):233–242, July 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Herr:1967:TSA

[HGC67] F. Herr, E. Greselin, and C. Chappel. Toxicology studies of antimycin, a fish eradicant. Transactions of the American Fisheries Society, 96(3):320–326, July 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hergenrader:1967:SCS

[HH67] Gary L. Hergenrader and Arthur D. Hasler. Seasonal changes in swimming rates of yellow perch in Lake Mendota as measured by sonar. Transactions of the American Fisheries Society, 96(4):373– 382, October 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hale:1969:SSA

[HH69] John G. Hale and Donald A. Hilden. Spawning and some aspects of early life history of brook trout, Salvelinus fontinalis (Mitchill), in the laboratory. Transactions of the American Fisheries Society, 98 (3):473–477, July 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Henderson:1966:UTU

[HHC66] H. Francis Henderson, Arthur D. Hasler, and Gerald G. Chipman. An ultrasonic transmitter for use in studies of movements of fishes. Transactions of the American Fisheries Society, 95(4):350–356, October 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hilsenhoff:1965:ETB

[Hil65] William L. Hilsenhoff. The effect of toxaphene on the benthos in a thermally-stratified lake. Transactions of the American Fisheries Society, 94(3):210–213, July 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hiltibran:1967:ESH

[Hil67] Robert C. Hiltibran. Effects of some herbicides on fertilized fish eggs and fry. Transactions of the American Fisheries Society, 96 (4):414–416, October 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hill:1968:OPS

[Hil68] Loren G. Hill. Oxygen preference in the spring cavefish, Chologaster agassizi. Transactions of the American Fisheries Society, 97(4): 448–454, October 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hiner:1961:PNP

[Hin61] Laurence E. Hiner. Propagation of northern pike. Transactions of the American Fisheries Society, 90(3):298–302, July 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hirschhorn:1966:EGP

[Hir66] George Hirschhorn. Effect of growth pattern and exploitation features on yields of pink salmon in the Northwestern North Pacific. Transactions of the American Fisheries Society, 95(1):39–51, January 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hitz:1962:SBR

[Hit62] Charles R. Hitz. Seasons of birth of rockfish (Sebastodes spp.) in Oregon coastal waters. Transactions of the American Fisheries Society, 91(2):231–233, April 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hoberman:1962:GRN

[HJ62] John M. Hoberman and Albert C. Jensen. The growth rate of New England Pollock. Transactions of the American Fisheries Society, 91(2):227–228, April 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Horak:1967:ICM

[HK67] Donald L. Horak and William D. Klein. Influence of capture methods on fishing success, stamina, and mortality of rainbow trout (Salmo gairdneri) in Colorado. Transactions of the American Fisheries Society, 96(2):220–222, April 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Heaton:1967:CSG

[HM67] Le Roy H. Heaton and Ivan B. Mcelwain. Complement in Salmo gairdneri and Salmo trutta. Transactions of the American Fisheries Society, 96(3):351, July 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Haw:1969:RUS

[HM69] Frank Haw and Stephen B. Mathews. Recreational use of surplus hatchery Coho salmon. Transactions of the American Fisheries Society, 98(3):487–496, July 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hnath:1969:TAA

[Hna69] John G. Hnath. Transfer of an adult acanthocephalan from one fish host to another. Transactions of the American Fisheries Society, 98(2):332, April 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hodder:1966:TFR

[Hod66] V. M. Hodder. Two further records of the American shad in Newfoundland waters. Transactions of the American Fisheries Society, 95(2):228–229, April 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hoff:1968:OLA

[Hof68] James G. Hoff. Occurrence of Liparis atlanticus in Aequipecten irradians in Buzzards Bay, Massachusetts. Transactions of the American Fisheries Society, 97(3):278, July 1968. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hoopes:1960:UMC

[Hoo60] David Townsend Hoopes. Utilization of mayflies and caddis flies by some Mississippi River fishes. Transactions of the American Fisheries Society, 89(1):32–34, January 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Horak:1966:DMD

[Hor66] Donald L. Horak. A densitometric method for determining bodyfat concentration in fish. Transactions of the American Fisheries Society, 95(3):318–320, July 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hoss:1964:AZF

[Hos64] Donald E. Hoss. Accumulation of zinc-65 by flounder of the genus Paralichthys. Transactions of the American Fisheries Society, 93 (4):364–368, October 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hoss:1967:MPL

[Hos67] Donald E. Hoss. Marking post-larval paralichthid flounders with radioactive elements. Transactions of the American Fisheries Society, 96(2):151–156, April 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Houde:1964:STC

[Hou64] Edward D. Houde. The strigeid trematode, Crassiphiala bulboglossa, from the blacknose dace, Rhinichthys atratulus. Transactions of the American Fisheries Society, 93(3):304–306, July 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Houser:1965:GPF

[Hou65] Alfred Houser. Growth of paddlefish in Fort Gibson Reservoir, Oklahoma. Transactions of the American Fisheries Society, 94(1): 91–93, January 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Houde:1967:FPY

[Hou67] Edward D. Houde. Food of pelagic young of the walleye, Stizostedion vitreum vitreum, in Oneida Lake, New York. Transactions of the American Fisheries Society, 96(1):17–24, January 1967. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hoffman:1965:BSU

[HP65] Glenn L. Hoffman and Robert E. Putz. The black-spot (Uvulifer ambloplitis: Trematoda: Strigeoidea) of centrarchid fishes. Transactions of the American Fisheries Society, 94(2):143–151, April 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hooper:1961:URH

[HPS61] F. F. Hooper, H. A. Podoliak, and S. F. Snieszko. Use of radioisotopes in hydrobiology and fish culture. Transactions of the American Fisheries Society, 90(1):49–57, January 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Harris:1968:SPS

[HR68] Alva H. Harris and Curt D. Rose. Shrimp predation by the sea catfish, Galeichthys felis. Transactions of the American Fisheries Society, 97(4):503–504, October 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hansen:1964:CSM

[HS64] Martin J. Hansen and Thomas M. Stauffer. Cadmium sulfide and mercuric sulfide for marking sea lamprey larvae. Transactions of the American Fisheries Society, 93(1):21–26, January 1964. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hess:1967:EES

[HS67] Paul W. Hess and Ruth A. Smith. Electrophoresis of elasmobranch serum proteins with cellulose acetate. Transactions of the American Fisheries Society, 96(2):131–133, April 1967. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hartman:1962:SBS

[HSH62] Wilbur L. Hartman, Charles W. Strickland, and David T. Hoopes. Survival and behavior of sockeye salmon fry migrating into Brooks Lake, Alaska. Transactions of the American Fisheries Society, 91(2):133–139, April 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Horak:1964:UVG

[HT64a] Donald L. Horak and Howard A. Tanner. The use of vertical gill nets in studying fish depth distribution, Horsetooth Reservoir, Colorado. Transactions of the American Fisheries Society, 93(2):137– 145, April 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Howell:1964:AES

[HT64b] John H. Howell and Paul M. Thomas. Anesthetic effect of 4styrylpyridine on lamprey and fish. Transactions of the American Fisheries Society, 93(2):206–208, April 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hubley:1961:ILS

[Hub61] Raymond C. Hubley, Jr. Incidence of lamprey scarring on fish in the Upper Mississippi River, 1956–58. Transactions of the American Fisheries Society, 90(1):83–85, January 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hubley:1963:MTC

[Hub63] Raymond C. Hubley, Jr. Movement of tagged channel catfish in the Upper Mississippi River. Transactions of the American Fisheries Society, 92(2):165–168, April 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hubbs:1968:OEC

[Hub68] Clark Hubbs. An opinion on the effects of cichlid releases in North America. Transactions of the American Fisheries Society, 97(2): 197–198, April 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hunt:1960:DRF

[Hun60] Burton P. Hunt. Digestion rate and food consumption of Florida gar, warmouth, and largemouth bass. Transactions of the American Fisheries Society, 89(2):206–211, April 1960. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hunt:1965:DWB

[Hun65a] Robert L. Hunt. Dispersal of wild brook trout during their first summer of life. Transactions of the American Fisheries Society, 94(2):186–188, April 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hunt:1965:FNP

[Hun65b] Robert L. Hunt. Food of northern pike in a Wisconsin trout stream. Transactions of the American Fisheries Society, 94(1):95–97, January 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Hunn:1969:CCR

[Hun69] Joseph B. Hunn. Chemical composition of rainbow trout urine following acute hypoxic stress. Transactions of the American Fisheries Society, 98(1):20–22, January 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Heard:1968:FTU

[HV68] William R. Heard and Louis E. Vogele. A flag tag for underwater recognition of individual fish by divers. Transactions of the American Fisheries Society, 97(1):55–57, January 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Herting:1967:RPF

[HW67] Gerald E. Herting and Arthur Witt, Jr. The role of physical fitness of forage fishes in relation to their vulnerability to predation by bowfin (Amia calva). Transactions of the American Fisheries Society, 96(4):427–430, October 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Haynes:1969:BNS

[HW69] Evan B. Haynes and Roland L. Wigley. Biology of the northern shrimp, Pandalus borealis, in the Gulf of Maine. Transactions of the American Fisheries Society, 98(1):60–76, January 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Herrmann:1962:IOC

[HWD62] Robert B. Herrmann, Charles E. Warren, and Peter Doudoroff. Influence of oxygen concentration on the growth of juvenile Coho salmon. Transactions of the American Fisheries Society, 91(2): 155–167, April 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Isaac:1963:SCF

[IB63] Gary W. Isaac and Carl E. Bond. Standing crops of fish in Oregon farm ponds. Transactions of the American Fisheries Society, 92(1):

25–29, January 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Iversen:1960:ABT

[II60] Edwin S. Iversen and C. P. Idyll. Aspects of the biology of the Tortugas pink shrimp, Penaeus duorarum. Transactions of the American Fisheries Society, 89(1):1–8, January 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Isaacson:1966:NLH

[II66] Peter A. Isaacson and Diane M. Isaacson. Notes on the life history of the black perch, Embiotoca jacksoni, Agassiz. Transactions of the American Fisheries Society, 95(1):107–109, January 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Iversen:1962:EAM

[IM62] E. S. Iversen and Alan W. Moffett. Estimation of abundance and mortality of a spotted seatrout population. Transactions of the American Fisheries Society, 91(4):395–398, October 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Isaacson:1964:LWR

[Isa64] Peter A. Isaacson. Length–weight relationship of the white croaker. Transactions of the American Fisheries Society, 93(3):302–303, July 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Isaacson:1965:PBP

[Isa65] Peter A. Isaacson. Pugheadedness in the black perch, Embiotoca jacksoni. Transactions of the American Fisheries Society, 94(1): 98, January 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Islam:1968:FMS

[IT68] Bhuiyan N. Islam and G. B. Talbot. Fluvial migration, spawning, and fecundity of Indus River Hilsa, Hilsa ilisha. Transactions of the American Fisheries Society, 97(4):350–355, October 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Incerpi:1969:FLS

[IW69] Angelo Incerpi and Kendall Warner. Fecundity of landlocked salmon, Salmo salar. Transactions of the American Fisheries Society, 98(4):720–723, October 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jensen:1967:ETG

[Jen67] Albert C. Jensen. Effects of tagging on the growth of cod. Transactions of the American Fisheries Society, 96(1):37–41, January 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jenkins:1969:OCC

[Jen69] Thomas M. Jenkins, Jr. Observations on color changes of brown and rainbow trout (Salmo trutta and S. gairdneri) in stream habitats, with description of an unusual color pattern in brown trout. Transactions of the American Fisheries Society, 98(3):517–519, July 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jernejcic:1969:UEC

[Jer69] Frank Jernejcic. Use of emetics to collect stomach contents of walleye and largemouth bass. Transactions of the American Fisheries Society, 98(4):698–702, October 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jensen:1960:OSC

[JF60] Albert C. Jensen and Raymond L. Fritz. Observations on the stomach contents of the silver hake. Transactions of the American Fisheries Society, 89(2):239–240, April 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jensen:1964:ETO

[JG64a] Loren D. Jensen and Arden R. Gaufin. Effects of ten organic insecticides on two species of stonefly naiads. Transactions of the American Fisheries Society, 93(1):27–34, January 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jensen:1964:LTE

[JG64b] Loren D. Jensen and Arden R. Gaufin. Long-term effects of organic insecticides on two species of stonefly naiads. Transactions of the American Fisheries Society, 93(4):357–363, October 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jenkins:1969:RTE

[JK69] Thomas M. Jenkins, Jr. and Aaron Klain. A regulated-temperature electric tool for marking fish. Transactions of the American Fisheries Society, 98(2):338–340, April 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jones:1963:PPA

[JM63] Bernard R. Jones and John B. Moyle. Populations of plankton animals and residual chlorinated hydrocarbons in soils of six Minnesota ponds treated for control of mosquito larvae. Transactions of the American Fisheries Society, 92(3):211–215, July 1963. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:1969:MLS

[JM69] Fritz H. Johnson and John B. Moyle. Management of a large shallow winterkill lake in Minnesota for the production of pike (Esox lucius). Transactions of the American Fisheries Society, 98(4): 691–697, October 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:1961:WES

[Joh61] Fritz H. Johnson. Walleye egg survival during incubation on several types of bottom in Lake Winnibigoshish, Minnesota, and connecting waters. Transactions of the American Fisheries Society, 90(3): 312–322, July 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:1962:STA

[Joh62] James H. Johnson. Sea temperatures and the availability of albacore off the coasts of Oregon and Washington. Transactions of the American Fisheries Society, 91(3):269–274, July 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:1965:EFP

[Joh65] Murray G. Johnson. Estimates of fish populations in warmwater streams by the removal method. Transactions of the American Fisheries Society, 94(4):350–357, October 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:1968:PFR

[Joh68a] Donald W. Johnson. Pesticides and fishes — a review of selected literature. Transactions of the American Fisheries Society, 97(4):

398–424, October 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:1968:ACC

[Joh68b] James E. Johnson. Albinistic carp, Cyprinus carpio, from Roosevelt Lake, Arizona. Transactions of the American Fisheries Society, 97(2):209–210, April 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jones:1966:TNT

[Jon66] Benjamin F. Jones. Two new tools for applying disk tags to fish. Transactions of the American Fisheries Society, 95(3):323–325, July 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Joyner:1961:EZE

[Joy61] Timothy Joyner. Exchange of zinc with environmental solutions by the brown bullhead. Transactions of the American Fisheries Society, 90(4):444–448, October 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Jensen:1961:MTH

[JW61] Albert C. Jensen and John P. Wise. Movement of tagged halibut off New England — II. Transactions of the American Fisheries Society, 90(4):489–490, October 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Johnson:1962:EAS

[JW62] Merle W. Johnson and Leonard Wroblewski. Errors associated with a systematic sampling creel census. Transactions of the American Fisheries Society, 91(2):201–207, April 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Katz:1961:ATS

[Kat61] Max Katz. Acute toxicity of some organic insecticides to three species of salmonids and to the threespine stickleback. Transactions of the American Fisheries Society, 90(3):264–268, July 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kilambi:1969:FTS

[KB69] Raj V. Kilambi and Raymond E. Baglin, Jr. Fecundity of the threadfin shad, Dorosoma petenense, in Beaver and Bull Shoals Reservoirs. Transactions of the American Fisheries Society, 98(2): 320–322, April 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Katz:1961:TES

[KC61] Max Katz and George G. Chadwick. Toxicity of endrin to some Pacific Northwest fishes. Transactions of the American Fisheries Society, 90(4):394–397, October 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Krumholz:1968:CMF

[KC68] Louis A. Krumholz and Harold S. Cavanah. Comparative morphometry of freshwater drum from two Midwestern localities. Transactions of the American Fisheries Society, 97(4):429–441, October 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kallman:1962:DDT

[KCN62] Burton J. Kallman, Oliver B. Cope, and Richard J. Navarre. Distribution and detoxication of toxaphene in Clayton Lake, New Mexico. Transactions of the American Fisheries Society, 91(1):14–22, January 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kiser:1963:ERZ

[KDO63] R. W. Kiser, John R. Donaldson, and Paul R. Olson. The effect of rotenone on zooplankton populations in freshwater lakes. Transactions of the American Fisheries Society, 92(1):17–24, January 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Keeton:1965:ASM

[Kee65] Dee Keeton. Application of Stoeltzner’s method to determine growth of fish scales. Transactions of the American Fisheries Society, 94(1):93–94, January 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kelley:1962:SMF

[Kel62] John W. Kelley. Sexual maturity and fecundity of the largemouth bass, Micropterus salmoides (Lacépède), in Maine. Transactions of the American Fisheries Society, 91(1):23–28, January 1962. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kelly:1967:MFM

[Kel67] William H. Kelly. Marking freshwater and a marine fish by injected dyes. Transactions of the American Fisheries Society, 96(2):163– 175, April 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kemmerer:1968:MDF

[Kem68] Andrew J. Kemmerer. A method to determine fertilization requirements of a small sport fishing lake. Transactions of the American Fisheries Society, 97(4):425–428, October 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ketchen:1968:TMF

[Ket68] K. S. Ketchen. Trends in marine fisheries along the Pacific Coast of North America. Transactions of the American Fisheries Society, 97(1):82–88, January 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kevern:1966:FRC

[Kev66] N. R. Kevern. Feeding rate of carp estimated by a radioisotopic method. Transactions of the American Fisheries Society, 95(4): 363–371, October 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Khandker:1964:SSY

[Kha64] N. Alam Khandker. Sponge as shelter for young spiny lobsters. Transactions of the American Fisheries Society, 93(2):204, April 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Klein:1966:MTC

[Kle66] W. D. Klein. Mortality of trout caught on artificial lures and released by fishermen. Transactions of the American Fisheries Society, 95(3):326–328, July 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Krumholz:1964:CFP

[KM64] Louis A. Krumholz and W. L. Minckley. Changes in the fish population in the Upper Ohio River following temporary pollution abatement. Transactions of the American Fisheries Society, 93(1):1–5, January 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kendle:1965:DHO

[KM65] Earl R. Kendle and Larry A. Morris. A device for holding objects in the stomachs of fish. Transactions of the American Fisheries Society, 94(2):193–194, April 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Knight:1963:ELD

[Kni63] Alexis E. Knight. The embryonic and larval development of the rainbow trout. Transactions of the American Fisheries Society, 92 (4):344–355, October 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Klein:1969:EIF

[KOB69] W. D. Klein, O. W. Olsen, and D. C. Bowden. Effects of intestinal fluke, Crepidostomum farionis, on rainbow trout, Salmo gairdnerii. Transactions of the American Fisheries Society, 98(1):1–6, January 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Konar:1969:LST

[Kon69] S. K. Konar. Laboratory studies on two organophosphorus insecticides, DDVP and phosphamidon, as selective toxicants. Transactions of the American Fisheries Society, 98(3):430–437, July 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kramer:1960:FYG

[KS60] Robert H. Kramer and Lloyd L. Smith, Jr. First-year growth of the largemouth bass, Micropterus salmoides (Lacépède), and some related ecological factors. Transactions of the American Fisheries Society, 89(2):222–233, April 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kramer:1962:FYC

[KS62] Robert H. Kramer and Lloyd L. Smith, Jr. Formation of year classes in largemouth bass. Transactions of the American Fisheries Society, 91(1):29–41, January 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kramer:1965:ESW

[KS65] Robert H. Kramer and Lloyd L. Smith, Jr. Effects of suspended wood fiber on brown and rainbow trout eggs and alevins. Transactions of the American Fisheries Society, 94(3):252–258, July 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kleerekoper:1961:DPA

[KTW61] H. Kleerekoper, Grace Taylor, and R. Wilton. Diurnal periodicity in the activity of Petromyzon marinus and the effects of chemical stimulation. Transactions of the American Fisheries Society, 90(1): 73–78, January 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Kitchell:1968:RGD

[KW68] J. F. Kitchell and J. T. Windell. Rate of gastric digestion in pumpkinseed sunfish, Lepomis gibbosus. Transactions of the American Fisheries Society, 97(4):489–492, October 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Laarman:1969:ELF

[Laa69] Percy W. Laarman. Effects of a limited food supply on growth rates of Coho salmon and steelhead trout. Transactions of the American Fisheries Society, 98(3):393–397, July 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lackey:1968:VGN

[Lac68] Robert T. Lackey. Vertical gill nets for studying depth distribution of small fish. Transactions of the American Fisheries Society, 97 (3):296–299, July 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lackey:1969:FIS

[Lac69] Robert T. Lackey. Food interrelationships of salmon, trout, alewives, and smelt in a Maine lake. Transactions of the American Fisheries Society, 98(4):641–646, October 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lambou:1963:DSR

[Lam63] Victor W. Lambou. Distribution of sex ratios of bluegill and redear sunfish. Transactions of the American Fisheries Society, 92(4):435– 436, October 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lambou:1965:OSD

[Lam65] Victor W. Lambou. Observations on size distribution and spawning behavior of threadfin shad. Transactions of the American Fisheries Society, 94(4):385–386, October 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lander:1963:GLR

[Lan63] Robert H. Lander. Girth-length relationships in sockeye and chum salmon. Transactions of the American Fisheries Society, 92(3): 305–307, July 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Larkins:1964:SEF

[Lar64] Herbert A. Larkins. Some epipelagic fishes of the North Pacific Ocean, Bering Sea, and Gulf of Alaska. Transactions of the American Fisheries Society, 93(3):286–290, July 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Latta:1962:PMB

[Lat62] William C. Latta. Periodicity of mortality of brook trout during first summer of life. Transactions of the American Fisheries Society, 91(4):408–411, October 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Latta:1963:SEN

[Lat63] William C. Latta. Semiannual estimates of natural mortality of hatchery brook trout in lakes. Transactions of the American Fisheries Society, 92(1):53–59, January 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Latta:1965:RYY

[Lat65] William C. Latta. Relationship of young-of-the-year trout to mature trout and groundwater. Transactions of the American Fisheries Society, 94(1):32–39, January 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Laurence:1969:EEL

[Lau69] Geoffrey C. Laurence. The energy expenditure of largemouth bass larvae, Micropterus salmoides, during yolk absorption. Transactions of the American Fisheries Society, 98(3):398–405, July 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lewis:1966:FSB

[LB66] Robert M. Lewis and Rupert R. Bonner, Jr. Fecundity of the striped bass, Roccus saxatilis (Walbaum). Transactions of the American Fisheries Society, 95(3):328–331, July 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Larimore:1968:ETA

[LD68] R. Weldon Larimore and Michael J. Duever. Effects of temperature acclimation on the swimming ability of smallmouth bass fry. Transactions of the American Fisheries Society, 97(2):175–184, April 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Leader:1963:I

[Lea63] William C. Beckman Discussion Leader. Introduction. Transactions of the American Fisheries Society, 92(3):259, July 1963. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Legler:1967:CIS

[LED67] Donald W. Legler, E. Edward Evans, and Harry K. Dupree. Comparative immunology: Serum complement of freshwater fishes. Transactions of the American Fisheries Society, 96(3):237–242, July 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lewis:1962:SMD

[Lew62] Robert M. Lewis. Sexual maturity as determined from ovum diameters in striped bass from North Carolina. Transactions of the American Fisheries Society, 91(3):279–282, July 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lewis:1965:EMT

[Lew65] Robert M. Lewis. The effect of minimum temperature on the survival of larval Atlantic menhaden, Brevoortia tyrannus. Transactions of the American Fisheries Society, 94(4):409–412, October 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lewis:1966:EST

[Lew66] Robert M. Lewis. Effects of salinity and temperature on survival and development of larval Atlantic menhaden, Brevoortia tyrannus. Transactions of the American Fisheries Society, 95(4):423– 426, October 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lewis:1969:PFI

[Lew69] Stephen L. Lewis. Physical factors influencing fish populations in pools of a trout stream. Transactions of the American Fisheries Society, 98(1):14–19, January 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ludke:1968:SER

[LFB68] J. Larry Ludke, Denzel E. Ferguson, and W. David Burke. Some endrin relationships in resistant and susceptible populations of golden shiners, Notemigonus crysoleucas. Transactions of the American Fisheries Society, 97(3):260–263, July 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lewis:1961:FCL

[LGLB61] William M. Lewis, Gerald E. Gunning, Edward Lyles, and W. Leigh Bridges. Food choice of largemouth bass as a function of availability and vulnerability of food items. Transactions of the American Fisheries Society, 90(3):277–280, July 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lauer:1962:ICF

[LGS^{+}62] Gerald J. Lauer, Alfred R. Grzenda, Donald W. Shanklin, Hubert J. Webb, Robert E. O’Brien, Alfred R. Grzenda, Gerald J. Lauer, H. Page Nicholson, and H. Page Nicholson. Insecticide contamination in a farm pond: Part I — origin and duration, Part II — biological effects. Transactions of the American Fisheries Society, 91(2):213–222, April 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lewis:1964:VFO

[LH64] William M. Lewis and Don R. Helms. Vulnerability of forage organisms to largemouth bass. Transactions of the American Fisheries Society, 93(3):315–318, July 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lewis:1968:ETS

[LH68] Robert M. Lewis and William F. Hettler, Jr. Effect of temperature and salinity on the survival of young Atlantic menhaden, Brevoortia tyrannus. Transactions of the American Fisheries Society, 97(4): 344–349, October 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lewis:1968:LFD

[LHK68] William M. Lewis, Roy Heidinger, and Mark Konikoff. Loss of fishes over the drop box spillway of a lake. Transactions of the American Fisheries Society, 97(4):492–494, October 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lindroth:1963:SCS

[Lin63] Arne Lindroth. Salmon conservation in Sweden. Transactions of the American Fisheries Society, 92(3):286–291, July 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Latta:1961:NUD

[LM61] William C. Latta and Gerald F. Myers. Night use of a directcurrent electric shocker to collect trout in lakes. Transactions of the American Fisheries Society, 90(1):81–83, January 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lemke:1963:SEA

[LM63] Armond E. Lemke and Donald I. Mount. Some effects of alkyl benzene sulfonate on the bluegill, Lepomis macrochirus. Transactions of the American Fisheries Society, 92(4):372–378, October 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lauer:1966:PCS

[LNCT66] Gerald J. Lauer, H. Page Nicholson, William S. Cox, and John I. Teasley. Pesticide contamination of surface waters by sugar cane farming in Louisiana. Transactions of the American Fisheries Society, 95(3):310–316, July 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Logan:1963:WOB

[Log63] Sidney M. Logan. Winter observations on bottom organisms and trout in Bridger Creek, Montana. Transactions of the American Fisheries Society, 92(2):140–145, April 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Logie:1968:TAC

[Log68] R. R. Logie. Trends in the Atlantic Coast fisheries of Canada. Transactions of the American Fisheries Society, 97(1):89–92, January 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lowe:1964:CES

[Low64] Jack I. Lowe. Chronic exposure of spot, Leiostomus xanthurus, to sublethal concentrations of toxaphene in seawater. Transactions of the American Fisheries Society, 93(4):396–399, October 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lowry:1965:MCT

[Low65] Gerald R. Lowry. Movement of cutthroat trout, Salmo clarki clarki (Richardson) in three Oregon coastal streams. Transactions of the American Fisheries Society, 94(4):334–338, October 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lennon:1960:SCA

[LP60] Robert E. Lennon and Phillip S. Parker. The stoneroller, Campostoma anomalum (Rafinesque), in Great Smoky Mountains National Park. Transactions of the American Fisheries Society, 89 (3):263–270, July 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Leman:1966:SPM

[LP66] Bernie Leman and G. J. Paulik. Spill-pattern manipulation to guide migrant salmon upstream. Transactions of the American Fisheries Society, 95(4):397–407, October 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lux:1960:SFI

[LS60] Fred E. Lux and Lloyd L. Smith, Jr. Some factors influencing seasonal changes in angler catch in a Minnesota lake. Transactions of the American Fisheries Society, 89(1):67–79, January 1960. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Larsen:1961:MMT

[LS61] Howard N. Larsen and S. F. Snieszko. Modification of the microhematocrit technique with trout blood. Transactions of the American Fisheries Society, 90(2):139–142, April 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lewis:1963:RSC

[LSL63] William M. Lewis, Robert C. Summerfelt, and Alvin Lopinot. Results of stocking catchable-sized warmwater fishes in a lake with an established fish population. Transactions of the American Fisheries Society, 92(3):235–238, July 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lux:1960:LSY

[Lux60] Fred E. Lux. Length shrinkage of yellowtail flounder between live and landed condition. Transactions of the American Fisheries Society, 89(4):373–374, October 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lux:1969:LWR

[Lux69] F. E. Lux. Length–weight relationships of six New England flatfishes. Transactions of the American Fisheries Society, 98(4):617– 621, October 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Lagler:1962:PDV

[LW62] Karl F. Lagler and Asa T. Wright. Predation of the Dolly Varden, Salvelinus malma, on young salmons, Oncorhynchus spp., in an estuary of Southeastern Alaska. Transactions of the American Fisheries Society, 91(1):90–93, January 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Macphee:1966:IDA

[Mac66] Craig Macphee. Influence of differential angling mortality and stream gradient on fish abundance in a trout-sculpin biotope. Transactions of the American Fisheries Society, 95(4):381–387, October 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Magnuson:1969:DFC

[Mag69] John J. Magnuson. Digestion and food consumption by skipjack tuna (Katsuwonus pelamis). Transactions of the American Fisheries Society, 98(3):379–392, July 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Manz:1964:PDU

[Man64] Jerry V. Manz. A pumping device used to collect walleye eggs from offshore spawning areas in Western Lake Erie. Transactions of the American Fisheries Society, 93(2):204–206, April 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Manion:1967:DFL

[Man67] Patrick J. Manion. Diatoms as food of larval sea lampreys in a small tributary of Northern Lake Michigan. Transactions of the American Fisheries Society, 96(2):224–226, April 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Manion:1968:PSL

[Man68] Patrick J. Manion. Production of sea lamprey larvae from nests in two Lake Superior streams. Transactions of the American Fisheries Society, 97(4):484–486, October 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Martin:1966:SFH

[Mar66] N. V. Martin. The significance of food habits in the biology, exploitation, and management of Algonquin Park, Ontario, lake trout. Transactions of the American Fisheries Society, 95(4):415– 422, October 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Marak:1967:EEL

[Mar67] Robert R. Marak. Eggs and early larval stages of the offshore hake, Merluccius albidus. Transactions of the American Fisheries Society, 96(2):227–228, April 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mullan:1968:FLP

[MAR68] James W. Mullan, Richard L. Applegate, and William C. Rainwater. Food of logperch (Percina caprodes), and brook silverside (Labidesthes sicculus), in a new and old Ozark reservoir. Transactions of the American Fisheries Society, 97(3):300–305, July 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Marcy:1969:ADS

[Mar69] Barton C. Marcy, Jr. Age determinations from scales of Alosa pseudoharengus (Wilson) and Alosa aestivalis (Mitchill) in Connecticut waters. Transactions of the American Fisheries Society, 98(4):622–630, October 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mayhew:1969:AGF

[May69] J. K. Mayhew. Age and growth of flathead catfish in the Des Moines River, Iowa. Transactions of the American Fisheries Society, 98(1):

118–121, January 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McIntire:1962:EAF

[MB62] C. David McIntire and Carl E. Bond. Effects of artificial fertilization on plankton and benthos abundance in four experimental ponds. Transactions of the American Fisheries Society, 91(3): 303–312, July 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCrimmon:1965:AVR

[MB65] Hugh R. McCrimmon and Bryant Bidgood. Abnormal vertebrae in the rainbow trout with particular reference to electrofishing. Transactions of the American Fisheries Society, 94(1):84–88, January 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mason:1967:CSW

[MBD67] John W. Mason, Oscar M. Brynildson, and Paul E. Degurse. Comparative survival of wild and domestic strains of brook trout in streams. Transactions of the American Fisheries Society, 96(3): 313–319, July 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McFadden:1962:ECS

[MC62] James T. McFadden and Edwin L. Cooper. An ecological comparison of six populations of brown trout (Salmo trutta). Transactions of the American Fisheries Society, 91(1):53–62, January 1962. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCann:1969:CPM

[MC69] James A. McCann and Donna F. Cruse. Computer program for mark and recovery population estimates (CDC 3600 computer). Transactions of the American Fisheries Society, 98(2):332–334, April 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McFadden:1962:SRW

[MCA62] James T. McFadden, Edwin L. Cooper, and John K. Andersen. Sex ratio in wild populations of brown trout. Transactions of the American Fisheries Society, 91(1):94–95, January 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCarraher:1960:PHE

[McC60] D. B. McCarraher. Pike hybrids (Esox lucius × E. vermiculatus) in a Sandhill Lake, Nebraska. Transactions of the American Fisheries Society, 89(1):82–83, January 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCarraher:1961:ERN

[McC61a] D. B. McCarraher. Extension of the range of northern pike (Esox lucius). Transactions of the American Fisheries Society, 90(2): 227–228, April 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCombie:1961:GNS

[McC61b] A. M. McCombie. Gill-net selectivity of lake whitefish from Goderich–Bayfield Area, Lake Huron. Transactions of the American Fisheries Society, 90(3):337–340, July 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCarraher:1962:NPE

[McC62] D. B. McCarraher. Northern pike, Esox lucius, in alkaline lakes of Nebraska. Transactions of the American Fisheries Society, 91(3): 326–329, July 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McConnell:1963:PPF

[McC63a] William J. McConnell. Primary productivity and fish harvest in a small desert impoundment. Transactions of the American Fisheries Society, 92(1):1–12, January 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCrimmon:1963:RTL

[McC63b] Hugh H. McCrimmon. The recovery of tagged lake trout, Salvelinus namaycush, from Lake Simcoe, Ontario. Transactions of the American Fisheries Society, 92(3):315–317, July 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McComish:1967:FHB

[McC67] Thomas S. McComish. Food habits of bigmouth and smallmouth buffalo in Lewis and Clark Lake and the Missouri River. Transactions of the American Fisheries Society, 96(1):70–74, January 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McDowall:1968:INA

[McD68] R. M. McDowall. Interactions of the native and alien faunas of New Zealand and the problem of fish introductions. Transactions of the American Fisheries Society, 97(1):1–11, January 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McFadden:1969:KST

[McF69] James T. McFadden. Keynote Symposium: Trends in Freshwater Sport Fisheries of North America. Transactions of the American Fisheries Society, 98(1):136–150, January 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McLarney:1968:SHM

[McL68] William O. McLarney. Spawning habits and morphological variation in the coastrange sculpin, Cottus aleuticus, and the prickly sculpin, Cottus asper. Transactions of the American Fisheries Society, 97(1):46–48, January 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mclain:1968:EBT

[MD68] Alberton L. Mclain and Frederick H. Dahl. An electric beam trawl for the capture of larval lampreys. Transactions of the American Fisheries Society, 97(3):289–293, July 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meehan:1961:UFS

[Mee61] William R. Meehan. Use of a fishwheel in salmon research and management. Transactions of the American Fisheries Society, 90 (4):490–494, October 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meffert:1968:URL

[Mef68] P. Meffert. Ultrasonic recorder for locomotor activity studies. Transactions of the American Fisheries Society, 97(1):12–17, January 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meister:1962:ASP

[Mei62] Alfred L. Meister. Atlantic salmon production in Cove Brook, Maine. Transactions of the American Fisheries Society, 91(2): 208–212, April 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meyer:1962:LHT

[Mey62] William H. Meyer. Life history of three species of redhorse (Moxostoma) in the Des Moines River, Iowa. Transactions of the American Fisheries Society, 91(4):412–419, October 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meyer:1965:EUG

[Mey65] Fred P. Meyer. The experimental use of Guthion(R) as a selective fish eradicator. Transactions of the American Fisheries Society, 94 (3):203–209, July 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCombie:1960:SGN

[MF60] A. M. McCombie and F. E. J. Fry. Selectivity of gill nets for lake whitefish, Coregonus clupeaformis. Transactions of the American Fisheries Society, 89(2):176–184, April 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Minckley:1964:IAS

[MGC64] W. L. Minckley, Richard H. Goodyear, and James E. Craddock. Incidence of aberrant scalation in catostomid fishes from Doe Run, Meade County, Kentucky. Transactions of the American Fisheries Society, 93(2):202–203, April 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Musick:1968:VAH

[MH68] John A. Musick and James G. Hoff. Vertebral anomalies in humpbacked specimens of menhaden, Brevoortia tyrannus. Transactions of the American Fisheries Society, 97(3):277–278, July 1968. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Minckley:1967:EEI

[MI67] W. L. Minckley and Murray Itzkowitz. Ecology and effects of intestinal gas accumulation in a pupfish (genus Cyprinodon). Transactions of the American Fisheries Society, 96(2):216–218, April 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miller:1962:ORC

[Mil62] Jack G. Miller. Occurrence of ripe chain pickerel in the fall. Transactions of the American Fisheries Society, 91(3):323, July 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miller:1967:FTS

[Mil67] Robert Victor Miller. Food of the threadfin shad, Dorosoma petenense, in Lake Chicot, Arkansas. Transactions of the American Fisheries Society, 96(3):243–246, July 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Minckley:1962:SFJ

[Min62] W. L. Minckley. Spring foods of juvenile blue catfish from the Ohio River. Transactions of the American Fisheries Society, 91 (1):95, January 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mcerlean:1968:CSA

[MK68] Andrew J. Mcerlean and Victor S. Kennedy. Comparison of some anesthetic properties of benzocaine and MS-222. Transactions of the American Fisheries Society, 97(4):496–498, October 1968. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Minckley:1969:LHP

[MK69] C. O. Minckley and Harold E. Klaassen. Life history of the plains killifish, Fundulus kansae (Garman), in the Smoky Hill River, Kansas. Transactions of the American Fisheries Society, 98(3): 460–465, July 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moyle:1960:CLA

[ML60] John B. Moyle and Richard Lound. Confidence limits associated with means and medians of series of net catches. Transactions of the American Fisheries Society, 89(1):53–58, January 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Morris:1968:EMS

[MLRW68] Larry A. Morris, Ralph N. Langemeier, Thomas R. Russell, and Arthur Witt, Jr. Effects of main stem impoundments and channelization upon the limnology of the Missouri River, Nebraska. Transactions of the American Fisheries Society, 97(4):380–388, October 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCaig:1960:GES

[MM60] Robert S. McCaig and James W. Mullan. Growth of eight species of fishes in Quabbin Reservoir, Massachusetts, in relation to age of reservoir and introduction of smelt. Transactions of the American Fisheries Society, 89(1):27–31, January 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Monan:1969:DDD

[MMPS69] Gerald E. Monan, Robert J. McConnell, John R. Pugh, and Jim R. Smith. Distribution of debris and downstream-migrating salmon in the Snake River above Brownlee Reservoir. Transactions of the American Fisheries Society, 98(2):239–244, April 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Muir:1965:MMS

[MNB65] B. S. Muir, G. J. Nelson, and K. W. Bridges. A method for measuring swimming speed in oxygen consumption studies on the aholehole Kuhlia sandvicensis. Transactions of the American Fisheries Society, 94(4):378–382, October 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Muncy:1963:TTI

[MO63] Robert J. Muncy and Abe D. Oliver, Jr. Toxicity of ten insecticides to the red crawfish, Procambarus clarki (Girard). Transactions of the American Fisheries Society, 92(4):428–431, October 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moe:1967:PSM

[Moe67] Martin A. Moe, Jr. Prolonged survival and migration of three tagged reef fishes in the Gulf of Mexico. Transactions of the American Fisheries Society, 96(2):228–229, April 1967. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Momot:1965:FHB

[Mom65] Walter T. Momot. Food habits of the brook trout in West Lost Lake. Transactions of the American Fisheries Society, 94(2):188– 191, April 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Momot:1967:EBT

[Mom67] Walter T. Momot. Effects of brook trout predation on a crayfish population. Transactions of the American Fisheries Society, 96(2): 202–209, April 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moody:1960:RAL

[Moo60] Harold L. Moody. Recaptures of adult largemouth bass from the St. Johns River, Florida. Transactions of the American Fisheries Society, 89(3):295–300, July 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mount:1961:DSC

[Mou61] Donald I. Mount. Development of a system for controlling dissolved-oxygen content of water. Transactions of the American Fisheries Society, 90(3):323–327, July 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mount:1964:AIS

[Mou64a] Donald I. Mount. Additional information on a system for controlling the dissolved oxygen content of water. Transactions of the American Fisheries Society, 93(1):100–103, January 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mount:1964:ATZ

[Mou64b] Donald I. Mount. An autopsy technique for zinc-caused fish mortality. Transactions of the American Fisheries Society, 93(2):174–182, April 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Massmann:1960:DYA

[MP60] William H. Massmann and Anthony L. Pacheco. Disappearance of young Atlantic croakers from the York River, Virginia. Transactions of the American Fisheries Society, 89(2):154–159, April 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McErlean:1961:TDP

[MP61] Andrew J. McErlean and Ronald C. Phillips. A technique for direct photography of translucent objects. Transactions of the American Fisheries Society, 90(4):500–501, October 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Monan:1967:ECE

[MPS67] Gerald E. Monan, John R. Pugh, and Jim R. Smith. Efficiency of a combined electrode and louver array in guiding juvenile steelhead trout (Salmo gairdneri). Transactions of the American Fisheries Society, 96(4):422–423, October 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mueller:1962:EHM

[MR62] John W. Mueller and Louis C. Rockett. Effect of harvest, migration, and stocking on rainbow trout spawning potential in a Wyoming lake. Transactions of the American Fisheries Society, 91(1):63–68, January 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Macphee:1969:CSL

[MR69] Craig Macphee and Richard Ruelle. A chemical selectively lethal to squawfish (Ptychocheilus oregonensis and p. umpquae). Transactions of the American Fisheries Society, 98(4):676–684, October 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mraz:1964:AGR

[Mra64] Donald Mraz. Age and growth of the round whitefish in Lake Michigan. Transactions of the American Fisheries Society, 93(1): 46–52, January 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McFadden:1964:SSO

[MRC64] James T. McFadden, James R. Ryckman, and Gerald P. Cooper. A survey of some opinions of Michigan sport fishermen. Transactions of the American Fisheries Society, 93(2):183–193, April 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Moss:1961:DOR

[MS61] D. D. Moss and D. C. Scott. Dissolved-oxygen requirements of three species of fish. Transactions of the American Fisheries Society, 90 (4):377–393, October 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Meehan:1962:SDM

[MS62] William R. Meehan and Donald B. Siniff. A study of the downstream migrations of anadromous fishes in the Taku River, Alaska. Transactions of the American Fisheries Society, 91(4):399–407, October 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McErlean:1964:ASS

[MS64] Andrew J. McErlean and C. Lavett Smith. The age of sexual succession in the protogynous hermaphrodite Mycteroperca microlepis. Transactions of the American Fisheries Society, 93(3): 301–302, July 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Macleod:1966:EPF

[MS66] John Cameron Macleod and Lloyd L. Smith, Jr. Effect of pulpwood fiber on oxygen consumption and swimming endurance of the fathead minnow, Pimephales promelas. Transactions of the American Fisheries Society, 95(1):71–84, January 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Mount:1967:MEA

[MS67] Donald I. Mount and Charles E. Stephan. A method for establishing acceptable toxicant limits for fish — malathion and the butoxyethanol ester of 2,4-D. Transactions of the American Fisheries Society, 96(2):185–193, April 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Maloney:1962:LBP

[MSS62] J. E. Maloney, D. R. Schupp, and W. J. Scidmore. Largemouth bass population and harvest, Gladstone Lake, Crow Wing County, Minnesota. Transactions of the American Fisheries Society, 91(1): 42–52, January 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

McCarraher:1968:SEO

[MT68] D. B. McCarraher and Robert Thomas. Some ecological observations on the fathead minnow, Pimephales promelas, in the alkaline waters of Nebraska. Transactions of the American Fisheries Society, 97(1):52–55, January 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Murphy:1962:EMD

[Mur62] Garth I. Murphy. Effect of mixing depth and turbidity on the productivity of fresh-water impoundments. Transactions of the American Fisheries Society, 91(1):69–76, January 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Murray:1968:ENC

[Mur68] A. R. Murray. Estuarine net counting fence for trapping Atlantic salmon. Transactions of the American Fisheries Society, 97(3): 282–286, July 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Miller:1966:WTD

[MZC66] C. W. Miller, B. M. Zuckerman, and A. J. Charig. Water translocation of diazinon-c^{14} and parathion-s^{35} off a model cranberry bog and subsequent occurrence in fish and mussels. Transactions of the American Fisheries Society, 95(4):345–349, October 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nakamura:1965:FFH

[Nak65] Eugene L. Nakamura. Food and feeding habits of skipjack tuna (Katsuwonus pelamis) from the Marquesas and Tuamotu Islands. Transactions of the American Fisheries Society, 94(3):236–242, July 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Navarre:1962:NSH

[Nav62] Richard J. Navarre. A new stream habitat improvement structure in New Mexico. Transactions of the American Fisheries Society, 91(2):228–229, April 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Needham:1965:ENP

[NB65] P. R. Needham and R. J. Behnke. The effect of nematode (Philonema) and cestode (Diphyllobothrium) parasites in rainbow trout of Tebay Lake, Alaska. Transactions of the American Fisheries Society, 94(2):184–186, April 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Narver:1965:EFD

[ND65] David W. Narver and Michael L. Dahlberg. Estuarine food of Dolly Varden at chignik, Alaska. Transactions of the American Fisheries Society, 94(4):405–408, October 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nelson:1968:LHB

[Nel68a] Joseph S. Nelson. Life history of the brook silverside, Labidesthes sicculus, in Crooked Lake, Indiana. Transactions of the American Fisheries Society, 97(3):293–296, July 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nelson:1968:ELC

[Nel68b] William R. Nelson. Embryo and larval characteristics of sauger, walleye, and their reciprocal hybrids. Transactions of the American Fisheries Society, 97(2):167–174, April 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nelson:1968:REL

[Nel68c] William R. Nelson. Reproduction and early life history of sauger, Stizostedion canadense, in Lewis and Clark Lake. Transactions of the American Fisheries Society, 97(2):159–166, April 1968. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Niggol:1969:TMA

[Nig69] Karl Niggol. Thermal marking of adult Chinook salmon, Oncorhynchus tshawytscha. Transactions of the American Fisheries Society, 98(2):331–332, April 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nilsson:1963:IBT

[Nil63] Nils-Arvid Nilsson. Interaction between trout and char in Scandinavia. Transactions of the American Fisheries Society, 92(3): 276–285, July 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Norden:1967:AGF

[Nor67] Carroll R. Norden. Age, growth and fecundity of the alewife, Alosa pseudoharengus (Wilson), in Lake Michigan. Transactions of the American Fisheries Society, 96(4):387–393, October 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Normandeau:1969:LHE

[Nor69a] Donald A. Normandeau. Life history and ecology of the round whitefish Prosopium cylindraceum (Pallas), of Newfound Lake, Bristol, New Hampshire. Transactions of the American Fisheries Society, 98(1):7–13, January 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Norris:1969:DGF

[Nor69b] David O. Norris. Depression of growth following radiothyroidectomy of larval Chinook salmon and steelhead trout. Transactions of the American Fisheries Society, 98(1):104–106, January 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Neuhold:1960:ESF

[NS60] John M. Neuhold and William F. Sigler. Effects of sodium fluoride on carp and rainbow trout. Transactions of the American Fisheries Society, 89(4):358–370, October 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Netsch:1962:CLH

[NW62] Lt. Norval F. Netsch and Arthur Witt, Jr. Contributions to the life history of the longnose gar, (Lepisosteus osseus) in Missouri. Transactions of the American Fisheries Society, 91(3):251–262, July 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Northcote:1963:UCS

[NW63] T. G. Northcote and D. W. Wilkie. Underwater census of stream fish populations. Transactions of the American Fisheries Society, 92(2):146–151, April 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Nakamura:1961:IGT

[NY61] Eugene L. Nakamura and Heeny S. H. Yuen. Incidence of the giant trematode, hirudinella marina garcin, in skipjack tuna, Euthynnus pelamis (Linnaeus), from marquesan and Hawaiian waters. Transactions of the American Fisheries Society, 90(4):419–423, October 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Oviatt:1968:JLS

[OG68] Candace Oviatt and George W. Gray, Jr. Juvenile lookdowns, Selene vomer, in Wickford Cove, Narragansett Bay, Rhode Island. Transactions of the American Fisheries Society, 97(1):64, January 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Olla:1967:LEA

[OMK67] Bori L. Olla, Warren W. Marchioni, and Harvey M. Katz. A large experimental aquarium system for marine pelagic fishes. Transactions of the American Fisheries Society, 96(2):143–150, April 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Orcutt:1968:CST

[OPA68] Donald R. Orcutt, Ben R. Pulliam, and Arthur Arp. Characteristics of steelhead trout redds in Idaho streams. Transactions of the American Fisheries Society, 97(1):42–45, January 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

OConnor:1962:DAS

[OS62a] Joel S. O’Connor and Saul B. Saila. A developing aquatic sciences information retrieval system. Transactions of the American Fisheries Society, 91(2):151–154, April 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Olson:1962:HBS

[OS62b] Donald E. Olson and Warren J. Scidmore. Homing behavior of spawning walleyes. Transactions of the American Fisheries Society, 91(4):355–361, October 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Olson:1963:HTS

[OS63] Donald E. Olson and Warren J. Scidmore. Homing tendency of spawning white suckers in Many Point Lake, Minnesota. Transactions of the American Fisheries Society, 92(1):13–16, January 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Oglesby:1963:SBC

[OW63] Ray T. Oglesby and Charles M. Weiss. Some biochemical constituents of fish tissue, with special reference to brains. Transactions of the American Fisheries Society, 92(2):168–170, April 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Olla:1969:BWF

[OWW69] Bori L. Olla, Robert Wicklund, and Stuart Wilk. Behavior of winter flounder in a natural habitat. Transactions of the American Fisheries Society, 98(4):717–720, October 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Park:1969:SCD

[Par69] Donn L. Park. Seasonal changes in downstream migration of agegroup 0 Chinook salmon in the Upper Columbia River. Transactions of the American Fisheries Society, 98(2):315–317, April 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Patten:1962:CPU

[Pat62] Benjamin G. Patten. Cottid predation upon salmon fry in a Washington stream. Transactions of the American Fisheries Society, 91 (4):427–429, October 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Patriarche:1968:PTE

[Pat68a] Mercer H. Patriarche. Production and theoretical equilibrium yields for the bluegill (Lepomis macrochirus) in two Michigan lakes. Transactions of the American Fisheries Society, 97(3):242–251, July 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Patriarche:1968:REF

[Pat68b] Mercer H. Patriarche. Rate of escape of fish from trap nets. Transactions of the American Fisheries Society, 97(1):59–61, January 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Patten:1969:ESA

[Pat69] Bernard C. Patten. Ecological systems analysis and fisheries science. Transactions of the American Fisheries Society, 98(3):570– 581, July 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Paulik:1962:UCR

[Pau62] G. J. Paulik. Use of the Chapman — robson survival estimate for single- and multi- release tagging experiments. Transactions of the American Fisheries Society, 91(1):95–98, January 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Paulik:1969:CSM

[Pau69] G. J. Paulik. Computer simulation models for fisheries research, management, and teaching. Transactions of the American Fisheries Society, 98(3):551–559, July 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pereyra:1967:TCS

[PB67a] Walter T. Pereyra and Dina Barzel. Trawl catch summarization program, IBM 7090/7094, Fortran IV. Transactions of the American Fisheries Society, 96(3):363–364, July 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Powers:1967:EFG

[PB67b] James E. Powers and Anne LaBastille Bowes. Elimination of fish in the Giant Grebe Refuge, Lake Atitlan, Guatemala, using the fish toxicant, Antimycin. Transactions of the American Fisheries Society, 96(2):210–213, April 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Phillips:1969:OTS

[PBK69] Philip J. Phillips, W. David Burke, and Elizabeth J. Keener. Observations on the trophic significance of jellyfishes in Mississippi Sound with quantitative data on the associative behavior of small fishes with Medusae. Transactions of the American Fisheries Society, 98(4):703–712, October 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Phillips:1966:IMR

[PC66] Robert W. Phillips and Errol W. Claire. Intragravel movement of the reticulate sculpin, Cottus perplexus and its potential as a predator on salmonid embryos. Transactions of the American Fisheries Society, 95(2):210–212, April 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Power:1967:WRM

[PC67] G. Power and J. R. Colman. Winter rotenoning as a means of collecting and studying trout in an ice and snow covered Canadian stream. Transactions of the American Fisheries Society, 96(2): 222–223, April 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pierce:1967:IMM

[PD67] Frances Pierce and C. J. Dicostanzo. Instantaneous mortality and maturation schedules IBM 1620, Fortran II. Transactions of the American Fisheries Society, 96(3):366, July 1967. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pearcy:1962:TLF

[Pea62] William G. Pearcy. A tail-less flounder. Transactions of the American Fisheries Society, 91(2):233–234, April 1962. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pereyra:1965:CLW

[Per65] Walter T. Pereyra. Carapace length-width relationship for male and female Tanner crab. Transactions of the American Fisheries Society, 94(3):269–270, July 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Paulik:1964:AGB

[PG64] G. J. Paulik and Lawrence E. Gales. Allometric growth and the Beverton and Holt yield equation. Transactions of the American Fisheries Society, 93(4):369–381, October 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Paulik:1965:WLR

[PG65] G. J. Paulik and Lawrence E. Gales. Weighted linear regression for two variables, IBM 709, Fortran II. Transactions of the American Fisheries Society, 94(2):196, April 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pippy:1969:RRP

[PH69] John H. C. Pippy and Gerard M. Hare. Relationship of river pollution to bacterial infection in salmon (Salmo salar) and suckers (Catostomus commersoni). Transactions of the American Fisheries Society, 98(4):685–690, October 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Phillips:1969:AFE

[Phi69] Gary L. Phillips. Accuracy of fecundity estimates for the minnow Chrosomus erythrogaster (Cyprinidae). Transactions of the American Fisheries Society, 98(3):524–526, July 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pickering:1962:TOP

[PHL62] Q. H. Pickering, C. Henderson, and A. E. Lemke. The toxicity of organic phosphorus insecticides to different species of warmwater fishes. Transactions of the American Fisheries Society, 91(2):175– 184, April 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Peterson:1966:NCT

[PJS^{+}66] Durey H. Peterson, Herbert K. Jäger, George M. Savage, George N. Washburn, and Harry Westers. Natural coloration of trout using xanthophylls. Transactions of the American Fisheries Society, 95 (4):408–414, October 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Poole:1963:SSU

[PKW63] Richard L. Poole, Donald R. Kuester, and Arthur Witt, Jr. The status of spearfishing and the use of SCUBA in fish management programs in the freshwaters of the United States. Transactions of the American Fisheries Society, 92(1):30–33, January 1963. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Phinney:1967:MMY

[PMD67] Duane E. Phinney, Denny M. Miller, and Michael L. Dahlberg. Mass-marking young salmonids with fluorescent pigment. Transactions of the American Fisheries Society, 96(2):157–162, April 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Podoliak:1961:RBW

[Pod61] Henry A. Podoliak. Relation between water temperature and metabolism of dietary phosphorus by fingerling brook trout. Transactions of the American Fisheries Society, 90(4):398–403, October 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Poole:1969:SWF

[Poo69] John C. Poole. A study of winter flounder mortality rates in Great South Bay, New York. Transactions of the American Fisheries Society, 98(4):611–616, October 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Powell:1964:FMM

[Pow64a] Guy C. Powell. Fishing mortality and movements of adult male king crabs, Paralithodes camtschatica (Tilesius), released seaward from Kodiak Island, Alaska. Transactions of the American Fisheries Society, 93(3):295–300, July 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Power:1964:TPS

[Pow64b] G. Power. A technique for preparing scale smears. Transactions of the American Fisheries Society, 93(2):201–202, April 1964. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Powell:1965:TKC

[Pow65] Guy C. Powell. Tagged king crab recaptured six years after release in the North Pacific. Transactions of the American Fisheries Society, 94(1):95, January 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Patten:1969:RBN

[PR69] B. G. Patten and D. T. Rodman. Reproductive behavior of northern squawfish, Ptychocheilus oregonensis. Transactions of the American Fisheries Society, 98(1):108–111, January 1969. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Priegel:1963:DSG

[Pri63a] Gordon R. Priegel. Dispersal of the shortnose gar, Lepisosteus platostomus, into the Great Lakes drainage. Transactions of the American Fisheries Society, 92(2):178, April 1963. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Priegel:1963:FWS

[Pri63b] Gordon R. Priegel. Food of walleye and sauger in Lake Winnebago, Wisconsin. Transactions of the American Fisheries Society, 92(3): 312–313, July 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Priegel:1964:ESD

[Pri64] Gordon R. Priegel. Early scale development in the walleye. Transactions of the American Fisheries Society, 93(2):199–200, April 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Priegel:1966:ESD

[Pri66] Gordon R. Priegel. Early scale development in the freshwater drum, Aplodinotus grunniens Rafinesque. Transactions of the American Fisheries Society, 95(4):434–436, October 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Priegel:1967:FFD

[Pri67] Gordon R. Priegel. Food of the freshwater drum, Aplodinotus grunniens, in Lake Winnebago, Wisconsin. Transactions of the American Fisheries Society, 96(2):218–220, April 1967. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Priegel:1969:ARG

[Pri69a] Gordon R. Priegel. Age and rate of growth of the freshwater drum in Lake Winnebago, Wisconsin. Transactions of the American Fisheries Society, 98(1):116–118, January 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Priegel:1969:FGY

[Pri69b] Gordon R. Priegel. Food and growth of young walleyes in Lake Winnebago, Wisconsin. Transactions of the American Fisheries Society, 98(1):121–124, January 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Purkett:1961:RED

[Pur61] Charles A. Purkett, Jr. Reproduction and early development of the paddlefish. Transactions of the American Fisheries Society, 90(2): 125–129, April 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Purkett:1963:PFO

[Pur63] Charles A. Purkett, Jr. The paddlefish fishery of the Osage River and the Lake of the Ozarks, Missouri. Transactions of the American Fisheries Society, 92(3):239–244, July 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Parisot:1963:NGH

[PYB63] Thomas J. Parisot, William T. Yasutake, and Vernon Bressler. A new geographic and host record for infectious pancreatic necrosis. Transactions of the American Fisheries Society, 92(1):63–66, January 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Pycha:1961:RCW

[Pyc61] Richard L. Pycha. Recent changes in the walleye fishery of Northern Green Bay and history of the 1943 year class. Transactions of the American Fisheries Society, 90(4):475–488, October 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Perlmutter:1968:UFE

[PYS68] Alfred Perlmutter, Man-Lim Yu, and Robert E. Servis. Use of fish eggs for direct assay of chromatographically separated components on silica gel plates. Transactions of the American Fisheries Society, 97(2):204–207, April 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Qadri:1961:FDL

[Qad61] S. U. Qadri. Food and distribution of lake whitefish in Lac la Ronge, Saskatchewan. Transactions of the American Fisheries Society, 90 (3):303–307, July 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rahrer:1965:AGM

[Rah65] Jerold F. Rahrer. Age, growth, maturity, and fecundity of humper lake trout, Isle Royale, Lake Superior. Transactions of the American Fisheries Society, 94(1):75–83, January 1965. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rahrer:1967:GLT

[Rah67] Jerold F. Rahrer. Growth of lake trout in Lake Superior before the maximum abundance of sea lampreys. Transactions of the American Fisheries Society, 96(3):268–277, July 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rahrer:1968:MAL

[Rah68] Jerold F. Rahrer. Movements of adult lake trout in Lake Superior. Transactions of the American Fisheries Society, 97(4):481–484, October 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Raymond:1968:MRY

[Ray68] Howard L. Raymond. Migration rates of yearling Chinook salmon in relation to flows and impoundments in the Columbia and Snake Rivers. Transactions of the American Fisheries Society, 97(4):356– 359, October 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Raymond:1969:EJD

[Ray69] Howard L. Raymond. Effect of John Day Reservoir on the migration rate of juvenile Chinook salmon in the Columbia River. Transactions of the American Fisheries Society, 98(3):513–514, July 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Randall:1960:OEE

[RB60] John E. Randall and Vernon E. Brock. Observations on the ecology of epinepheline and lutjanid fishes of the Society Islands, with emphasis on food habits. Transactions of the American Fisheries Society, 89(1):9–16, January 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Robson:1961:CCM

[RC61] D. S. Robson and D. G. Chapman. Catch curves and mortality rates. Transactions of the American Fisheries Society, 90(2):181– 189, April 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rupp:1965:SCF

[RD65] Robert S. Rupp and Stuart E. Deroche. Standing crops of fishes in three small lakes compared with C^{14} estimates of net primary productivity. Transactions of the American Fisheries Society, 94 (1):9–25, January 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reed:1967:AGP

[Ree67a] Roger J. Reed. Age and growth of Prince of Wales, Alaska, Dolly Varden, Salvelinus malma (Walbaum), and rainbow trout, Salmo gairdneri Richardson. Transactions of the American Fisheries Society, 96(2):223–224, April 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reed:1967:OFA

[Ree67b] Roger J. Reed. Observation of fishes associated with spawning salmon. Transactions of the American Fisheries Society, 96(1): 62–67, January 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reed:1969:ASF

[Ree69] James R. Reed. Alarm substances and fright reaction in some fishes from the Southeastern United States. Transactions of the American Fisheries Society, 98(4):664–668, October 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Regier:1962:EMC

[Reg62a] Henry A. Regier. On estimating mortality coefficients in exploited fish populations, given two censuses. Transactions of the American Fisheries Society, 91(3):283–294, July 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Regier:1962:VSM

[Reg62b] Henry A. Regier. Validation of the scale method for estimating age and growth of bluegills. Transactions of the American Fisheries Society, 91(4):362–374, October 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Reimers:1963:BCW

[Rei63] Norman Reimers. Body condition, water temperature, and overwinter survival of hatchery-reared trout in Convict Creek, California. Transactions of the American Fisheries Society, 92(1):39–46, January 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Renfro:1963:GBM

[Ren63] William C. Renfro. Gas-bubble mortality of fishes in Galveston Bay, Texas. Transactions of the American Fisheries Society, 92 (3):320–322, July 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rose:1968:AGD

[RH68] Curt D. Rose and W. W. Hassler. Age and growth of the Dolphin, Coryphaena hippurus (Linnaeus), in North Carolina waters. Transactions of the American Fisheries Society, 97(3):271–276, July 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rose:1969:AST

[RH69] Curt D. Rose and W. W. Hassler. Application of survey techniques to the Dolphin, Coryphaena hippurus, fishery of North Carolina. Transactions of the American Fisheries Society, 98(1):94–103, January 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ricker:1962:CRS

[Ric62] W. E. Ricker. Connecticut River shad. Transactions of the American Fisheries Society, 91(1):98–99, January 1962. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Richards:1967:AGF

[Ric67] C. E. Richards. Age, growth and fecundity of the cobia, Rachycentron canadum, from Chesapeake Bay and adjacent mid-Atlantic waters. Transactions of the American Fisheries Society, 96(3): 343–350, July 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ridenhour:1960:DPS

[Rid60] Richard L. Ridenhour. Development of a program to sample young fish in a lake. Transactions of the American Fisheries Society, 89(2):185–192, April 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Richards:1966:CPJ

[RM66] Sarah W. Richards and A. Marshall McBean. Comparison of postlarvae and juveniles of Fundulus heteroclitus and Fundulus majalis (Pisces: Cyprinodontidae). Transactions of the American Fisheries Society, 95(2):218–226, April 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ross:1960:BDY

[RNBH60] A. John Ross, Phyllis R. Nordstrom, Jack E. Bailey, and John R. Heaton. A bacterial disease of yellow perch (Peres flavescens). Transactions of the American Fisheries Society, 89(3):310–312, July 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Robel:1962:WIC

[Rob62] Robert J. Robel. Weight increases of carp in populations of different densities. Transactions of the American Fisheries Society, 91(2): 234–236, April 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Roberson:1967:OCS

[Rob67] Kenneth Roberson. An occurrence of Chinook salmon beach spawning in Lake Washington. Transactions of the American Fisheries Society, 96(4):423–424, October 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Roessler:1965:AVF

[Roe65] Martin Roessler. An analysis of the variability of fish populations taken by otter trawl in Biscayne Bay, Florida. Transactions of the American Fisheries Society, 94(4):311–318, October 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Roe:1966:PCN

[Roe66] Richard B. Roe. Potentially commercial nephropsids from the Western Atlantic. Transactions of the American Fisheries Society, 95(1):92–98, January 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rolan:1968:ECH

[Rol68] Robert G. Rolan. Electrophoretic comparison of human and catfish plasma proteins. Transactions of the American Fisheries Society, 97(4):455–459, October 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Roos:1960:PYC

[Roo60] John F. Roos. Predation of young Coho salmon on sockeye salmon fry at chignik, Alaska. Transactions of the American Fisheries Society, 89(4):377–379, October 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ropes:1961:LHC

[Rop61] John W. Ropes. Longevity of the horseshoe crab, Limulus polyphemus (L). Transactions of the American Fisheries Society, 90(1): 79–80, January 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rothschild:1961:PSE

[Rot61] Brian J. Rothschild. Production and survival of eggs of the American smelt, Osmerus mordax (Mitchill), in Maine. Transactions of the American Fisheries Society, 90(1):42–48, January 1961. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rothschild:1963:CSM

[Rot63] Brian J. Rothschild. A critique of the scale method for determining the age of the alewife, Alosa pseudoharengus (Wilson). Transactions of the American Fisheries Society, 92(4):409–413, October 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Richards:1969:MUR

[RP69] William J. Richards and Barbara J. Palko. Methods used to rear the thread herring, Opisthonema oglinum, from fertilized eggs. Transactions of the American Fisheries Society, 98(3):527–529, July 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ross:1960:VLD

[RPR60] A. J. Ross, J. Pelnar, and R. R. Rucker. A virus-like disease of Chinook salmon. Transactions of the American Fisheries Society, 89(2):160–163, April 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Richards:1968:RAS

[RPS68] William J. Richards, Barbara J. Palko, and Edwin L. Scott. A research aquarium suitable for shipboard use. Transactions of the American Fisheries Society, 97(3):286–287, July 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Robson:1964:SSP

[RR64] D. S. Robson and H. A. Regier. Sample size in Petersen mark– recapture experiments. Transactions of the American Fisheries Society, 93(3):215–226, July 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Robson:1966:ETL

[RR66] D. S. Robson and H. A. Regier. Estimates of tag loss from recoveries of fish tagged and permanently marked. Transactions of the American Fisheries Society, 95(1):56–59, January 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rickey:1963:CLW

[RS63] Roy Rickey and W. L. Sheridan. Carapace length-width relationship for the king crab. Transactions of the American Fisheries Society, 92(3):318–320, July 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Radtke:1967:HCT

[RT67] Larry D. Radtke and Jerry L. Turner. High concentrations of total dissolved solids block spawning migration of striped bass, Roccus saxatilis, in the San Joaquin River, California. Transactions of the American Fisheries Society, 96(4):405–407, October 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rupp:1961:MPF

[Rup61] Robert S. Rupp. Measurement of potential fishing quality. Transactions of the American Fisheries Society, 90(2):165–169, April 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Rupp:1965:SSS

[Rup65] Robert S. Rupp. Shore-spawning and survival of eggs of the American smelt. Transactions of the American Fisheries Society, 94(2): 160–168, April 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Radulescu:1968:THE

[RVIS68] I. Radulescu, D. G. Vasiliu, E. Ilie, and S. F. Snieszko. Thyroid hyperplasia of the eastern brook trout, Salvelinus fontinalis, in Romania. Transactions of the American Fisheries Society, 97(4): 486–488, October 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Roithmayr:1963:SOB

[RW63] Charles M. Roithmayr and Richard A. Waller. Seasonal occurrence of Brevoortia patronus in the Northern Gulf of Mexico. Transactions of the American Fisheries Society, 92(3):301–302, July 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ryder:1964:CCO

[Ryd64] R. A. Ryder. Chemical characteristics of Ontario lakes as related to glacial history. Transactions of the American Fisheries Society, 93 (3):260–268, July 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ryder:1965:MEP

[Ryd65] R. A. Ryder. A method for estimating the potential fish production of north-temperate lakes. Transactions of the American Fisheries Society, 94(3):214–218, July 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ross:1963:HRT

[RYW63] A. J. Ross, W. T. Yasutake, and G. R. White. Hermaphroditism in rainbow trout. Transactions of the American Fisheries Society, 92 (3):313–315, July 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shetter:1962:EFO

[SA62] David S. Shetter and Gaylord R. Alexander. Effects of a flies-only restriction on angling and on fall trout populations in Hunt Creek, Montmorency County, Michigan. Transactions of the American Fisheries Society, 91(3):295–302, July 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shetter:1965:RAU

[SA65] David S. Shetter and Gaylord R. Alexander. Results of angling under special and normal trout fishing regulations in a Michigan trout stream. Transactions of the American Fisheries Society, 94 (3):219–226, July 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shetter:1966:ATP

[SA66] David S. Shetter and Gaylord R. Alexander. Angling and trout populations on the North Branch of the Au Sable River, Crawford and Otsego Counties, Michigan, under special and normal regulations, 1958–63. Transactions of the American Fisheries Society, 95(1):85–91, January 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Saila:1962:PHB

[Sai62] Saul B. Saila. Proposed hurricane barriers related to winter flounder movements in Narragansett Bay. Transactions of the American Fisheries Society, 91(2):189–195, April 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sakuda:1966:CAO

[Sak66a] Henry M. Sakuda. Condition of American oyster, Crassostrea virginica, in West Loch, Pearl Harbor, Hawaii. Transactions of the American Fisheries Society, 95(4):426–429, October 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sakuda:1966:RCA

[Sak66b] Henry M. Sakuda. Reproductive cycle of American oyster, Crassostrea virginica in West Loch, Pearl Harbor, Hawaii. Transactions of the American Fisheries Society, 95(2):216–218, April 1966. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sanders:1966:TSP

[SC66] Herman O. Sanders and Oliver B. Cope. Toxicities of several pesticides to two species of cladocerans. Transactions of the American Fisheries Society, 95(2):165–169, April 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stock:1969:SET

[SC69] Jay N. Stock and Oliver B. Cope. Some effects of TEPA, an insect chemosterilant, on the guppy, Poecilia reticulata. Transactions of the American Fisheries Society, 98(2):280–287, April 1969. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schultz:1960:EED

[Sch60] Edward E. Schultz. Establishment and early dispersal of a loach, Misgurnus anguillicaudatus (Cantor), in Michigan. Transactions of the American Fisheries Society, 89(4):376–377, October 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schwartz:1962:APH

[Sch62] Frank J. Schwartz. Artificial pike hybrids, Esox americanus vermiculatus × E. lucius. Transactions of the American Fisheries Society, 91(2):229–230, April 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schaefer:1965:PHS

[Sch65a] Milner B. Schaefer. The potential harvest of the sea. Transactions of the American Fisheries Society, 94(2):123–128, April 1965. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schofield:1965:WQR

[Sch65b] Carl L. Schofield, Jr. Water quality in relation to survival of brook trout, Salvelinus fontinalis (Mitchill). Transactions of the American Fisheries Society, 94(3):227–235, July 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schoumacher:1965:MWS

[Sch65c] Roger Schoumacher. Movement of walleye and sauger in the Upper Mississippi River. Transactions of the American Fisheries Society, 94(3):270–271, July 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schumacher:1966:SIM

[Sch66] Robert E. Schumacher. Successful introduction of Mysis relicta (Loven) into a Minnesota lake. Transactions of the American Fisheries Society, 95(2):216, April 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schmittou:1967:SRB

[Sch67] H. R. Schmittou. Sex ratios of bluegill in four populations. Transactions of the American Fisheries Society, 96(4):420–421, October 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schaefer:1968:MEE

[Sch68a] Milner B. Schaefer. Methods of estimating effects of fishing on fish populations. Transactions of the American Fisheries Society, 97 (3):231–241, July 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Scholl:1968:RDM

[Sch68b] Russell L. Scholl. A rapid decalcifying method for sectioning channel catfish pectoral spines. Transactions of the American Fisheries Society, 97(2):210–211, April 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schoumacher:1968:SOF

[Sch68c] Roger Schoumacher. Some observations on flathead catfish in the Mississippi River bordering Iowa. Transactions of the American Fisheries Society, 97(1):65–66, January 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schwartz:1963:AGF

[SD63] Frank J. Schwartz and Ben W. Dutcher. Age, growth, and food of the oyster toadfish near solomons, Maryland. Transactions of the American Fisheries Society, 92(2):170–173, April 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schumacher:1960:APS

[SE60] Robert E. Schumacher and Samuel Eddy. The appearance of pink salmon, Oncorhynchus gorbuscha (Walbaum), in Lake Superior. Transactions of the American Fisheries Society, 89(4):371–373, October 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Seal:1966:TCA

[Sea66] Hilary L. Seal. Testing for contagion in animal populations. Transactions of the American Fisheries Society, 95(4):436–437, October 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Springer:1963:ACT

[SF63] Victor G. Springer and John H. Finucane. The African cichlid, Tilapia heudeloti Dumeril, in the commercial fish catch of Florida. Transactions of the American Fisheries Society, 92(3):317–318, July 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Saila:1969:TGM

[SF69] S. B. Saila and J. M. Flowers. Toward a generalized model of fish migration. Transactions of the American Fisheries Society, 98(3): 582–588, July 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Saila:1969:FAL

[SFH69] S. B. Saila, J. M. Flowers, and J. T. Hughes. Fecundity of the American lobster, Homarus americanus. Transactions of the American Fisheries Society, 98(3):537–539, July 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Skulberg:1967:SRT

[SG67] Anton Skulberg and Magne Grande. Susceptibility of rainbow trout Salmo gairdneri Richardson, to Clostridium botulinum toxins. Transactions of the American Fisheries Society, 96(1):67–70, January 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schumacher:1962:TGP

[SH62a] Robert E. Schumacher and John G. Hale. Third generation pink salmon, Oncorhynchus gorbuscha (Walbaum), in Lake Superior. Transactions of the American Fisheries Society, 91(4):421–422, October 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smitherman:1962:APS

[SH62b] R. Oneal Smitherman and F. Eugene Hester. Artificial propagation of sunfishes, with meristic comparisons of three species of Lepomis and five of their hybrids. Transactions of the American Fisheries Society, 91(4):333–341, October 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stark:1967:ECC

[SH67] A. E. Stark and Mary C. Halden. Estimation of catch composition from sample surveys by stepwise application of raising factors, CDC 3200, Fortran IV. Transactions of the American Fisheries Society, 96(2):230, April 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stauffer:1969:MRS

[SH69] Thomas M. Stauffer and Martin J. Hansen. Mark retention, survival, and growth of jaw-tagged and fin-clipped rainbow trout. Transactions of the American Fisheries Society, 98(2):225–229, April 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shaw:1969:PPS

[Sha69a] W. N. Shaw. The past and present status of off-bottom oyster culture in North America. Transactions of the American Fisheries Society, 98(4):755–761, October 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shaw:1969:OST

[Sha69b] William N. Shaw. Oyster setting in two adjacent tributaries of Chesapeake Bay. Transactions of the American Fisheries Society, 98(2):309–314, April 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shetter:1960:RES

[She60] David S. Shetter. The relative effectiveness of salmon eggs and earthworms as bait for trout in the Sturgeon River, Michigan. Transactions of the American Fisheries Society, 89(3):280–286, July 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shetter:1961:SBT

[She61] David S. Shetter. Survival of brook trout from egg to fingerling stage in two Michigan trout streams. Transactions of the American Fisheries Society, 90(3):252–258, July 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shetter:1962:RAL

[She62] David S. Shetter. Recoveries by anglers of legal-sized trout stocked during different seasons of the year in Michigan streams. Transactions of the American Fisheries Society, 91(2):145–150, April 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shell:1963:ECD

[She63] E. W. Shell. Effects of changed diets on the growth of channel catfish. Transactions of the American Fisheries Society, 92(4): 432–434, October 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shetter:1967:EJT

[She67] David S. Shetter. Effects of jaw tags and fin excision upon the growth, survival, and exploitation of hatchery rainbow trout fingerlings in Michigan. Transactions of the American Fisheries Society, 96(4):394–399, October 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sheldon:1968:DGM

[She68a] Andrew L. Sheldon. Drift, growth, and mortality of juvenile sculpins in Sagehen Creek, California. Transactions of the American Fisheries Society, 97(4):495–496, October 1968. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shetter:1968:OMW

[She68b] David S. Shetter. Observations on movements of wild trout in two Michigan stream drainages. Transactions of the American Fisheries Society, 97(4):472–480, October 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shomura:1964:ETL

[Sho64] Richard S. Shomura. Effectiveness of tilapia as live bait for skipjack tuna fishing. Transactions of the American Fisheries Society, 93 (3):291–294, July 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Siefert:1965:ESD

[Sie65] Richard E. Siefert. Early scale development in the white crappie. Transactions of the American Fisheries Society, 94(2):182, April 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Siefert:1968:RBI

[Sie68] Richard E. Siefert. Reproductive behavior, incubation and mortality of eggs, and postlarval food selection in the white crappie. Transactions of the American Fisheries Society, 97(3):252–259, July 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Siefert:1969:CSW

[Sie69] Richard E. Siefert. Characteristics for separation of white and black crappie larvae. Transactions of the American Fisheries Society, 98(2):326–328, April 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Silliman:1969:ACS

[Sil69] Ralph P. Silliman. Analog computer simulation and catch forecasting in commercially fished populations. Transactions of the American Fisheries Society, 98(3):560–569, July 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1963:SWE

[SK63] Lloyd L. Smith, Jr. and Robert H. Kramer. Survival of walleye eggs in relation to wood fibers and sphaerotilus natans in the Rainy River, Minnesota. Transactions of the American Fisheries Society, 92(3):220–234, July 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1964:SSL

[SK64] Lloyd L. Smith, Jr. and Robert H. Kramer. The spottail shiner in Lower Red Lake, Minnesota. Transactions of the American Fisheries Society, 93(1):35–45, January 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1965:SWF

[SK65] Lloyd L. Smith, Jr. and Robert H. Kramer. Survival of walleye fingerlings in conifer groundwood fiber. Transactions of the American Fisheries Society, 94(4):402–404, October 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1966:LTE

[SKO66] Lloyd L. Smith, Jr., Robert H. Kramer, and Donovan M. Oseid. Long-term effects of conifer-groundwood paper fiber on walleyes. Transactions of the American Fisheries Society, 95(1):60–70, January 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Summerfelt:1968:WSE

[SL68] Robert C. Summerfelt and William M. Lewis. A water sampler employing a solenoid tripping mechanism. Transactions of the American Fisheries Society, 97(3):287–289, July 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Solon:1968:CFA

[SLN68] Joseph M. Solon, Jeffrey L. Lincer, and John H. Nair III. A continuous flow, automatic device for short-term toxicity experiments. Transactions of the American Fisheries Society, 97(4):501–502, October 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Springer:1961:TGB

[SM61] Victor G. Springer and Andrew J. McErlean. Tagging of great barracuda, Sphyraena barracuda (Walbaum). Transactions of the American Fisheries Society, 90(4):497–500, October 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Seaburg:1964:FHD

[SM64] Keith G. Seaburg and John B. Moyle. Feeding habits, digestive rates, and growth of some Minnesota warmwater fishes. Transactions of the American Fisheries Society, 93(3):269–285, July 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Swee:1966:RBC

[SM66] U. Boon Swee and Hugh R. McCrimmon. Reproductive biology of the carp, Cyprinus carpio L., in Lake St. Lawrence, Ontario. Transactions of the American Fisheries Society, 95(4):372–380, October 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Summerfelt:1969:ALH

[SM69] Robert C. Summerfelt and Charles O. Minckley. Aspects of the life history of the sand shiner, Notropis stramineus (Cope), in the Smoky Hill River, Kansas. Transactions of the American Fisheries Society, 98(3):444–453, July 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stewart:1967:ATI

[SMB67] Nelson E. Stewart, Raymond E. Millemann, and Wilbur P. Breese. Acute toxicity of the insecticide Sevin(R) and its hydrolytic product 1-naphthol to some marine organisms. Transactions of the American Fisheries Society, 96(1):25–30, January 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1964:SDC

[Smi64] Stanford H. Smith. Status of the deepwater cisco population of Lake Michigan. Transactions of the American Fisheries Society, 93(2):155–163, April 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1967:ELL

[Smi67] Allen J. Smith. The effect of the lamprey larvicide, 3trifluoromethyl-4-nitrophenol, on selected aquatic invertebrates. Transactions of the American Fisheries Society, 96(4):410–413, October 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Simon:1968:HOS

[SN68] Raymond C. Simon and Richard E. Noble. Hybridization in Oncorhynchus (Salmonidae). I. Viability and inheritance in artificial crosses of chum and pink salmon. Transactions of the American Fisheries Society, 97(2):109–118, April 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sokoloski:1969:SEE

[Sok69] Adam A. Sokoloski. Some elements of an evaluation of the effects of legal factors on the utilization of fishery resources. Transactions of the American Fisheries Society, 98(2):360–367, April 1969. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1961:FRC

[SP61] Lloyd L. Smith, Jr. and Richard L. Pycha. Factors related to commercial production of the walleye in Red Lakes, Minnesota. Transactions of the American Fisheries Society, 90(2):190–217, April 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shelton:1966:SES

[SP66a] J. M. Shelton and R. D. Pollock. Siltation and egg survival in incubation channels. Transactions of the American Fisheries Society, 95(2):183–187, April 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stober:1966:MPP

[SP66b] Q. J. Stober and W. R. Payne, Jr. A method for preparation of pesticide-free fish food from commercial fish food pellets. Transactions of the American Fisheries Society, 95(2):212–214, April 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sheri:1969:AFS

[SP69a] A. N. Sheri and G. Power. Annulus formation on scales of the white perch, Morone americanus (Gmelin), in the Bay of Quinte, Lake Ontario. Transactions of the American Fisheries Society, 98(2): 322–326, April 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Smith:1969:FSB

[SP69b] Philip W. Smith and Lawrence M. Page. The food of spotted base in streams of the Wabash River drainage. Transactions of the American Fisheries Society, 98(4):647–651, October 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Speece:1969:UTO

[Spe69] R. E. Speece. U-tube oxygenation for economical saturation of fish hatchery water. Transactions of the American Fisheries Society, 98(4):789–795, October 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Schoeneman:1961:MDM

[SPJ61] Dale E. Schoeneman, Richard T. Pressey, and Charles O. Junge, Jr. Mortalities of downstream migrant salmon at McNary Dam. Transactions of the American Fisheries Society, 90(1):58–72, January 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Saunders:1962:PAS

[SS62a] J. W. Saunders and M. W. Smith. Physical alteration of stream habitat to improve brook trout production. Transactions of the American Fisheries Society, 91(2):185–188, April 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Saunders:1962:TBT

[SS62b] J. W. Saunders and M. W. Smith. Transplantation of brook trout (Salvelinus fontinalis (Mitchill)) within a small stream system.

Transactions of the American Fisheries Society, 91(4):388–394, October 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sundararaj:1962:FSS

[SS62c] Bangalore I. Sundararaj and Royal D. Suttkus. Fecundity of the spotted seatrout, Cynoscion nebulosus (Cuvier), from Lake Borgne Area, Louisiana. Transactions of the American Fisheries Society, 91(1):84–88, January 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Swingle:1966:PAP

[SS66] Wayne E. Swingle and H. S. Swingle. Population analysis programs, IBM 1620, Fortran/Format. Transactions of the American Fisheries Society, 95(2):232–233, April 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Surber:1963:LSE

[ST63] Eugene W. Surber and Thomas O. Thatcher. Laboratory studies of the effects of alkyl benzine sulfonate (ABS) on aquatic invertebrates. Transactions of the American Fisheries Society, 92(2): 152–160, April 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stauffer:1962:DLL

[Sta62] Thomas M. Stauffer. Duration of larval life of sea lampreys in Carp Lake River, Michigan. Transactions of the American Fisheries Society, 91(4):422–423, October 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stevens:1966:ASQ

[Ste66] Joe B. Stevens. Angler success as a quality determinant of sport fishery recreational values. Transactions of the American Fisheries Society, 95(4):357–362, October 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stevens:1969:MEV

[Ste69] Joe B. Stevens. Measurement of economic values in sport fishing: an economist’s views on validity, usefulness, and propriety. Transactions of the American Fisheries Society, 98(2):352–357, April 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stickney:1967:AST

[Sti67] Alden P. Stickney. Aquarium susceptibility of tagged and untagged Atlantic herring to predation. Transactions of the American Fisheries Society, 96(3):359–361, July 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Stober:1969:UNS

[Sto69] Quentin J. Stober. Underwater noise spectra, fish sounds and response to low frequencies of cutthroat trout (Salmo clarki) with reference to orientation and homing in Yellowstone Lake. Transactions of the American Fisheries Society, 98(4):652–663, October 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Strawn:1961:GLB

[Str61] Kirk Strawn. Growth of largemouth bass fry at various temperatures. Transactions of the American Fisheries Society, 90(3):334– 335, July 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Sumner:1962:MGC

[Sum62] F. H. Sumner. Migration and growth of the coastal cutthroat trout in Tillamook County, Oregon. Transactions of the American Fisheries Society, 91(1):77–83, January 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Summerfelt:1964:NMP

[Sum64] Robert C. Summerfelt. A new microsporidian parasite from the golden shiner, Notemigonus crysoleucas. Transactions of the American Fisheries Society, 93(1):6–10, January 1964. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Summerfelt:1966:HSP

[Sum66] Robert C. Summerfelt. Homology of serum proteins of golden shiner (Notemigonus crysoleucas) and man. Transactions of the American Fisheries Society, 95(3):272–279, July 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Swarts:1969:BSS

[Swa69] William Swarts. Blood studies of some marine teleosts. Transactions of the American Fisheries Society, 98(2):328–331, April 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Silver:1963:DOR

[SWD63] Stuart J. Silver, Charles E. Warren, and Peter Doudoroff. Dissolved oxygen requirements of developing steelhead trout and Chinook salmon embryos at different water velocities. Transactions of the American Fisheries Society, 92(4):327–343, October 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Shumway:1964:IOC

[SWD64] Dean L. Shumway, Charles E. Warren, and Peter Doudoroff. Influence of oxygen concentration and water movement on the growth of steelhead trout and Coho salmon embryos. Transactions of the American Fisheries Society, 93(4):342–356, October 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Swedberg:1968:FGF

[Swe68] Donald V. Swedberg. Food and growth of the freshwater drum in Lewis and Clark Lake, South Dakota. Transactions of the American Fisheries Society, 97(4):442–447, October 1968. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Swingle:1960:CET

[Swi60] H. S. Swingle. Comparative evaluation of two tilapias as pondfishes in Alabama. Transactions of the American Fisheries Society, 89(2): 142–148, April 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Swift:1963:IOC

[Swi63] D. R. Swift. Influence of oxygen concentration on growth of brown trout, Salmo trutta L. Transactions of the American Fisheries Society, 92(3):300–301, July 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Swingle:1964:LWRa

[Swi64a] Wayne E. Swingle. Length–weight relationships I, IBM 1620, Fortran/Format. Transactions of the American Fisheries Society, 93 (3):318–319, July 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Swingle:1964:LWRb

[Swi64b] Wayne E. Swingle. Length–weight relationships II, IBM 1620, Fortran/Format. Transactions of the American Fisheries Society, 93 (3):319, July 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tagatz:1969:SRT

[Tag69] Marlin E. Tagatz. Some relations of temperature acclimation and salinity to thermal tolerance of the blue crab, Callinectes sapidus. Transactions of the American Fisheries Society, 98(4):713–716, October 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Talbot:1964:DBM

[Tal64] Gerald B. Talbot. Drift bottle modifications for air drops. Transactions of the American Fisheries Society, 93(2):203–204, April 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Talbot:1967:UCS

[Tal67] Gerald B. Talbot. Urinary calculi in striped marlin. Transactions of the American Fisheries Society, 96(3):355–357, July 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tanner:1960:SCA

[Tan60] Howard A. Tanner. Some consequences of adding fertilizer to five Michigan trout lakes. Transactions of the American Fisheries Society, 89(2):198–205, April 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tarrant:1964:REC

[Tar64] Robert M. Tarrant, Jr. Rate of extinction of a conditioned response in juvenile sockeye salmon. Transactions of the American Fisheries Society, 93(4):399–401, October 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tarrant:1966:TPE

[Tar66] Robert M. Tarrant, Jr. Threshold of perception of eugenol in juvenile sockeye salmon. Transactions of the American Fisheries Society, 95(1):112–115, January 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tarter:1969:SAR

[Tar69] Donald C. Tarter. Some aspects of reproduction in the western blacknose dace, Rhinichthys atratulus meleagris Agassiz, in Doe Run, Meade County, Kentucky. Transactions of the American Fisheries Society, 98(3):454–459, July 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Taub:1963:RML

[Tau63] Stephen H. Taub. Recovery of a marked lake trout after 21 years. Transactions of the American Fisheries Society, 92(4):432, October 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thomas:1969:EYS

[TBG69] Allan E. Thomas, Joe L. Banks, and Donald C. Greenland. Effect of yolk sac absorption on the swimming ability of fall Chinook salmon. Transactions of the American Fisheries Society, 98(3): 406–410, July 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thomas:1969:HCC

[TEB69] Allan E. Thomas, Joseph W. Elliott, and Joe L. Banks. Hematological and chemical characteristics associated with precocious male Chinook salmon fingerlings. Transactions of the American Fisheries Society, 98(1):23–26, January 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Teter:1960:BFL

[Tet60] Harold E. Teter. The bottom fauna of Lake Huron. Transactions of the American Fisheries Society, 89(2):193–197, April 1960. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thorsteinson:1962:HPP

[Tho62] Fredrik V. Thorsteinson. Herring predation on pink salmon fry in a Southeastern Alaska estuary. Transactions of the American Fisheries Society, 91(3):321–323, July 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tabb:1962:EHD

[TJ62] Durbin C. Tabb and Albert C. Jones. Effect of Hurricane Donna on the aquatic fauna of North Florida Bay. Transactions of the American Fisheries Society, 91(4):375–378, October 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Turner:1967:OPI

[TR67] William R. Turner and Richard B. Roe. Occurrence of the parasitic isopod Olencira praegustator in the yellowfin menhaden, Brevoortia smithi. Transactions of the American Fisheries Society, 96(3):357– 359, July 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Thompson:1967:PDV

[TT67] Richard B. Thompson and Dennis F. Tufts. Predation by Dolly Varden and northern squawfish on hatchery-reared sockeye salmon in Lake Wenatchee, Washington. Transactions of the American Fisheries Society, 96(4):424–427, October 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Turner:1960:SCF

[Tur60] William R. Turner. Standing crops of fishes in Kentucky farm ponds. Transactions of the American Fisheries Society, 89(4):333– 337, October 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Turner:1969:LHM

[Tur69] William R. Turner. Life history of menhadens in the Eastern Gulf of Mexico. Transactions of the American Fisheries Society, 98(2): 216–224, April 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Tabb:1969:RMA

[TYII69] D. C. Tabb, W. T. Yang, C. P. Idyll, and E. S. Iversen. Research in marine aquaculture at the Institute of Marine Sciences, University of Miami. Transactions of the American Fisheries Society, 98(4): 738–742, October 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Uzmann:1966:CMS

[UD66] Joseph R. Uzmann and John Douglas. Clinostomum marginatum in steelhead trout (Salmo gairdneri) and cutthroat trout (Salmo clarki) in a Western Washington lake. Transactions of the American Fisheries Society, 95(1):35–38, January 1966. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Uzmann:1965:EHJ

[UPH65] J. R. Uzmann, G. J. Paulik, and S. H. Hayduk. Experimental hexamitiasis in juvenile Coho salmon (Oncorhynchus kisutch) and steelhead trout (Salmo gairdneri). Transactions of the American Fisheries Society, 94(1):53–61, January 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanValin:1968:SEM

[VAE68] Charles C. Van Valin, Austin K. Andrews, and Lafayette L. Eller. Some effects of mirex on two warm-water fishes. Transactions of the American Fisheries Society, 97(2):185–196, April 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanOosten:1961:RAG

[Van61] John Van Oosten. Records, ages, and growth of the mooneye, Hiodon tergisus, of the Great Lakes. Transactions of the American Fisheries Society, 90(2):170–174, April 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VanMeir:1969:EEA

[Van69] Lawrence W. Van Meir. An economic evaluation of alternative management systems for commercial fisheries. Transactions of the American Fisheries Society, 98(2):347–350, April 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Vanderwalker:1962:DFB

[VC62] John G. Vanderwalker and Edward Chin. A device for feeding brine shrimp to fishes. Transactions of the American Fisheries Society, 91(2):230–231, April 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Verduin:1969:CRM

[Ver69] Jacob Verduin. Critique of research methods involving plastic bags in aquatic environments. Transactions of the American Fisheries Society, 98(2):335–336, April 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Vibert:1963:NEF

[Vib63] Richard Vibert. Neurophysiology of electric fishing. Transactions of the American Fisheries Society, 92(3):265–275, July 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Vincent:1960:SID

[Vin60] Robert E. Vincent. Some influences of domestication upon three stocks of brook trout (Salvelinus fontinalis Mitchill). Transactions of the American Fisheries Society, 89(1):35–52, January 1960. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Vanicek:1969:LHC

[VK69] C. David Vanicek and Robert H. Kramer. Life history of the Colorado squawfish, Ptychocheilus lucius, and the Colorado chub, Gila robusta, in the Green River in Dinosaur National Monument, 1964– 1966. Transactions of the American Fisheries Society, 98(2):193– 208, April 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

VonGeldern:1961:ADM

[Von61] C. E. Von Geldern, Jr. Application of the DeLury method in determining the angler harvest of stocked catchable-sized trout. Transactions of the American Fisheries Society, 90(3):259–263, July 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Voigtlander:1967:AGP

[VR67] Clyde W. Voigtlander and Alan C. Roochvarg. Age and growth program, Control Data 3600, Fortran 63–Fortran IV. Transactions of the American Fisheries Society, 96(3):364–366, July 1967. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wagner:1968:EST

[Wag68] Harry H. Wagner. Effect of stocking time on survival of steelhead trout, Salmo gairdnerii, in Oregon. Transactions of the American Fisheries Society, 97(4):374–379, October 1968. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wagner:1969:ILI

[Wag69a] Charles C. Wagner. Incidence of leech infestation on blacknose dace. Transactions of the American Fisheries Society, 98(1):115– 116, January 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wagner:1969:ESL

[Wag69b] Harry H. Wagner. Effect of stocking location of juvenile steelhead trout, Salmo gairdnerii, on adult catch. Transactions of the American Fisheries Society, 98(1):27–34, January 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wallis:1960:MSF

[Wal60] Orthello L. Wallis. Management of sport fishing in national parks. Transactions of the American Fisheries Society, 89(2):234–238, April 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Walburg:1961:NMA

[Wal61] Charles H. Walburg. Natural mortality of American shad. Transactions of the American Fisheries Society, 90(2):228–230, April 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Walburg:1963:PPR

[Wal63] Charles H. Walburg. Parent–progeny relation and estimation of optimum yield for American shad in the Connecticut River. Transactions of the American Fisheries Society, 92(4):436–439, October 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Walters:1969:GCS

[Wal69] Carl J. Walters. A generalized computer simulation model for fish population studies. Transactions of the American Fisheries Society, 98(3):505–512, July 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Warner:1962:CHR

[War62] Kendall Warner. Contribution of hatchery-reared salmon to the fishery of the Fish River Lakes, Maine. Transactions of the American Fisheries Society, 91(1):99–102, January 1962. CODEN TAF-SAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Warner:1963:NSS

[War63] Kendall Warner. Natural spawning success of landlocked salmon, Salmo salar. Transactions of the American Fisheries Society, 92(2):161–164, April 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Waters:1960:DPE

[Wat60] Thomas F. Waters. The development of population estimate procedures in small trout lakes. Transactions of the American Fisheries Society, 89(3):287–294, July 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Waters:1962:MEP

[Wat62] Thomas F. Waters. A method to estimate the production rate of a stream bottom invertebrate. Transactions of the American Fisheries Society, 91(3):243–250, July 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Waters:1964:RDS

[Wat64a] Thomas F. Waters. Recolonization of denuded stream bottom areas by drift. Transactions of the American Fisheries Society, 93(3): 311–315, July 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Watson:1964:DAY

[Wat64b] John E. Watson. Determining the age of young herring from their otoliths. Transactions of the American Fisheries Society, 93(1): 11–20, January 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Watson:1965:TMS

[Wat65] John E. Watson. A technique for mounting and storing herring otoliths. Transactions of the American Fisheries Society, 94(3): 267–268, July 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wells:1963:FBC

[WB63] La Rue Wells and Alfred M. Beeton. Food of the bloater, Coregonus hoyi, in Lake Michigan. Transactions of the American Fisheries Society, 92(3):245–255, July 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wilson:1969:EHD

[WB69] Dennis C. Wilson and Carl E. Bond. The effects of the herbicides Diquat(R) and Dichlobenil (Casoron(R)) on pond invertebrates. Part I. Acute toxicity. Transactions of the American Fisheries Society, 98(3):438–443, July 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wilton:1962:ECS

[WCH62] John W. Wilton, Ronald W. Crawford, and Don Hunsaker II. Effect of copper sulfate on angling success. Transactions of the American Fisheries Society, 91(4):424–425, October 1962. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weaver:1965:OSA

[Wea65] Charles R. Weaver. Observations on the swimming ability of adult American shad (Alosa sapidissima). Transactions of the American Fisheries Society, 94(4):382–385, October 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Webster:1962:ASF

[Web62] Dwight A. Webster. Artificial spawning facilities for brook trout, Salvelinus fontinalis. Transactions of the American Fisheries Society, 91(2):168–174, April 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weiss:1961:PEO

[Wei61] Charles M. Weiss. Physiological effect of organic phosphorus insecticides on several species of fish. Transactions of the American Fisheries Society, 90(2):143–152, April 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wells:1966:SDD

[Wel66] LaRue Wells. Seasonal and depth distribution of larval bloaters (Coregonus hoyi) in Southeastern Lake Michigan. Transactions of the American Fisheries Society, 95(4):388–396, October 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Welker:1967:MMC

[Wel67] Bill Welker. Movements of marked channel catfish in the Little Sioux River, Iowa. Transactions of the American Fisheries Society, 96(3):351–353, July 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Werner:1967:IMB

[Wer67] Robert G. Werner. Intralacustrine movements of blue-gill fry in Crane Lake, Indiana. Transactions of the American Fisheries Society, 96(4):416–420, October 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Westgard:1964:PBA

[Wes64] Richard L. Westgard. Physical and biological aspects of gas-bubble disease in impounded adult Chinook salmon at McNary Spawning Channel. Transactions of the American Fisheries Society, 93(3): 306–309, July 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Warner:1963:STF

[WF63] Kendall Warner and Owen C. Fenderson. The salmon and trout fishery of the Fish River Lakes, Maine. Transactions of the American Fisheries Society, 92(3):193–201, July 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Warner:1961:BSR

[WH61] Kendall Warner and Keith Havey. Body-scale relationships in landlocked salmon, Salmo salar. Transactions of the American Fisheries Society, 90(4):457–461, October 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

West:1966:IHC

[WH66] Jerry L. West and F. Eugene Hester. Intergeneric hybridization of centrarchids. Transactions of the American Fisheries Society, 95 (3):280–288, July 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Whitney:1969:ITB

[Whi69a] Richard R. Whitney. Inferences on tuna behavior from data in Fishermen’s logbooks. Transactions of the American Fisheries Society, 98(1):77–93, January 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Whitney:1969:SFR

[Whi69b] Richard R. Whitney. Schooling of fishes relative to available light. Transactions of the American Fisheries Society, 98(3):497–504, July 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Whitworth:1964:ORF

[WI64] W. R. Whitworth and W. H. Irwin. Oxygen requirements of fishes in relation to exercise. Transactions of the American Fisheries Society, 93(2):209–212, April 1964. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Warner:1969:CSS

[WI69] Kendall Warner and Angelo Incerpi. Current status of the salmon, Salmo salar, fisheries of two fish river lakes, Maine. Transactions of the American Fisheries Society, 98(1):45–51, January 1969. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wilimovsky:1969:D

[Wil69] Norman J. Wilimovsky. Discussion. Transactions of the American Fisheries Society, 98(2):350–351, April 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Winther:1966:JAT

[Win66] John K. Winther. Journal account of trout angling in Michigan, 1903–05. Transactions of the American Fisheries Society, 95(1): 101–104, January 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Waters:1961:ISB

[WK61] Thomas F. Waters and Robert J. Knapp. An improved stream bottom fauna sampler. Transactions of the American Fisheries Society, 90(2):225–226, April 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Welch:1963:TTD

[WKH63] Eugene B. Welch, Max Katz, and Marvin Hull. A trout with two dorsal fins. Transactions of the American Fisheries Society, 92(1): 67–68, January 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

William:1963:CEG

[WL63] M. William and Sue D. Lewis. Control of epizootics of Gyrodactylus elegans in golden shiner populations. Transactions of the American Fisheries Society, 92(1):60–62, January 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Webster:1960:OCL

[WLWY60] Dwight A. Webster, William A. Lund, Jr., Richard W. Wahl, and William D. Youngs. Observed and calculated lengths of lake trout (Salvelinus namaycush) in Cayuga Lake, New York. Transactions of the American Fisheries Society, 89(3):274–279, July 1960. CO-DEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wolfert:1963:MWT

[Wol63] David R. Wolfert. The movements of walleyes tagged as yearlings in Lake Erie. Transactions of the American Fisheries Society, 92 (4):414–420, October 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Woods:1969:FMB

[Woo69] C. S. Woods. A fish measuring board with double vernier. Transactions of the American Fisheries Society, 98(3):529–532, July 1969. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Warner:1960:EIB

[WP60] Kendall Warner and Ivan R. Porter. Experimental improvement of a bulldozed trout stream in Northern Maine. Transactions of the American Fisheries Society, 89(1):59–63, January 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Weis:1967:TMM

[WP67] Judith Shulman Weis and Alfred Perlmutter. Technique for microinjecting materials into the eggs of fishes. Transactions of the American Fisheries Society, 96(1):57–60, January 1967. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wagner:1962:SLL

[WS62] Wilbert C. Wagner and Thomas M. Stauffer. Sea lamprey larvae in lentic environments. Transactions of the American Fisheries Society, 91(4):384–387, October 1962. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wigley:1965:SFH

[WT65] Roland L. Wigley and Roger B. Theroux. Seasonal food habits of highlands ground haddock. Transactions of the American Fisheries Society, 94(3):243–251, July 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Ward:1963:TBS

[WV63] F. J. Ward and L. A. Verhoeven. Two biological stains as markers for sockeye salmon fry. Transactions of the American Fisheries Society, 92(4):379–383, October 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wagner:1963:SMR

[WWC63] Harry H. Wagner, Richard L. Wallace, and Homer J. Campbell. The seaward migration and return of hatchery-reared steelhead trout, Salmo gairdneri Richardson, in the Alsea River, Oregon.

Transactions of the American Fisheries Society, 92(3):202–210, July 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Whitmore:1960:ARS

[WWD60] Cecil M. Whitmore, Charles E. Warren, and Peter Doudoroff. Avoidance reactions of salmonid and centrarchid fishes to low oxygen concentrations. Transactions of the American Fisheries Society, 89(1):17–26, January 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Wood:1960:CHS

[WYHW60] E. M. Wood, W. T. Yasutake, J. E. Halver, and A. N. Woodall. Chemical and histological studies of wild and hatchery salmon in fresh water. Transactions of the American Fisheries Society, 89(3): 301–307, July 1960. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Yashouv:1963:IFP

[Yas63] A. Yashouv. Increasing fish production in ponds. Transactions of the American Fisheries Society, 92(3):292–297, July 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Yuen:1966:SSY

[Yue66] Heeny S. H. Yuen. Swimming speeds of yellowfin and skipjack tuna. Transactions of the American Fisheries Society, 95(2):203– 209, April 1966. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zein-Eldin:1961:PPE

[ZE61] Zoula P. Zein-Eldin. Plankton pigments in East lagoon, Galveston, Texas. Transactions of the American Fisheries Society, 90(1):32– 41, January 1961. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zein-Eldin:1965:EIB

[ZEK65] Zoula P. Zein-Eldin and Edward F. Klima. Effects of injected biological stains on oxygen uptake by shrimp. Transactions of the American Fisheries Society, 94(3):277–278, July 1965. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).

Zurbuch:1963:DLR

[Zur63] Peter E. Zurbuch. Dissolving limestone from revolving drums in flowing water. Transactions of the American Fisheries Society, 92(2):173–178, April 1963. CODEN TAFSAI. ISSN 0002-8487 (print), 1548-8659 (electronic).